### v2

# Nearby Special Lagrangians

Abstract
Let be a Calabi–Yau manifold and a compact connected embedded special Lagrangian without boundary. We combine methods from symplectic topology and geometric analysis, to study those compact irreducibly immersed special Lagrangians which are close to and have unobstructed Floer cohomology. We prove that if the fundamental group has no non-abelian free subgroup then every such Lagrangian is unbranched (that is, the projection is a covering map); moreover, if is abelian then is close to We prove a stronger uniqueness theorem when is finite. The conditions, the Floer cohomology condition and the special Lagrangian condition are all essential as we show by counterexamples.
Contents
1 Introduction
We begin with a definition we use throughout the paper.
Definition 1.1.
A Calabi–Yau manifold is a Kähler manifold equipped with a holomorphic volume form A special Lagrangian in a Calabi–Yau manifold is a Lagrangian of phase in the sense of Definition 2.32.
Remark 1.2.
The definition agrees with that in the literature on special Lagrangians except that Joyce [34] and others call an almost Calabi–Yau manifold to indicate that this need not be Ricci-flat. We do more symplectic topology in the present paper and therefore leave out the word almost so as not to be confused with almost complex structures. Special Lagrangians are in general defined as Lagrangians of any constant phase. We mention explicitly what the phases are if they are not
Let be a Calabi–Yau manifold and a closed connected embedded special Lagrangian. Closed means compact without boundary. The nearby problem for is solved by McLean [43]; that is, those special Lagrangians which are close to form a smooth manifold whose tangent space at is isomorphic to the (de Rham) cohomology group We study the nearby problem, which is to classify those closed special Lagrangians contained in a given neighbourhood of Donaldson and He [15, 29, 30] have proved recently an existence theorem about this. We prove on the other hand a uniqueness and non-existence theorem, Theorem 1.3 below.
We make now the definitions we use to state our results. Let be a Weinstein neighbourhood of (which may be embedded in the cotangent bundle Let be a Lagrangian immersion. We suppose that every self-intersection point of is a transverse double point unless otherwise mentioned. In many places we refer only to the image and call this simply the Lagrangian. We say that it is closed if is compact without boundary. We say that it is irreducibly immersed if the domain is connected.
We say that a Lagrangian immersion is unbranched if the composition with the projection map is a finite cover . This is equivalent to saying that there exists a finite cover such that the pull-back of under the induced map is a union of smooth perturbations of (if is embedded, this union is disjoint). Here a smooth perturbation means the graph over of some closed -form (so this is a Lagrangian). We say that the Lagrangian is branched if it is not unbranched.
We give also a brief account of Floer cohomology groups and Fukaya categories which we use. We assign to the symplectic manifold above a filtered strict category This is a version of the Fukaya categories defined by Akaho–Joyce [9] and Fukaya, Oh, Ohta and Ono [21] which we explain in §2. The objects of are closed Lagrangians in equipped with some additional data, the most important of which are bounding cochains in the sense of Fukaya, Oh, Ohta and Ono [21]. It is a property of the Lagrangians that they have these data; that is, some Lagrangians do have the additional data, in which case we say that they have unobstructed Floer cohomology or unobstructed for short, and others have no such data at all, in which case we say that they have obstructed Floer cohomology or obstructed for short. Floer cohomology groups are the morphism groups of the cohomology category of the strict category The symbol is a standard notation for these.
As we are working not in the whole but in the Weinstein neighbourhood the zero-section (and any Lagrangian that is close to it) has unobstructed in because this does not bound any non-constant holomorphic discs in (it has possibly obstructed in the Fukaya category of the whole ).
Also as is an exact symplectic manifold we can do the virtual counts in any characteristic [23]. The advantage of having is that we can cancel out the effect of pseudo-holomorphic discs whose contributions are multiples of for instance, with we can do without relative spin structures. Thus by varying we can include more Lagrangians with unobstructed. The zero-section has unobstructed for any because it has essentially no holomorphic discs. On the other hand, we shall need the nearby special Lagrangians to have unobstructed only for some In these circumstances our theorem may be stated as follows:
Theorem 1.3.
Let be a Calabi–Yau manifold and a closed connected embedded special Lagrangian; then the following hold.
Let the fundamental group be finite and let be any Weinstein neighbourhood of Then every closed special Lagrangian in with unobstructed is necessarily of phase and equal to
If has no non-abelian free subgroup, there exists a Weinstein neighbourhood of such that every closed irreducibly immersed special Lagrangian with unobstructed is unbranched. Moreover, if is abelian then is a McLean perturbation of
Here is a sketch of the proof. Suppose first that is abelian as in the latter part of (ii). As has unobstructed there is a corresponding object of the Fukaya category and we can define the Floer cohomology group which is a -graded vector space. Since the Lagrangian is special it follows that is non-zero and supported in a single degree; the non-zero property will be proved in Corollary 2.56, and the single degree property will be proved in Lemma 3.4. We use then the known techniques for dealing with which we will explain in §2. The key fact is that is a certain representation of the fundamental group and that this representation determines up to quasi-isomorphism. As is abelian there is a one-dimensional sub-representation of This may be realized by a McLean perturbation of with some local system over it. Then by a Thomas–Yau theorem, Proposition 2.67 (i), we can show that the original special Lagrangian agrees with the perturbation of
The same method applies to which has no non-abelian free subgroup as in (ii) above. In this case we shall need to pass to a finite-index subgroup of for the algebraic reason. Denote by the corresponding finite cover, and by the induced finite cover between the cotangent bundles. The McLean perturbation will then be defined in and the original will be the image of under So is unbranched.
Part (i) is proved in §3.1 with simpler algebraic arguments. The conclusion is stronger in two respects: we do not have to shrink the Weinstein neighbourhood and we can include special Lagrangians of any phase.
Apart from the hypothesis Theorem 1.3 (i) is stronger than the result of Tsai–Wang [61, 62] which we recall in Proposition 4.1. They prove this by a simple Riemannian geometry method, supposing that is Ricci-flat and Ricci-positive. Then by Myers’ theorem the fundamental group is necessarily finite as in (i) above. On the other hand, Tsai–Wang’s method applies to any singular special Lagrangians in the sense of geometric measure theory [42]. The hypothesis of our theorem presupposes that the Lagrangians are not very singular, about which we discuss more in Remark 3.18 after we prove the theorem.
The hypothesis on is essential to Theorem 1.3 (ii). For instance, if is a compact Riemann surface of genus then has free non-abelian subgroups and the cotangent bundle of such a surface admits many branched nearby special Lagrangians given by the hyperKähler rotations of Hitchin spectral curves [31, §5.1] (where the hyperKähler metric is given by Feix [17] and Kaledin [35]). We give in Example 5.12 a higher-dimensional version of this. He [29] constructs also branched special Lagrangians which are locally modelled upon the product of and a -fold spectral curve. His construction applies to many -homology -spheres [30] and their fundamental groups have non-abelian free subgroups (more precisely, his examples do not include finite quotients of or cf. Proposition 5.16). Thus no contradiction occurs with Theorem 1.3 (ii).
The abelian condition is essential to the latter part of Theorem 1.3 (ii). We prove in Corollary 4.7 that there are multi-valued graphs over a certain flat manifold This has a non-trivial finite cover by and the fundamental group is accordingly not abelian.
We show also that the other hypotheses of Theorem 1.3 are also indispendable. For this we use Lagrangian surgeries obtained from Morse -forms on We prove in Example 5.10 that there are examples satisfying the conditions of Theorem 1.3 (ii) except that the Lagrangians should be special. We prove in Corollary 5.3 that these Lagrangian are not special even in the weak categorical sense.
We examine now the hypothesis. This would in general be hard to verify directly (because the Maurer–Cartan equation is defined by using all the holomorphic curve moduli spaces). On the other hand, so far as we know there are no examples in which we cannot apply Theorem 1.3 just because the Lagrangians have obstructed (which would be hard to verify for the same reason). We change the point of view in Corollary 5.13: we use Theorem 1.3 to find examples with obstructed. These are branched special Lagrangians and we argue indirectly: if they had unobstructed, they would satisfy the conclusion of Theorem 1.3, which is impossible because they are branched; they have thus obstructed. This has a further consequence, Corollary 5.15, which is about Morse -forms and perhaps of interest in itself.
We begin in §2 with the more careful treatment of Fukaya categories. In §3 we prove Theorem 1.3. In §4 we recall Tsai–Wang’s result and give the examples related to it. This section is essentially about minimal submanifolds and geometric measure theory, which has nothing directly to do with Floer cohomology groups or Fukaya categories. In §5 we give the branched examples and prove the relevant results.
Acknowledgements. This paper is based upon the discussions we had in Fall 2016 at the Institute for Advanced Study in Princeton. We are both grateful to David Treumann for pointing out that spectral curves are the typical counter example to our results, to Dominic Joyce for helping us to find the examples with obstructed, and to the referees for their useful comments. The second author would also like to thank Mark Haskins and Martin Li for interesting conversations which inspired him to begin this work.
The first author was supported by the Erik Ellentuck Fellowship, the IAS Fund of Math, NSF grants DMS-1308179, DMS-1609148 and DMS-1564172, DMS-2103805 and by the Simons Foundation through its “Homological Mirror Symmetry” Collaboration grant. The second author was supported by the WPI project at the Kavli Institute for Physics and Mathematics of Universe, by the JSPS grants 16K17587, 18J00075 and 21K13788, and by the NSFC grant 11950410501.
2 Fukaya Categories
In this section we explain the definition and properties of the Fukaya categories. The key result is the split-generation theorem for cotangent bundles, including non-exact Lagrangians; see Theorem 2.84. This is known for exact Lagrangians and we extend it to non-exact Lagrangians. We must include non-exact Lagrangians because of the McLean perturbations, which are non-exact except the trivial perturbations.
We use pseudo-holomorphic curve moduli spaces to define the Fukaya categories. As we include non-exact Lagrangians we cannot follow the direct perturbation method of Seidel [53] and we follow instead the virtual perturbation method of Akaho–Joyce [9] and Fukaya, Oh, Ohta and Ono [21]. But still we cannot use their results as they are. The reason is that we shall be concerned with wrapped Fukaya categories and therefore need to include countably many Lagrangians. This causes a new problem which we explain now. Recall that we can deal only with finitely many kinds of Kuranishi spaces in doing the virtual counts [21, §7.2.3] and the results are accordingly a finite approximation of what we want. These are called gapped categories in our simplified notation, Definition 2.14 (or gapped algebras in their notation [9, 21]). We pass the the limit, called gapped categories, by the obstruction theory arguments as in the literature [9, 21]. But their statements are all about single or finitely many Lagrangians, so we make the version for countably many Lagrangians; see Theorem 2.26.
The gapped categories are curved categories, including the terms. We make them into strict categories by adding bounding cochains, following the references [21, 9]. We define then the wrapped Fukaya categories by the localization method, which has been used already in the literature [26] (the idea is due to Paul Seidel and the first author). The advantage of this is that we can make the definition only from the algebraic properties of Hamiltonian continuations, without going back to the study of pseudo-holomorphic curve moduli spaces.
We prove the split-generation theorem by a method of Paul Seidel and the first author (also from their unpublished work). This is related to the proof of Fukaya, Seidel and Smith [24] for exact Lagrangians (they assume that the ground field has characteristic ). They use Lefschetz fibrations, Seidel long exact sequences, and equivariant Fukaya categories, and we replace the last ingredient with the use of Viterbo restriction functors. When working with non-exact Lagrangians, the first two ingredients have already appeared in the literature, so we focus on a discussion of Viterbo restriction functors in the non-exact setting. After providing the argument for this generalisation, the remainder of the proof extends fairly immediately.
There is another method [2] of proving the split-generation theorem, using symplectic cohomology (for time-one periodic Hamiltonian orbits). One should then prove that it has the same relation as in the exact case to the algebraic invariants obtained from the Fukaya category. One could do this in closed symplectic manifolds (as will appear in [6]) and for monotone Lagrangians (as is done already in [49]) but not yet in the circumstances above.
We begin in §2.1 with the treatment of Novikov fields. The major difference from the references [21, 9] is that we extend the definition of Novikov fields used in the references [21, 9]; see Definition 2.2. We do this because we will use Lie–Kolchin’s theorem, Theorem 3.17, for the proof of Theorem 1.3. We shall therefore need the Novikov fields to be algebraically closed, and our definition will be suitable for this purpose.
In §2.2 we define the gapped categories and provide the obstruction theory arguments. Apart from the issue mentioned above, the major difference from the references [21, 9] is that we extend the definition of gapped algebras; see Remark 2.15. We do this because we include filtered local systems (in the sense of Definition 2.12) in the Fukaya categories.
The gapped structure will be used only for the process of defining the curved categories, which we shall in practice be able to treat as filtered categories (in the sense of Definition 2.14) forgetting the gap conditions.
In §§2.3 and 2.4 we define the curved category of a symplectic manifold In §2.5 we define the strict categories including the bounding cochains. In §2.6 we recall generalized Thomas–Yau theorems. In §2.7 we explain the effect of symplectic and Hamiltonian diffeomorphisms upon the Fukaya categories. In §2.8 we introduce the wrapped Fukaya category by the localisation method. In §2.9 we define the exact Fukaya category and show that its cohomology category is equivalent to that in the literature. In §2.10 we turn to the study of cotangent bundles; we prove two lemmas we shall need to deal with McLean perturbations, which we will use to prove Theorem 1.3 (ii). In §2.11 we prove the split-generation theorem for the Fukaya categories in cotangent bundles. In §2.12 we study the Yoneda functors obtained from the split-generation theorem.
2.1 Novikov Fields
We begin by recalling the definition of non-Archimedean valued fields. We call them valued fields for short.
Definition 2.1.
By a valued field we mean the pair of a field and a function called a valuation, such that and for we have and By a valued field extension of we mean a valued field such that is a sub-field of and the two valuations agree on We use the same symbol for different valuations.
Every valued field has a sub-ring defined by which is a local ring with maximal ideal in other words, is a unit if and only if
Let be a valued field and define a function by for Then for the strong triangle inequality holds and there is a distance function defined by making into a metric space We call a complete valued field if is complete.
Definition 2.2.
Define the minimum Novikov field to be the Novikov field in the literature [9, 21] without the Maslov index factor in their notation. We begin by recalling this. Let be a field and its group ring. Denote by the formal variable corresponding to so that every element may be written as with a finite subset of Define a function by Then the function defined by for satisfies the strong triangle inequalities, making into a metric space; and is its completion. The convolution products extend to the completion, making into a ring. This is in fact a field because of the inversion formula for with The function extends to the completion, making it into a complete valued field We denote an element of by an infinite sum with a discrete subset of
By a Novikov field we mean a complete valued field extension of We call the ground field of the Novikov field.
We prove now that algebraically closed Novikov fields exist for arbitrary characteristic. We give two proofs: one is Krasner’s lemma, Lemma 2.5, a general fact about valued fields, and the other is an explicit construction in Proposition 2.6 (ii).
Let be a valued field. Define a subring defined by This has a unique maximal ideal and the residue field The value group is a subgroup of defined by We recall the following basic facts:
Proposition 2.3.
Let be a complete valued field, and an algebraic extension; then the valuation of extends uniquely to so that:
if has finite degree, the valuation on is complete; and
if has infinite degree, the valuation of is strictly not complete.
Also, if is finite then
More precisely, there exist two intermediate fields of the maximal unramified field and the maximal tamely-ramified field such that:
if has characteristic then and if then is a power of
Proof.
We consult Neukirch [45]; the referred statements in what follows are all from Chapter II of this book.
The unique extension is by Theorem 4.8. By Proposition 4.9 the extended valuation is equivalent to the norm on a direct sum of whose elements have only finitely many nonzero components even if they are infinite dimensional. So they are complete if and only if finite-dimensional.
The inequality (2.1) is from Proposition 6.8. The field is defined in Definition 7.4, and in Definition 7.10 but under the hypothesis that has positive characteristic; if has characteristic we define The parts (i), (ii) and (iv) are then obvious. The part (iii) follows from Proposition 7.7; this is stated again in the positive-characteristic case, but its proof applies also to the characteristic-zero case. ∎
Here the extension is called the wild part of which is in general difficult to study. But it has degree a power of and for Galois extensions of such a degree, there is a classification result:
Proposition 2.4 (Artin–Schreier).
Let be any field of characteristic Then for every either:
the polynomial has a root in which case its roots are all in or
is irreducible, in which case is a degree- Galois extension of
Conversely:
every degree- Galois extension of is of the form for some or more generally
every finite Galois extension of of degree equal to a power of is a repeated extension of this kind.
Proof.
Finally we recall:
Lemma 2.5 (Krasner).
Every valued field has a complete algebraically-closed extension.
Proof.
We turn now to the second method. Let be a field. In Definition 2.2 we have defined the minimum Novikov field with ground field We define the maximum Novikov field by
where Recall that being well ordered means that every nonempty subset has a minimum. In particular, is half discrete in the sense that for every there exists such that
because we can take to be the minimum of The sums and products in are defined in the natural way: for we can define
because is well ordered; and
because is finite, which follows since are half discrete.
The valuation is defined by setting and by taking in (2.3) the least Then is a complete valued field. We have
We prove now the following proposition. Part (ii) with algebraically closed is the result we have promised to prove.
Proposition 2.6.
are both perfect.
is maximal in Krull’s sense [40]; that is, has no nontrivial valued field extension with and In particular, if is algebraically closed, so is
If is algebraically closed of characteristic so is
If has positive characteristic then is not algebraically closed, and its algebraic closure is strictly not complete.
Remark 2.7.
These results do not assert that is the algebraic closure of For rational-power series, the algebraic closure is described in Kedlaya [38].
Proof of Proposition 2.6.
For (i) let be either or By definition, a field is perfect if either it has characteristic or the ring homomorphism defined by is surjective. Suppose and consider the polynomial We show that this has a root in If it has root If we can write and Let be a root of which exists because is algebraically closed. Define
Then is a desired root.
The former part of (ii) is an old result [36, 40]. For the latter, if is algebraically closed, then is algebraically maximal and is already maximal in so has no nontrivial finite extension, or equivalently is algebraically closed.
The former part of (iii) is proved by Fukaya, Oh, Ohta and Ono [22, Lemma A.1] but we give another proof, using Proposition 2.3 with Suppose therefore that is a finite extension, and note that:
since is algebraically closed it follows that the inclusions in (ii) are all equalities, so
since it follows that the inclusions in (iii) are all equalities, so and
because has characteristic zero.
Consequently, that is, has no nontrivial finite extension, or equivalently, is algebraically closed.
For the latter part of (iii) let and consider the Artin–Schreier polynomial
If this had a root of the form then by induction on it must be of the form
so is not algebraically closed.
We show next that the algebraic closure of has infinite degree over By an Artin–Schreier element over we shall mean a root of an Artin–Schreier polynomial over Let be the field extension which contains all the Artin–Schreier elements over We prove the following three facts in this order:
if with and if are respectively roots of the polynomials then
if is an Artin–Schreier element over then where is a root of the polynomial for some with and every a root of the polynomial for some and and
if the algebraic closure of had finite degree over we should have where are such as in (II) and are under the same condition as in (II) but that may of course be different.
For (I) note that is a root of the polynomial so for some which implies as we want. For (II) write as a root of the polynomial and recall that has by definition only finitely many terms of negative powers with respect to that is, and Applying to this the part (I) repeatedly, we see that as we want. For (III) it is clear from definition that On the other hand, by our hypothesis is a finite extension and so is obtained from by adding finite Artin–Schreier elements. Hence using (II) repeatedly we get the other inclusion we want.
Now in (III) write as a root of the polynomial for some and Then for some the algebraic closure of so
But as is an infinite extension there is some that is not in Then by (2.11) the Artin–Schreier element a root of the polynomial is not in which contradicts the definition of
So the algebraic closure of has infinite degree over and by Proposition 2.3 (b) it is strictly not complete. ∎
Remark 2.8.
The rôle of Artin–Schreier polynonimals is essential to the proof above. In fact, if is algebraically closed then every finite normal extension is a repeated Artin–Schreier extension, as we show now. In the notation of Proposition 2.3 with we have again On the other hand, by (i) above, is automatically Galois. To this we can apply Proposition 2.4 so is a repeated Artin–Schreier extension as we want.
2.2 Gapped Categories
We introduce the notion of valued vector spaces.
Definition 2.9.
Let be a Novikov field. By a valued -vector space we mean the pair of a -vector space and a function called also the valuation, such that for we have and for we have where we use both and
Define a function by for Then the strong triangle inequality holds for and in particular, there is a distance function defined by making into a metric space We call a complete valued vector space if is a complete metric space. The valued vector space structure of extends to the metric space completion of making it into a complete valued vector space. We call this the completion of
Let be valued -vector spaces and a -linear map which is continuous with respect to the metric space structures above. Then there exists by that is, is the operator norm of the bounded operator We say that is filtered if We call it a filtered isomorphism if is a linear isomorphism with filtered. The latter is equivalent to or to because for linear maps between valued vector spaces.
Let be valued -vector spaces and a multi-linear map. Define then by
We say that is filtered if We denote by the set of filtered multi-linear maps from to The pair has the obvious structure of a valued -vector space. If is complete then so is
Let be a family of complete valued -vector spaces. Then there exists a valuation defined by The complete direct sum is the completion of the direct sum
We give a basic example of finite-dimensional valued -vector space.
Example 2.10.
Let be an integer and an -dimensional -vector space. We give a certain valuation. Let be a basis with which exists because for every Then has a valuation defined by for By the strong triangle inequality every valuation on satisfies for In particular, if is another basis with then for Swapping the rôles of we get the reverse inequality; that is, which we call the valuation on
We give three other equivalent definitions of the valuation on Let be any valuation. We use the subring defined in Definition 2.2. Regard as an module and define a submodule by Then is a submodule of and the quotient has the vector space structure over We prove that the following four conditions are equivalent: (i) is the valuation; (ii) is a rank- free module; (iii) is a finitely generated module; and (iv) the -vector space has dimension Clearly (i) implies (ii) and (ii) implies (iii). We prove that (iii) implies (iv). Take such that their images are a -basis. Then as is finitely generated we can use Nakayama’s lemma [41, Chapter X, Lemma 4.3] (for non-Noetherian local rings) which implies So and which is the condition (iv). We prove now that (iv) implies (i). Take such that their images are linearly independent over Then and we show that the equality holds. Otherwise there is some By re-ordering we can write where with Then
and have positive valuation. So in But none of is zero, which contradicts that are linearly independent. Thus (i)–(iv) are equivalent.
Remark 2.11.
Consider the category whose objects are -dimensional valued -vector spaces, whose morphisms are filtered homomorphisms. Then there is an equivalence from the category of rank- free modules to that of -dimensional valued vector spaces with valuations. The function between the object sets is defined by Each morphism is mapped to
When every valuation is the valuation. So the category of rank- free modules is equivalent to that of -dimensional filtered local systems.
We define then the filtered local systems.
Definition 2.12.
By a filtered local system over a topological space we mean one of non-zero finite-dimensional valued vector spaces whose parallel transports are filtered homomorphisms of valued vector spaces.
We make now the relevant definitions about discrete sub-monoids of
Definition 2.13.
Denote by the set of non-negative integers. Let be a discrete sub-monoid, that is, a discrete subset containing and closed under addition. By a decomposition of we mean with and with such that By a reduced decomposition of we mean and with and see Equation (2.15) below for a typical formula where this notion is used. Every decomposition with and for some defines indeed the reduced decomposition
Choose a total order on such that there is an ordered set isomorphism which we denote by and if are as above a decomposition of then
with equality if and only if at least of are equal to Setting and we get that is, is the least element of In other words, if and only if
We choose the order for every discrete sub-monoid so that if are two discrete sub-monoids with then the induced map will preserve the orders. Here is an explicit way of doing this. Set and for set
The maximum makes sense because is discrete and has the least positive element. It is clear that for For define by either or and Then we can check all the conditions above.
Let be a Novikov field, which we shall fix in this section. We define the notion of gapped categories, over The morphisms spaces are -graded. The structures are curved, including the terms.
Definition 2.14.
We begin with a filtered curved category which consists of an object set a family of -graded complete valued -vector spaces, and a family and such that
where and are any integers with For we have for every and (2.15) means
Suppose now given By a gapped category we mean an object set a family of -graded complete valued -vector spaces, a discrete sub-monoid and a family with and such that
where are as in (2.15) and the sum is taken over reduced decompositions ’s of
The operators are well-defined because by (2.13).
When the gapped category defines a filtered category of the same object set, the same morphism spaces, and the operators defined by where for and for The sum is well-defined because is complete.
Remark 2.15.
The definition of implies that every is a filtered homomorphism between valued vector spaces. This will be crucial to defining the Fukaya categories with filtered local systems. In the references [9, 21] they define to be linear over the residue field determined uniquely from the sum this is not true in our circumstances.
They also work with gapped algebras which are gapped categories with a single object; or they are expressible as the pair of a -graded complete valued -vector space and a family with satisfying the same equations. If is a gapped category then the complete direct sum, has a gapped algebra structure defined in the obvious way. Conversely, let be a gapped algebra with such that the induced map is zero unless and if this is the case then has image in Then and define a gapped category.
We define the truncations of gapped functors, and the full subcategories of gapped functors.
Definition 2.16.
Let with and a gapped category. The truncation of means the same object set the same morphism spaces and the subfamily these define a gapped category. By a full subcategory of we mean a subset the subfamily and the subfamily these define a gapped category.
We define the gapped functors,
Definition 2.17.
Let be two filtered categories. By a filtered functor from to we mean a function between the object sets and a family
and such that (i) for every the element has positive valuation, that is, and for with we have
where are as in (2.15) and are any integers with The right-hand side of (2.18) converges by (i) above.
Suppose now that are gapped categories, Note that any two discrete sub-monoids of are contained in some single discrete sub-monoid of By a gapped functor from to we mean a function between the object sets; a discrete sub-monoid with respect to which are gapped; and a family
with and such that
where are as in (2.16) and any decomposition of The operators are well-defined because by (2.13).
When the gapped functor defines a filtered functor between the corresponding filtered categories, with where for and for The sum is well-defined because is complete.
We define the identity functors of gapped categories,
Definition 2.18.
Let be a filtered category; we leave out for short. The identity functor is defined by taking the function to be the identity; for to be the identity for and for with
Suppose now that is a gapped category, Then the identity functor is defined by taking the function to be the identity; to be the identity for and for with
We define the composites of gapped functors,
Definition 2.19.
Let be filtered categories and filtered functors. Then the composite functor means the ordinary composite and the family with defined by
where are as in (2.18).
Suppose now that are filtered categories, and filtered functors. Then the composite functor means the ordinary composite a discrete sub-monoid with respect to which are gapped; and the family with and defined by
where are as in (2.20). This is independent of the choice of in the sense that for not in the smallest choice of
We define the homotopies of gapped functors, We make a rather strong hypothesis on the functions between object sets.
Definition 2.20.
Let be filtered categories and filtered functors. Suppose that induce the same function between the object sets, which we denote by for short. By a homotopy we mean a family
and such that (i) for every the element has positive valuation, that is, and (ii) for with we have
where are as in (2.18); and any integers with The right-hand side converges by (i) above.
Suppose now that are gapped categories, and gapped functors with the same function between the object sets. By a homotopy we mean a discrete sub-monoid with respect to which are both gapped and a family
with and such that
where are as in (2.20) and any decomposition of The operators are well-defined because are all by (2.13).
When the gapped homotopy defines a filtered homotopy between the corresponding filtered functors, with where for and for The sum is well-defined because is complete.
Remark 2.21.
We define the notion of being homotopic for gapped functors,
Definition 2.22.
Let be gapped categories and gapped functors which induce the same function between the object sets. We say that are homotopic if there is a homotopy from to
The following is a straightforward extension of Fukaya, Oh, Ohta and Ono’s result for gapped algebras.
Lemma 2.23 (cf. [21, Proposition 4.2.37]).
Let be gapped categories and suppose given a function Then being homotopic is an equivalence relation for those gapped functors from to inducing the given function ∎
We define the homotopy equivalences of gapped categories,
Definition 2.24.
Let be gapped categories of the same object set and a gapped functor which induces the identity on the object set. We call a homotopy equivalence if there exists a gapped functor which induces the identity on the object set and such that are homotopic to the identity functors of respectively. We say that are homotopy equivalent if there exists a homotopy equivalence
The following is also a straightforward extension of Fukaya, Oh, Ohta and Ono’s result for gapped algebras.
Lemma 2.25 (cf. [21, Corollary 4.2.44 and Remark 7.2.71]).
Being homotopy equivalent is an equivalence relation for gapped categories of the same object set.
Let be gapped categories of the same object set and a gapped functor which induces the identity on the object set. Then is a homotopy equivalence if and only if for any the co-chain map induces an isomorphism of the cohomology groups. ∎
Here is the obstruction theory result for gapped categories which extends that of Fukaya, Oh, Ohta and Ono [21, Theorem 7.2.72].
Theorem 2.26.
Let be a gapped category, and a full subcategory of the truncation of Let be a gapped category with and a homotopy equivalence which induces the identity on the object set. Then there exist a gapped category with which contains as a full subcategory in its truncation; and a gapped functor which induces the identity on the object set, whose truncation agrees with on
Proof.
Recall from definition that the truncation of is a family of co-chain complexes. We assign to a co-chain complex and a co-chain map For set and For with or set and define to be the identity map. In both cases the co-chain map induces an isomorphism of the cohomology groups.
We extend these co-chain complexes to a gapped category and these co-chain maps to a gapped functor where is the truncation of We do this by an induction on Suppose therefore that we have defined for with and for We extend these to a gapped category and a gapped functor
Denote by the element with We assign to some filtered homomorphisms which satisfy the equations. For we use the given and otherwise we introduce some new Define a degree-one -linear map from to itself by
This is well-defined because implies And Define an obstruction co-cycle by
where or Doing the same computation as Fukaya and others [21, (4.5.2.1)] we find that is, is certainly a co-cycle. Their computation [21, (4.5.2.4)] shows also that the homotopy equivalence induces an isomorphism from the -cohomology group to the -cohomology group mapping the cohomology class of to that of But the latter vanishes for the obvious reason: is already an category. Thus is a co-boundary, written as for some
Define then an obstruction co-cycle in by sending to
where or and are a decomposition of with so that As is a co-chain equivalence we can modify if we need, so that the corresponding obstruction co-cycle will vanish. So there is a gapped functor as we want, completing the induction.
We have thus obtained a gapped category and a gapped functor that is, are given for every with Extend these to with by repeating the obstruction theory arguments above; then the proof is complete. ∎
2.3 Fukaya Categories
We begin with the definition of ambient spaces in which we do Floer theory. The example we have in mind is cotangent bundles, but we make a more general definition:
Definition 2.28.
A compact symplectic manifold with contact boundary is a symplectic manifold which is given a -form near such that and the vector field given by is outwards pointing on . We call a Liouville -form of
In these circumstances there exists near a unique smooth function with values in such that
We call the radius function near We say that an almost complex structures is compatible with the contact boundary if it is compatible with in the usual sense, near and is invariant near under the flow of
If there exists a global -form with and satisfying the conditions above then we call a Liouville domain.
We consider the following class of Lagrangians:
Definition 2.29.
Let be a compact symplectic manifold with contact boundary, and a Liouville -form of it. Let be a compact manifold possibly with boundary and a Lagrangian immersion. By an abuse of notation we shall usually mention only the image in place of the immersion map, and say that is compact if so is
We say that is generic if its every self-intersection point is a transverse double point outside We say that has Legendrian collar if near and has regular value (this is automatically true if is closed).
Suppose now that and are two Lagrangians in given by immersions and We say that are generic if the immersion is generic.
We recall the notion of -graded Lagrangians in -graded symplectic manifolds. There are several ways of doing this [51]. We choose that which uses almost complex structures on the ambient symplectic manifold.
Definition 2.30.
A -grading of is the choice up to homotopy of a nowhere-vanishing section of the squared canonical bundle of relative to some -compatible almost complex structure
Remark 2.31.
A -grading of exist if and only if where denotes the first Chern class of
Given a quadratic volume form representing a -grading on , we associate to each Lagrangian immersion a unique smooth function with Here is the squared volume form of with respect to the metric induced from To define we do not need to be orientable.
Definition 2.32.
We say that a Lagrangian immersion is -graded if this is equipped with a lift of to the universal cover We say that is special of phase with respect to if is locally constant. When is connected, must be constant and we shall say has phase We shall also use this constant function for the grading of The shift operator maps the -graded Lagrangian to
Remark 2.33.
The Lagrangian immersion admits a -grading if and only if the group homomorphism vanishes.
The definition above of special Lagrangians is more general than that in the literature [34]. These agree when is integrable and is the square of some holomorphic volume form.
Definition 2.34.
Let be a generic Lagrangian immersion. We consider the fibre product
which consists of the diagonal identified with and the self-intersection pairs in Suppose that is -graded by a phase function and we define then a Maslov index function Set on the diagonal We associate to each pair an integer called the Maslov index. Put and take a -linear isomorphism which maps to to and to for some These are unique up to order so we can define
The definition of phase functions implies readily that is in fact an integer.
Suppose now that are two generic Lagrangians, -graded by and respectively. The fibre product is then a finite set and we define the Maslov index function by sending to the right-hand side of (2.27) above with in place of respectively.
Remark 2.35.
Example 2.36.
Take and Define a -grading by Take and For define
so Define a -grading by starting with on and lifting the path Then and in particular The Maslov index of at the intersection point is now equal to This agrees with the Morse index of the quadratic function on Our convention is thus the same as Seidel’s [51, §2d, (v)].
We extend the definition of relative spin structures to countably many Lagrangians. These will equip the pseudo-holomorphic curve moduli spaces with natural orientations.
Definition 2.37.
Let be a symplectic manifold, a compact manifold and a generic Lagrangian immersion; we allow to have boundary. By a relative spin structure of we mean the data where is a triangulation of an oriented vector bundle over the -skeleton of an orientation of a triangulation of such that induces a simplicial map and a spin structure of the bundle restricted to the -skeleton of
Suppose now that is a sequence of compact manifolds which may have boundary, and a generic Lagrangian immersion. By a relative spin structure of we mean a sequence such that each is a relative spin structure of each is a subcomplex of and the restriction of each to agrees with
Remark 2.38.
Suppose that that is defined over the whole and that each is a spin structure of over the whole Then each Stiefel–Whitney class is the pull-back of In this case we refer to as the background class of the relative spin structure.
We can in practice take to be embedded in the cotangent bundle over a closed oriented manifold and take where is the projection. So the background class is and the zero-section will be given a spin structure of
We introduce now the notion of branes:
Definition 2.39.
Let be a compact symplectic manifold with contact boundary, and a Livouville -form of it. By a brane on we mean a pair where
is a compact -graded generically-immersed Lagrangian with Legendrian collar; and
is a filtered local system over the domain of the immersion
We write only in place of when we want to save notation.
We choose countably many branes and a Novikov field.
Hypothesis 2.40.
Let be a compact symplectic manifold with contact boundary, and a Louville -form of it. Let be a countable set of branes such that any two distinct Lagrangian immersions in are generic in the sense of Definition 2.29; we allow some of the Lagrangian immersions to be the same. Fix a Novikov field When has characteristic we suppose that has no non-constant pseudo-holomorphic sphere (cf. [23]). When has characteristic we suppose that and the underlying Lagrangians have a relative spin structure with respect to some order of the Lagrangians.
We introduce also the homologically perturbed Floer complexes:
Definition 2.41.
Let be as in Hypothesis 2.40 and two elements. Denote by and the projections. The homologically perturbed Floer complex associated to the pair is the -graded finite-dimensional valued -vector space
where denotes the degree shift by Maslov indices. Since the fibres of are valued vector spaces it follows that so are and
Remark 2.42.
The definition (2.29) is actually the result of applying the filtered homological perturbation lemma [21, Theorem A] to the singular or de Rham model in the literature. More precisely, in any construction of Lagrangian Floer cohomology groups we have to choose a chain model of the Floer complex with Akaho–Joyce [9] and Fukaya, Oh, Ohta and Ono [21] both take the singular chain model. Fukaya [19] takes the de Rham model assuming that the ground field contains
We come now to the construction of curved Fukaya categories.
Theorem 2.43.
Let be as in Hypothesis 2.40 and put Then has the structure of a gapped category whose morphism spaces are given by (2.29) and which has the following properties:
the structure is curved in such a way that if if has no self-intersection point of index or and if does not bound any non-constant pseudo-holomorphic discs, then has curvature
for a pair , the differential has leading term and
for a triple which are supported on the same Lagrangian, the leading term of the product map
(2.30) vanishes with the following exceptions: on the diagonal component of this agrees with the usual cup product map
and on each self-intersection pair it is given by the map
2.4 Proof of Theorem 2.43
Definition 2.44.
Let be a manifold, and a filtered local system over The smooth singular chain complex is the module generated by pairs where is a smooth singular -simplex in and a (flat) section of of nonnegative valuation, modulo the relation
The differential on is defined by
where is the face inclusion map.
If has boundary, we define We regard this, by Poincaré duality, as a cochain complex.
Let be a compatible almost complex structure on in the sense of Definition 2.28. We introduce now the moduli spaces of -holomorphic discs. Let be underlying Lagrangians in
Definition 2.45.
We denote by the compact moduli space of finite-area stable -holomorphic discs in with boundary points, whose corresponding segments are ordered counterclockwise and lifted to More precisely, for each genus-zero prestable bordered Riemann surface we choose a continuous map as Akaho–Joyce do [9, Definition 4.1]; this is essentially unique. Denote by the points corresponding to the marked points on and by the domains of the immersed Lagrangians. For each holomorphic curve we choose a continuous map as Akaho–Joyce [9, Definition 4.2]. We require that the connected components of should map to respectively.
For we denote by the subset made from -holomorphic curves of area
We recall the maximum principle on Riemann surfaces with Neumann boundary conditions, under the assumption that the almost complex structure of compatible with the contact boundary:
Lemma 2.46 (Abouzaid–Seidel [8, Lemma 7.2]).
There exists a compact set of which contains the image of every non-constant -holomorphic curve included in ∎
Hence by the Gromov compactness theorem we get
Corollary 2.47.
Every is compact. ∎
Suppose now that are given local systems respectively. Note that for every there is an evaluation map
where So has local systems of the form for Using the parallel transport map we get a natural map of local systems
Let be a discrete sub-monoid which contains the areas of every pseudoholomorphic disc with boundary on the underlying Lagrangians of The gapped categories in what follows will be gapped with respect to We fix an order on the countable set and use the same order on the underlying Lagrangians of it. Here are the next three steps:
For each construct a gapped category whose objects are the first Lagrangians and whose morphism space between is (2.33). We denote this morphism space by As Fukaya, Oh, Ohta and Ono [23, Remark 13.5] point out we can work with the whole complex (2.33), not with countably generated sub-complexes of it as in the earlier works [9, 21].
Construct a homotopy equivalence from to the truncation of
Lift the gapped structure of to the gapped structure.
We do (I) as follows. By Corollary 2.47 every moduli space of the form is a finite union of Kuranishi spaces, which are oriented naturally by the relative spin structure. Suppose that for each we are given We choose a multivalued perturbation of the fibre product
where whenever Combining this perturbation with (2.35) we get an element of which we call the virtual chain for It has coefficients in the ground field by the result of Fukaya, Oh, Ohta and Ono [23]. Varying we get a map
One can prove that there is a consistent choice of the virtual chains above such that defines the gapped structure.
Step (II) is done by a standard cobordism (or bifurcation) argument. Step (III) is a consequence of Theorem 2.26. Once these have been done, we get a gapped category whose truncation is homotopy equivalent to Applying the homological perturbation lemma to we get a gapped category that we want. ∎
2.5 Bounding Cochains
We define the units and bounding cochains of filtered categories.
Definition 2.48.
Let be a filtered category. By units of this we mean a family of degree-zero morphisms such that for for with for with and for with and for containing at least one of
Let be a filtered category. We call this a strict category if for every If it is strict then define the cohomology category to be the associative category consisting of the same objects; the morphism space for and the product map induced from the operators. Define also the category to be the subcategory of consisting of the same objects and whose every morphism space is the degree-zero part
Let be filtered strict categories and a filtered functor. We call a strict functor if This induces then a functor between the cohomology categories, in the obvious way.
Let be any and a degree-one element with We call a bounding cochain if The sum converges because We define then a filtered strict category Define to be the set of pairs with and a bounding cochain. For set For define by
The sum converges because Since it follows that that is, is a filtered homomorphism. The equations of imply those of Since it follows that Thus is a filtered strict category.
If is a filtered category with units then the strict category corresponding to it has units too. This is a straightforward extension of the algebra case [21, Lemma 5.2.17].
We introduce now the strict category where ‘nc’ stands for not necessarily closed Lagrangians.
Definition 2.49.
We define the strict category to be that made from by using bounding cochains as above. We denote an object of by a triple where is an object of and a bounding cochain in We say that is supported on or that underlies We define the -graded -linear associative category to be the cohomology category of We denote by its morphism space between two objects
The following is a straightforward extension of the result of Fukaya, Oh, Ohta and Ono [21, §7.3] for embedded Lagrangians.
Theorem 2.50.
The filtered categories have units. ∎
Remark 2.51.
Strict unitality fails at the level of singular chains, because the perturbations of Kuranishi spaces of the form (2.36) need not be compatible with forgetting marked points. One can however prove that for there exists a homotopy unit of whose leading term is equal to the fundamental class of This may be done by taking fundamental chains of the underlying Lagrangians in and constructing in a consistent way the homotopies between those perturbed Kuranishi spaces for which some of in (2.36) are the fundamental chains, and those obtained from the forgetful maps. The homotopy units become the ordinary units after passing to the minimal models.
We recall the notion of zero objects in and give a corollary to Theorem 2.50.
Definition 2.52.
We call a zero object if the unit of is zero.
Corollary 2.53.
If is a zero object, the Lagrangian has at least one self-intersection point of index
Proof.
Since is -graded it follows that the unit in may be cancelled out only by elements of which come from self-intersection points of of index ∎
Corollary 2.53 gives another proof of the following fact.
Corollary 2.54 (Fukaya, Oh, Ohta and Ono [21, Theorem E]).
Every supported on an embedded Lagrangian is non-zero. ∎
Take as in Definition 2.30 so that we can speak of special Lagrangians in Consider nearly special Lagrangians which have phase sufficiently close (in the sense) to a constant function. The estimate (2.28) implies then
Proposition 2.55.
If is supported on a generically immersed nearly special Lagrangian then the Floer cochain group supported in degrees and accordingly, so is the cohomology group ∎
This and Corollary 2.53 imply
Corollary 2.56.
Every supported on a nearly special Lagrangian is non-zero. ∎
We define the Fukaya category which we use most often.
Definition 2.57.
We denote by the full subcategory of objects supported on closed Lagrangians; that is, the underlying Lagrangian of an object of is a Lagrangian immersion from a compact manifold without boundary.
We prove now a Poincaré duality theorem. We are in the right context for doing so; but we shall not need the result for the later treatment, and the reader in a hurry can proceed safely to Proposition 2.66 below.
Theorem 2.58.
Let be an object of whose local system has one-dimensional fibres. Then there exists a nondegenerate pairing
Proof.
We begin by recalling the notion of filtered bimodules. Let be a -graded -linear filtered curved algebra. By a bimodule over we mean a -graded module equipped with operations
which satisfy the bimodule version of the equation. An bimodule bimodule is filtered if it is equipped with a valuation, so that bimodule operations do not decrease valuations.
Example 2.59.
The algebra itself has the structure of a diagonal bimodule with operations
Example 2.60.
If is an bimodule over then its dual has the structure of a bimodule given by
and
Next let and apply to the single brane the proof of Theorem 2.43. Then we get a gapped algebra We use the same notation for the diagonal bimodule, and denote by its dual bimodule. Recall that the of has leading term equal to the usual differential, whose cohomology group is
We prove
Proposition 2.61.
There exists a -graded -linear natural isomorphism of graded vector spaces
Proof.
If has characteristic then the condition that be relatively spin includes the data of an orientation of the domain of the immersed Lagrangian , so there is a Poincaré duality isomorphism
where we do not have to change the local-system term because For an off-diagonal self-intersection pair there is also a pairing
which completes the proof. ∎
The following is the heart of the proof of Theorem 2.58.
Lemma 2.62.
Proof.
Recall that is by construction an inverse limit of some gapped algebras, which we denote by Regard this as a diagonal bimodule with operations of the form
and For each we construct a bimodule map which consists of a collection of
indexed by nonnegative integers and . We regard the output as an input by duality, so that we have to construct a compatible collection of maps
Here we use the hypothesis that the local system of has one-dimensional fibres.
In order to construct this map, we use the stable moduli spaces of pseudoholomorphic discs with boundary points, ordered as in Figure 2.1: the marked point on the right corresponds to the input, and on the left to the output.
We choose inductively a family of perturbation data for the fibre products of the form
where denotes the evaluation maps at the boundary marked points ordered along the boundary and are singular simplices in with We make these compatible with the previous perturbations of (2.36) and those for the lower-order maps relative to Then we get a bimodule homomorphism, which we denote by
From the bimodule maps of this form we construct an bimodule map by the inductions and obstruction theory arguments. The algebraic framework for this is given by Fukaya, Oh, Ohta and Ono [21, §5.2], used originally for the deformation invariance of Floer cohomology groups; which is in fact a homotopy equivalence of two bimodules [21, Theorem D].
Recall that we have constructed an homotopy equivalence between and the truncation of From this we get an algebra which is homotopy equivalent to and whose trunation is The bimodule map which induces an bimodule map (over the algebra )
which we denote still by In the previous stage we have constructed an bimodule map
This is homotopic to the reduction of (2.51) by a choice of homotopy between the defining perturbation data. Using homological perturbation, and proceeding by induction, we get the desired system of bimodule maps. ∎
Corollary 2.63.
is a homotopy equivalence of bimodules.∎
The bimodule homotopy equivalence induces for every bounding cochain another bimodule homotopy equivalence
relative to the -deformed structure; and for the term
is given by
Hence, passing to the cohomology groups, we get an isomorphism
which completes the proof of Theorem 2.58. ∎
Theorem 2.58 implies
Corollary 2.64.
Let be a non-zero object of whose local system has one-dimensional fibres. Then ∎
Remark 2.65.
This is proved already for embedded Lagrangians without local systems by Fukaya, Oh, Ohta and Ono [21, Theorem E]. The second author [33, Lemma 4.4] gives two other proofs, including immersed Lagrangians. One uses the cyclic symmetry [19] assuming that the ground field contains The other uses open-closed maps [21, Theorem 6.4.2].
We end by recalling a result of the second author [33, Corollary 4.7 (ii)] which we state in a slightly different way because we include local systems.
Proposition 2.66.
Let be isomorphic in and supported on closed special Lagrangians of phase respectively which need not be modulo Suppose Then ∎
Proof.
Corollary 2.64 gives a sufficient condition for the hypothesis above.
2.6 Thomas–Yau Theorems
We recall a few generalizations of Thomas–Yau’s theorem proved by the second author [33]. In this subsection we work with Calabi–Yau manifolds and integrable complex structures.
Proposition 2.67.
Let be a Calabi–Yau manifold of complex dimension Let be supported on closed irreducibly immersed special Lagrangians of phase respectively which need not be modulo Suppose that either or embedded. Then the following hold.
If then either or is supported in degrees
If are isomorphic in with then
Remark 2.68.
The point of (i) is that the degrees are strictly and It follows immediately from (2.28) that they belong to
Proof of Proposition 2.67.
Let be embedded. Then we can work in a Weinstein neighbourhood of which is a Stein manifold. This is the condition (ii) of the second author’s result [33, Theorem 6.2], which implies (i) above. We prove (ii) now. By Proposition 2.66 the two special Lagrangians have the same phase up to shift. Hence it follows by (i) that either or is supported in for some But the latter is impossible because the graded vector space has two non-trivial degrees with difference So ∎
We recall also the following version. We say that a Calabi–Yau manifold is real analytic if the Kähler form of is real analytic with respect to the underlying real analytic structure of the complex manifold
Proposition 2.69.
Let be a real analytic Calabi–Yau manifold and two closed irreducibly-immersed Lagrangians which are, near both special and graded by Then either agree with each other near or any pair of objects supported on them have Floer cohomology supported in degrees
Proof.
Suppose and denote by the set of intersection points at which have at least one common tangent space. The second author [33, Theorem 5.1(i)] proves then that there exists a neighbourhood of and arbitrarily-small Hamiltonian perturbations of respectively which intersect generically each other with no intersection point in of index or On the other hand, these have no intersection point of index or outside [33, Lemma 2.7]. So they have no intersection point at all of index or Since are both graded by it follows by (2.28) that every intersection point of the two perturbed Lagrangians has degree in So their Floer complex is supported in degrees which completes the proof. ∎
2.7 Hamiltonian Continuations
We return now to the symplectic topology context. We recall the definition of homotopies between strict functors. Let be strict categories. Then we can define the functor category which is a strict category whose objects are strict functors from to We say that two objects are homotopic if they are isomorphic in Here we do not impose any conditions on the object sets as we do in Definition 2.20.
We say that a strict functor is a homotopy equivalence if there exists such that is isomorphic in to the identity on and isomorphic in to the identity on Seidel [53, Theorem 2.9] proves that is an homotopy equivalence if and only if the induced functor is an equivalence of categories.
We explain now the effect of a symplectomorphism of upon the Fukaya categories. Recall from Definition 2.28 that the radius function is defined near Note also that every symplectictomorphism acts naturally upon the objects of
Lemma 2.70.
Let be a symplectomorphism isotopic to the identity and which agrees near with the Hamilotonian diffeomorphism of for some Suppose that induces a bijection of Then there exists an equivalence
such that for every the image is given by a certain bounding cochain for The functor is compatible with composition in the sense that, for a pair and of boundary-linear symplectormorphisms, the composite is naturally isomorphic to
Proof.
The strict category is made from a gapped category whose definition involves the choice of an almost complex structure and other data. We can then define another gapped category by using We want to show that the two gapped categories are homotopy equivalent. This is done by Fukaya, Oh, Ohta and Ono [21, Theorem A] and by Akaho–Joyce [9, Theorem 11.2(a)] for gapped algebras by interpolating between the almost complex structures; we do so while maintaining compatibility with the contact boundary, thus ensuring compactness. They use the obstruction theory arguments for gapped functors [21, Lemma 7.2.129] for gapped algebras. We extend them to gapped categories in the same way as in the proof of Theorem 2.26. Then we get a homotopy equivalence of the two gapped categories. Fukaya [20, Proposition 13.18] proves that this induces a homotopy equivalence of the corresponding strict categories. There is thus an equivalence (2.57) of the cohomology categories.
To find the natural isomorphism between and we begin with homotopies of gapped categories. We lift them to gapped categories by an obstruction theory for homotopies of homotopies. This is done by Fukaya, Oh, Ohta and Ono [21, Corollary 7.2.219] for gapped algebras. We extend it in the same way as in the proof of Theorem 2.26. Then we get a homotopy between and as gapped functors. This induces again a homotopy between the strict categories, which induces in turn a natural isomorphism that we want. ∎
There is more to say if the symplectomorphism in Lemma 2.70 is a Hamiltonian diffeomorphism:
Proposition 2.71.
Let be the Hamiltonian diffeomorphism of a smooth function which agrees near with for some Suppose that induces a bijection of Then there exists a natural transformation from to the identity functor; that is, for any there is a commutative diagram
If either or are supported on a compact Lagrangian, then these arrows are isomorphisms. If there is another such Hamiltonian diffeomorphism then
Proof.
We explain how to deduce the statement from the result of Fukaya, Oh, Ohta and Ono [21, Theorem G (G.4)] who prove that pairs of Hamiltonian diffeomorphisms induce bimodule homomorphisms of Floer groups (Akaho–Joyce explain in [9, Theorem 13.6] how to extend this to immersed Lagrangians, and it is also straightforward to include local systems). As in Lemma 2.70, the construction of the bimodule homomorphisms involves pseudo-holomorphic curve equations with respect to varying almost complex structures on the target, and we require their compatibility with the contact boundary to ensure compactness.
Define to be the image of the unit of under the bimodule homomorphism induced by the pair of Hamiltonian diffeomorphisms. We prove then that the diagram (2.58) commutes. The pair of Hamiltonian diffeomorphisms induces the bimodule homomorphism In particular, this is a left module homomorphism and there is a commutative diagram
Consider now the direct sum If these are supported on distinct Lagrangians respectively then regard as an object of supported on If they are supported on the same Lagrangian then regard it as an object of with local system where are the local systems of respectively. In either case the pair induces a bimodule homomorphism In particular, this is a right module homomorphism and there is a commutative diagram
We look at the following component, which commutes too:
This combined with (2.60) implies the commutative diagram (2.58).
We finally prove (2.59). Using (2.61) with in place of in place of and in place of we get a commutative diagram
Consider the unit of and and look at the image of this pair. Then we see that
But Fukaya, Oh, Ohta and Ono [21, Theorem G (G.4)] prove so the right-hand side of (2.63) is equal to This completes the proof. ∎
2.8 Wrapped Fukaya Categories
In general, if is an category and a set of morphisms, there exists a universal pair consisting of an category and an functor such that every element of becomes invertible in The key property that we shall use is the fact that this construction is compatible with passage to triangulated categories in the sense that there is a natural equivalence of triangulated categories
where the right-hand side is the localisation of a triangulated category.
The next result is a key computational tool for computing Fukaya categories from localisations, and also implies the above equivalence of triangulated categories in this specific context:
Proposition 2.72.
Let be an category, and an equivalence. For objects write the morphism space of the pair Let be a natural transformation from to the identity functor; and the localisation of by the set . Then for any two objects there exist natural isomorphisms
where is an inverse to Also for there exists a commutative diagram
where the horizontal arrows are products maps and the vertical isomorphisms are obtained from those of (2.65).
Proof.
Assuming that the category is triangulated, the localisation may be constructed as the categorical quotient of by the subcategory consisting of the cones of all morphisms . Morphisms in the quotient category are then computed by the direct limit of the groups as proved in [55, Lemma 7.18]. The fact that this is isomorphic to the direct limit of the groups is then a consequence of iteratively applying the functor, and the expression in terms of the direct limit of the groups is implied by the constancy of the colimit with respect to either variable. The expression for the product is an immediate consequence of functoriality. ∎
We shall also need the following observation:
Lemma 2.73.
Assume that are two sets of morphisms such that for each element , there exists a morphism such that is a composition of morphisms in . Then the natural functor is an equivalence.
Proof.
It suffices to prove that satisfies the universal property of the localisation away from , i.e. that every morphism in is already invertible in . This is immediate from the assumption because a composition of morphisms in becomes invertible in , and the only way for a product of morphisms to be invertible is if both factors are so. ∎
We now arrive at the definition of the wrapped category in the contact-type setting:
Definition 2.74.
Let be a set of Hamiltonians which are linear at the boundary in the sense that they agree with for some constant near . Assume that this set contains Hamiltonians whose slopes are arbitrarily large. Identify these with the time-one maps they generate, and suppose that if and then Then is defined as the localisation of by the set of continuation morphisms.
As a consequence of Lemma 2.73 this definition does not depend on the set of Hamiltonians which are chosen: if we add or subtract a Hamiltonian diffeomorphism to the set , then the resulting localisation does not change because of the existence of continuation maps to Hamiltonians of larger slope.
The closed Fukaya category is embedded naturally as a full subcategory of because the continuation morphisms are invertible in
2.9 Exact Lagrangians
Suppose now that the Liouville -form is given globally on so that we can speak of exact Lagrangians in Denote by the ground field of the Novikov field. Denote by the full subcategory of those objects for which:
the Lagrangian is embedded and exact; that is, has no self-intersection point and is an exact -form;
the local system is of the form for some local system on and
is given the trivial bounding cochain, which exists because is embedded and does not bound any noncontant pseudoholomorphic disc.
For take any primitive function with For we define the action function
by where are the primitives of respectively. We recall:
Proposition 2.75.
Let and their primitives respectively. As in the proof of Theorem 2.43 take on a compatible almost-complex structure define the moduli space and define the evaluation map
Then the fibre of this map over a point consists of -holomorphic discs of area
Proof.
This implies that the structure of is reducible to :
Corollary 2.76.
has the structure of a -linear category whose hom space between two objects is given by
and whose structure maps are defined by the same disc counts but without the factor. The -linear extension of this category is isomorphic to that of the subcategory with the same objects, under the maps
defined as the multiplication by on the component of each ∎
Proof.
Choose now a countable set of Hamiltonians as in Definition 2.74. Denote by the localisation of with respect to this. Using (2.72) we regard as a full subcategory of
Consider now a Fukaya category with the same object set as but defined by direct perturbations [53]. Denote by the localisation of this with respect to the same Hamiltonians. We prove that is equivalent to We consider only the cohomology categories rather than the categories, which will do for the treatment in the later sections.
Proposition 2.77.
is equivalent to
Proof.
We define a functor We leave out the local systems to save notation. For two exact embedded Lagrangians we define a -linear map Suppose either or has no boundary. Then is defined as for some Hamiltonian diffeomorphisms with compact support in More precisely, the Floer group is defined by the direct perturbations. But these may be regarded as a different choice of virtual chains, so the resulting Floer groups will be canonically isomorphic. This is done by Fukaya, Oh, Ohta and Ono [21, Theorem G] (in the context of the present paper the similar arguments have appeared in the proof of Lemma 2.70). There is thus a canonical isomorphism There is another isomorphism from which we get an isomorphism that we want.
Suppose have both Legendrian collar. Then the morphisms of form the wrapped Floer groups and Since the isomorphism exists for every we get an isomorphism
We prove that the -linear maps thus defined are compatible with the products. Let and suppose first that at least two of have no boundary. Write and Put Recall that the isomorphism is a bimodule isomorphism with respect to and It is in particular a left module isomorphism with respect to so the diagram
commutes. The upper arrow contains the product map and the lower arrow contains the product map These are thus compatible.
Suppose now that at least two of have Legendrian collar. Then and The isomorphisms and are compatible with the product maps, completing the proof. ∎
2.10 Cotangent Bundles
Let be a compact connected Riemannian manifold without boundary, and the unit disc sub-bundle. Recall that the wrapped Floer group is computed by Abbondandolo, Portaluri and Schwarz [1]; they construct a -graded ring isomorphism
where denotes the based-loop space and is non-positively graded; We shall not need an lift of (2.73) as treated by the first author [4].
To define the Fukaya categories we do the following:
We take the background class on to be the pull-back by the projection of the Stiefel–Whitney class So the zero-section has a relatively spin structure, and we give it a natural one which we shall define shortly below.
Take the -grading of to be that by the complexification of a squared real volume form on So there exist -gradings of We include in the object supported on with the trivial rank-one local system and the trivial bounding cochain.
We give the standard Liouville -form so that is an exact Lagrangian. We fix a point and include in the object supported on the fibre with the trivial relative spin structure, an arbitrary -grading, the trivial rank-one local system and the trivial bounding cochain.
Here is the natural relative spin structure on :
Lemma 2.78.
The Lagrangian has a relative spin structure such that the representation on is trivial. If has characteristic this representation is automatically trivial.
Proof.
If has characteristic then defines an object of the exact Fukaya category over and the representation on makes sense over But this is automatically trivial because has only one unit.
Suppose now that has characteristic and choose at first any relative spin structure on . Then defines an object of the exact Fukaya category over (it is only here in the present paper that we use Fukaya categories over a ring). The object defines a representation and such representations are classified by elements of Since the set of relative spin structures is a torsor over the same group we can then choose a possibly different relative spin structure which makes this representation trivial. ∎
We compute the representations corresponding to perturbations of the zero-section
Lemma 2.79.
Let be a closed -form on whose graph lies in the sub-bundle Let be a filtered local system over and regard this as an object of which makes sense because and no non-constant pseudo-holomorphic curve exists for Denote by the one-dimensional representation Regard as the representation of Then is isomorphic to the representation on corresponding to the operator under the isomorphism
Proof.
Let be any element. Then there exists a corresponding element of under the isomorphism Choose a Hamiltonian diffeomorphism such that intersects transversely at in the zero-section and there exists an element of corresponding to
Denote by the fibrewise translation by Put where is the projection This induces a category isomorphism and accordingly a commutative diagram
where we have written for the morphism spaces of Write for the morphism spaces of Since has no differential it follows then that the operator
agrees with the parallel transport map of The areas of the -holomorphic discs are equal to where the integrals are as in Figure 2.2.
The first integral is equal to The second integral vanishes. The third integral is equal to Thus the map
maps to So acts by upon As the functor is an isomorphism, has the same action upon which completes the proof. ∎
We prove another lemma we shall need to prove Theorem 1.3 (ii). For denote by the infinimum of the length of a piecewise smooth loop with and whose homology class modulo torsion agrees with If then may be represented by some geodesic with of length This may be proved by the standard method in Riemannian geometry, explained by Sakai [50, Chapter V, Lemma 1.5] for instance (although he deals with free homotopy classes in place of homology classes modulo torsion, this does not affect the proof). We call the least length of We say that a geodesic is based if For denote by the image of the natural projection and by the least length of
Lemma 2.80.
Let be an object of and put a finite-dimensional -vector space. Denote by the -linear representation defined by the operator and the isomorphism Let be a subgroup, a one-dimensional subspace and a sub-representation of Denote by the composite of group homomorphisms, the last being the original valuation of the Novikov field Then for every with we have
Remark 2.81.
We can also define the least length in the based homotopy class itself, without passing to the homology class modulo torsion, and the same estimate will still hold. But this is weaker than the estimate above because there are fewer loops contained in We shall need the stronger estimate for Proposition 3.6 below.
Proof of Lemma 2.80.
Define the wrapped Floer group by choosing a countable set such that (i) every based geodesic may be lifted to a Hamilonian chord for some and (ii) there exists a constant such that for every Hamiltonian chord for we have Part (i) is possible by the following fact: denote by the squared distance function from the zero-section and let be smooth with derivative then Hamiltonian chords of are the lifts of based geodesics in Part (ii) is possible as in the symplectic cohomology case treated by Venkatesh [64, §2.2].
Take any with Choose a geodesic loop based at of length and with Abusing notation denote by the based homotopy class of Using the -vector space isomorphism we see then that Choose now some whose Hamiltonian flow contains the lift of Recall that and that is a filtered homomorphism. Then and
Take any Then so we can apply (2.74) to in place of which implies
where the first equality follows since is a group homomorphism and the last from the definition of Dividing (2.75) by and letting tend to infinity we get Replacing by we get the same inequality with in place of Consequently ∎
Remark 2.82.
The proof above of Lemma 2.80 implies also
Corollary 2.83.
In the circumstances of Lemma 2.80 suppose that the underlying Lagrangian of the object lies within distance from the zero-section Then for every
Proof.
Consider the fibrewise rescale of by The pull-back by this of the canonical -form on is and the estimate above follows from the proof above. ∎
2.11 The Split-Generation Theorem
Let be a strict category with units and full subcategories. We say that is split-generated by if the image of under lies in the full subcategory split-generated by Here is defined for instance by Seidel [53, Lemma 4.7]. The full subcategory split-generated by is defined by Seidel [53, after Lemma 4.8].
Take , with respect to a fixed background class on that we omit from the notation, and denote by the full subcategory consisting of the single object We prove
Theorem 2.84.
The Fukaya category is split-generated by
We begin by studying the Viterbo restriction functors. We work in a slightly more general context than above. Let be a Liouville domain and a Liouville subdomain (i.e. a codimension submanifold of , with smooth contact boundary). We assume that we have fixed a background class in for the construction of Fukaya categories on , and use the restriction of this class to when discussing Fukaya categories on . We shall assume that all objects of are supported on Lagrangians which are equipped with a primitive function for the Liouville -form, which is constant along the intersection with . In this setting, Abouzaid and Seidel [8] construct a Viterbo restriction functor
On the other hand, the maximum principle for pseudoholomorphic discs in bounded by Lagrangians in yields an functor
on compact Lagrangians. The compatibility between these two functors is given by the following result:
Proposition 2.85.
The left and right modules over associated to the image of each object of in are equivalent to the pullbacks of the associated modules over under the Viterbo restriction map.
Proof.
We make a little modification of Abouzaid–Seidel’s techniques [8]. We begin by constructing a Hamiltonian such that for each exact Lagrangian whose primitive vanishes away from , the generators of the Floer complex which lie away from have strictly positive action. In [8] one constructs a family depending on a parameter by rescaling along the Liouville flow inside , yielding a function that does not depend on away from ; this has the property that the action of all generators in are scaled by , while those outside have constant action. We consider instead a family parametrised by which is constant in , and obtained by rescaling outside; now the action in is constant, but the action outside is scaled.
Consider now a closed (unobstructed) Lagrangian in ; we shall omit all notation for the choices of background classes, local systems, and bounding cochains because they are not relevant to our argument, and discuss only the left module structure, since the argument for the right module structure is exactly the same. The goal is to show that the module structure of over the wrapped Floer cochains of in is equivalent to the module structure obtained by pullback from the corresponding module structure after restriction to . This amounts to constructing a homotopy between the given module structure and one for which we can prove the vanishing of any module operation with input including one chord with boundary on lying away from .
The essential fact that we shall use is that, since is exact, Stokes’s theorem implies that the topological energy of each holomorphic polygon contributing to the module operations is the sum of the negative of the action of the inputs chords with endpoints on , the difference between the values of the primitive of chosen on at the two intersection points with corresponding to the input and output, and the integral along the boundary component mapping to . Since is Lagrangian, the restriction of to is closed, so this last integral depends only on the homotopy class of the path.
We now consider the homotopy of -modules associated to the above family of Hamiltonians indexed by a finite interval . We claim that this defines a homotopy between the initial module structure, and one with the property that the action for a given set of inputs, one of which lies outside of , has valuation which is bounded by a quantity which goes to in the limit . The argument is exactly the same as the proof of [8, Lemma 7.4]: all of these terms in the expression for the topological energy of a polygon are independent of , except those corresponding to inputs lying outside ; since these are rescaled by and are negative, we find that the topological energy of a potential polygon is negative for sufficiently large , hence that the moduli space is empty.
Essentially the same argument as in [8] then implies that the homotopy associated to the entire path is well-defined: even though the indexing family is non-compact, for any given collection of inputs, the contribution with a fixed valuation will be supported in a finite interval, which implies convergence by the completeness of the Novikov field. ∎
We shall use the following consequence:
Corollary 2.86.
If is an object of whose image in lies in the category split-generated by a collection of exact Lagrangians, then lies in the subcategory of split-generated by their images under the restriction functor.
Proof.
Let and denote the left and right modules over the exact category associated to . The criterion for split-generation is that the map from the tensor product of these modules to the self-Floer cohomology of
hits the unit on cohomology. By the above result, these modules are isomorphic to the pullbacks of the modules and over the image of (which we call associated to ). The proof amounts to the homotopy commutativity of the following triangle
which follows by the same argument as the above proposition: the map in Equation (2.78) is obtained by counting discs with an arbitrary number of inputs that are chords with boundary on , and two of which are intersection points of and . The output is a marked point along the interval mapping to . As before, the integrated maximum principle implies that the image of the discs whose inputs are chords that are contained in cannot escape this region, so that the goal is to show that the family associated to rescaling the Hamiltonian outside of by a parameter has the property that there are no discs with prescribed inputs and homotopy class of the path along whenever is sufficiently large. ∎
We return now to the case where is a sufficiently small disc sub-bundle in In [24] Fukaya, Seidel and Smith construct a Lefschetz fibration with total space a Liouville subdomain of (which they call ), whose critical points are of either disjoint from or correspond to the critical points of a Morse function on it (see the discussion preceding [24, Lemma 7]). Taking the inverse image of a small region in the base of the Lefschetz fibration produces a subdomain of , containing the -section, and which has the property that all its Lefschetz thimbles intersect the -section transversely at a point, i.e. are isotopic to cotangent fibres. Combining this with Seidel’s doubling construction [53, §18] we get
a Liouville manifold (which Seidel denotes by ) containing as a Liouville subdomain and
a collection of Lagrangian spheres (which Seidel denotes by ), whose intersection with are Lefschetz thimbles associated to a Morse function on , and with the property that there is a composition of Dehn twists about such that is disjoint from (up to Hamiltonian isotopy); see the first paragraph in the proof of [53, Lemma 18.15].
Since we are interested in studying Fukaya categories that are possibly twisted by background classes, we need to know that any background class on extends to a background class on :
Lemma 2.87.
The restriction map admits a splitting which vanishes on every sphere.
Proof.
Seidel’s construction exhibits as a branched double cover of a Liouville domain containing as a deformation retract. A splitting is then provided by the double covering map. ∎
From these properties we get:
Lemma 2.88.
If is a sufficiently small Weinstein neighbourhood of , then
the intersection of each Lagrangian with is Hamiltonian isotopic to a cotangent fibre, and
the image of every closed Lagrangian in under is disjoint from (up to Hamiltonian isotopy).∎
Fix a background class on and the class induced in by the splitting above. Introduce the Fukaya category which contains the Lagrangians with the trivial relative spin structures (which exists because every is the double of a disc in ), the trivial rank-one local systems and the trivial bounding cochains.
Remark 2.89.
The key result about this Fukaya category is:
Lemma 2.90.
Every object of supported in lies in the category split-generated by the spheres .
Proof.
The proof is the same as that for exact Lagrangians by Seidel [53, Proposition 18.17], using Oh’s result [46, Section 9] for Seidel sequences in the nonexact case. Oh proves indeed that given two objects and a Lagrangian sphere , with corresponding Dehn twist , we have a null homotopy for the composition
where the first map is given by composition and the second by multiplication with a distinguished closed morphism in . So the cone of this morphism lies in the category generated by .
Using this sequence repeatedly we see that if is an object of supported in then there is some closed morphism from to whose cone lies in the category generated by But the Lagrangians underlying and may be made disjoint after a Hamiltonian isotopy, so is exact and the cone of is isomorphic to a direct sum of and (with shifts). This implies that is a summand of the object lying in the category generated by . ∎
We are ready to prove the split-generation result:
Proof of Theorem 2.84.
Lemma 2.87 implies that any background class on is the restriction of a background class on . If , the Lagrangian spheres are automatically relatively spin. In the remaining case, we note that the only possible background class is represented by the Poincaré dual of the zero section, which corresponds to a cotangent fibre; and its pullback to is represented by the inverse image of a cotangent fibre, whose intersection with vanishes with coefficients because the submanifolds intersect an even number of times. The result then follows by applying Lemma 2.90 and Corrolary 2.86, using as well the fact that all cotangent fibres are equivalent in the wrapped Fukaya category, so that one suffices. ∎
2.12 Yoneda Functors
We recall now some standard facts about about Yoneda functors. We follow Seidel [53] with minor modifications. For a strict category with units denote by the category of left modules (Seidel [53, (2f)] works with right modules in place of left modules, and with cohomology units in place of strict units). Seidel [53, (1l) and (2g)] defines the Yoneda functor which is a strict functor, and proves that this induces a fully faithful functor between the cohomology categories [53, Corollary 2.13].
Seidel [53, (2.13)] proves also that if are strict categories and an functor then the pull-back functor and the Yoneda functors fit into the diagram
which commute after passing to the cohomology categories.
Lemma 2.91.
Let be a strict category with units. Let be full subcategories such that is split-generated by Consider the composite of the inclusion functor the Yoneda functor and the pull-back functor Then the induced functor is fully faithful.
Proof.
We repeat the argument by Seidel [53, Lemma 4.7]. Denote by the full subcategory split-generated by Then the composite functor has image in Using (2.81) with we get a commutative diagram
Here is induced from the Yoneda functor and accordingly fully faithful. As (2.82) commutes the composite factoring through and is fully faithful. But is split-generated by so is fully faithful too. The arrow is induced from the Yoneda functor and therefore fully faithful too. The composite is thus fully faithful. ∎
We return now to the circumstances of Theorem 2.84 with and the full subcategory made of the single object Lemma 2.91 implies then
Corollary 2.92.
The induced functor is fully faithful.
The rest of the section is devoted to showing that this functor is compatible in a certain sense with passing to finite covers of Denote by the endomorphism algebra of the single object Then is the category of modules over Write to save notation. The Yoneda functor is then an functor Suppose now that is a finite cover and any point. We define then a corresponding functor and compare this with that on
We denote by the same the induced map between the cotangent bundles, which is also a finite cover. We pull back to the branes on More precisely, the pull-back of an underlying Lagrangian in is possibly disconnected, in which case we include all the connected components of it; for instance, we include the fibre We pull back to the relative spin structure on with which we introduce the Fukaya category
We define then an functor which acts as the set-theoretic push-forward upon the underlying Lagrangians. Choose a compatible almost complex structure on and pull this back to Fix for a moment and recall that are made from gapped categories. We define a gapped functor between these which respects all the operators, simply by pushing forward the pseudo-holomorphic discs. We define in the same way an inverse system of homotopy equivalences with which we can pass to the limits. Then we get a gapped functor between the curved categories Including the bounding cochains we get an functor that we want.
We pass now to the wrapped categories. By definition there exist a localization functor and accordingly a composite functor We pull back to the set of Hamitlonians used to define with which we define the wrapped category We show that the functor descends to Let be the time-one diffeomorphism of a Hamiltonian we have chosen, which is the pull-back of some Hamiltonian Denote by the same the time-one diffeomorphism Then the push-forward of is equal to and the push-forward of the continuation morphism is the continuation morphism The latter is already invertible in and hence we get by the universal property the functor that we want.
We define now a pull-back functor between the Fukaya categories whose underlying Lagrangians are compact without boundary. Take with a Lagrangian, a local system and a bounding cochain. We define a gapped homomorphism
that assigns to each singular chain in the sum of all possible inverse images in Using the fact that discs are simply connected, we see that the moduli space of discs in is a cover of the moduli space of discs in , and hence that a choice of virtual chains for determines the corresponding choice for by passing to the cover. This implies that (2.83) is compatible with the gapped algebra structures on the singular chain complexes The same argument shows that the homotopies required to construct the inverse system are also compatible with the pull-back by So we can take the limit with respect to and (2.83) is a gapped homomorphism. This maps the bounding cochain to a bounding cochain in which defines an functor that we want.
Note that for every Lagrangian we have Hence we get a commutative diagram
On the other hand, there are a pull-back functor and the same functor for in place of We use also (2.81) with From these we get a digram
which commutes after passing to the cohomology categories, where denotes the pull-back by Combining this with (2.84) we get a diagram
which commutes after passing to the cohomology categories; note that we have got rid of the horizontal arrow in the left direction so that the horizontal arrows of (2.86) are both in the right direction. We have thus proved
Proposition 2.93.
For every finite cover the pull-back functor the push-forward algebra homomorphism and the Yoneda functors fit into the diagram (2.86) which commutes after passing to the cohomology categories. ∎
3 Proof of Theorem 1.3
Let be a compact connected orientable Riemannian manifold without boundary. We introduce a word we use in what follows.
Definition 3.1.
A standard Calabi–Yau structure near the zero-section is a pair consisting of a compatible (integrable) complex structure and a -holomorphic volume form relative to which is a special Lagrangian.
We begin in §3.1 with the proof of Theorem 1.3 (i). In §3.2 we prove the key lemma for Theorem 1.3 (ii). In §3.3 we prove Theorem 1.3 (ii) when is abelian. In §3.4 we prove a few facts we shall need to deal with finite covers of In §3.5 we prove Theorem 1.3 (ii) in the general case.
3.1 Proof of Theorem 1.3 (i)
Let be a Weinstein neighbourhood of and regard this as embedded in with a standard Calabi–Yau structure. Suppose first that is trivial.
Proposition 3.2.
If the fundamental group is trivial then the Fukaya category is generated by the single object the zero-section.
Proof.
Let be an object of and consider its image under the Yoneda functor Applying the homological transfer lemma, we may replace by an algebra whose underlying graded vector space is given by its cohomology, which is isomorphic to and hence is supported in non-positive degrees. It follows (by the degree-filtration argument [5, Appendix A]) that the Yoneda image of is a repeated extension of ungraded modules over the degree algebra But is isomorphic to so the image of is a repeated extension of ungraded -vector spaces. As is supported on a compact Lagrangian without boundary, the number of these extensions is finite and each ungraded piece is finite-dimensional over Recall now that intersects transversely and exactly at one point and that is one-dimensional over Then each ungraded piece of the Yoneda image of may be written as for some finite-dimensional -vector space As the functor is fully faithful, the object is isomorphic in the cohomology Fukaya category to a repeated extension of objects of the form This completes the proof. ∎
We use an algebraic lemma by the first author [5, Lemma A.4], which goes back to the first general results on Arnold’s nearby Lagrangian problem [25]. The original statement is over but the proof shows that it applies also to :
Lemma 3.3.
Let be a -graded -linear triangulated category generated by one object whose endomorphism algebra is supported in non-negative degrees. Then every object with supported in non-negative degrees is isomorphic in to for some integer and for some finite-dimensional -vector space ∎
Let be supported on a closed special Lagrangian relative to a standard Calabi–Yau structure. Then is a non-zero object and by Proposition 2.55 the graded group is supported in non-negative degrees. Proposition 3.2 and Lemma 3.3 imply then that is isomorphic in to some with the trivial bounding cochain. Since is non-zero it follows that is a non-zero vector space. Note also that there are vector space isomorphisms
which is non-zero (by Corollary 2.64 or because is isomorphic to the ordinary cohomology group). Hence it follows by Proposition 2.67 (ii) that
If is not trivial but finite, taking the universal cover we find so ∎
3.2 The Key Lemma
Put Fix a point and fix a grading of by phase at Then the following holds, for which we do not need any hypothesis on :
Lemma 3.4.
For every standard Calabi–Yau structure near there exists a neighbourhood of such that for every supported on a closed special Lagrangian graded by the graded vector space is supported in degree
Remark 3.5.
Here the speciality is essential because otherwise may be supported in two or more degrees; see Example 5.10.
There is in fact an analytical theorem by Almgren [10], explained also by De Lellis and Spadaro [14, Theorem 2.4]. Suppose that the nearby special Lagrangians are in a given homology class. Then they are close to the zero-section as varifolds in the sense of geometric measure theory. Then there exists a small set in such that if is outside it then near each special Lagrangian is close to the zero-section (the same arguments apply also to area-minimizing currents). This implies that the Maslov index of the pair is zero, so the result holds at the chain level. But we do not want to fix the homology class of and the cohomology level result will do for us.
Proof of Lemma 3.4.
Suppose first that the Calabi–Yau structure near is real analytic. For each we define a compactly-supported Hamiltonian perturbation of which is special Lagrangian near To start with, take a -linear isomorphism which maps to to the symplectic form to and the holomorphic volume form to where are the coordinates of Take and with so that the plane is special Lagrangian with respect to the flat Calabi–Yau structure. Take a small ball in about and embed it into near mapping to Denote the image of the ball by the same for short. Take a neighbourhood of in over which is a product bundle. We can suppose that the ball lies in the product bundle Denote by the vertical projection of the product bundle. Put which is a ball in about the origin. Choose a neighbourhood of which is so small that
We perturb into a special Lagrangian with respect to the Calabi–Yau structure of There are two ways of doing this. The first is to use Cartan–Kähler’s theorem as Harvey and Lawson do [28, Chapter III, Proof of Theorem 5.5], which we can do because we have supposed that our Calabi–Yau structure is real analytic. The second is to use the implicit function theorem again as Harvey and Lawson do [28, Chapter III, Proof of Corollary 2.14]. In either case we get a special Lagrangian near We can suppose that this is a graph over Using a cut-off function we get then a compactly-supported Hamiltonian perturbation of such that is a special Lagrangian in
Grade the Lagrangians and by phase and the Lagrangians and by phase at For define an affine function by For define an affine function by Using these we get a Hamiltonian isotopy from to the plane which does not cross As there is a graded Hamiltonian isotopy from to This induces a -graded Hamiltonian isotopy from to Suppose now that is supported on a closed special Lagrangian Since it follows then that the special Lagrangian part of does not agree with We can thus apply Proposition 2.69 to and consequently is supported in degrees So is supported in degrees Since this holds for every it follows that as we want.
Finally, if the Kähler form near is not real analytic we can approximate it by a real analytic one (after taking a Kähler potential) and find a compactly-supported Hamiltonian diffeomorphism such that is real analytic. Applying to this the result above, we get a corresponding neighbourhood of Put Let be supported on a closed special Lagrangian (relative to ) and denote by the push-forward. Then is supported in the single degree But as is compactly supported, we have which completes the proof. ∎
The effect of the lemma above is that we do not have to deal with the extension problem which we explain now. Let be an object of As in the proof of Theorem 1.3 (i) the Yoneda module is a repeated extension of ungraded modules over that is, representations of These extensions may be nontrivial. In Example 5.10, for instance, the direct sum of and the graph of are different from their surgery. More precisely, these define non-isomorphic objects of They are made up of the same representations but with different extensions.
3.3 The abelian Case
We prove Theorem 1.3 (ii) when is abelian. We work in a Weinstein neighbourhood of which we embed in Denote by the metric on induced from and by the standard Calabi–Yau structure near induced from Define a function by Denote by the linear elliptic operator where is computed with respect to This is the linearized operator we use to prove McLean’s theorem [34, Proposition 2.13]. By a version of Hodge theory the de Rham cohomology group is isomorphic to the vector subspace By the de Rham theorem there exists such that the following holds: suppose given and a group homomorphism with for then may be represented by some closed -form on with and where the pointwise norm is computed with respect to By McLean’s theorem [34, 43] there exists such that the following holds: let be a closed -form on with and then the graph of is Hamiltonian equivalent to some special Lagrangian near with respect to Denote by the disc sub-bundle of radius which is so small too that we can use Lemma 3.4.
Let be a closed -unobstructed special Lagrangian and let be supported on Corollary 2.56 implies then that is a non-zero object. By Lemma 3.4 the graded vector space is supported in a single degree. Denote by the representation made from the products in By hypothesis the group is abelian so has a one-dimensional sub-representation Recall that and denote by the closed -form on which represents the de Rham class and satisfies the equation Using the notation of Lemma 2.79 write where is some group homomorphism This defines a rank-one local system over As in Remark 2.11 the category of rank-one free modules is equivalent to that of one-dimensional valued vector spaces. Hence we get a filtered local system over which we denote by the same Lemma 2.79 implies then that is isomorphic to the representation on As is a sub-representation of there exists a degree-zero non-zero morphism from to Thus
3.4 Finite Covers and McLean’s Theorem
Let be a compact Riemannian manifold without boundary, and a point. Recall that for we have defined before stating Lemma 2.80. Let be a finite cover and denote by the homomorphism induced by We use the -vector space isomorphism
Proposition 3.6.
For every integer there exists independent of the finite cover and so large that every linear map with may be represented by a closed -form with norm with respect to the induced metric
Proof.
The proof contains several key steps. We begin by choosing some geodesics in and a triangulation of We prove in Lemma 3.7 a uniform estimate for the geodesics; uniform means that the relevant constants are independent of the finite cover We choose then a subgroup of -cycles which generate We prove in Lemma 3.8 a uniform estimate for the elements of We pass then to the -coefficients and introduce a -cocycle of cohomology class We prove in Corollary 3.9 and Lemma 3.10 that satisfies a uniform estimate. We use the de Rham theorem to define a closed -form corresponding to which will satisfy the uniform estimate we want.
We turn now to the formal treatment. Recall that the fundamental group of the compact manifold is generated by some finitely many elements For with represent by a geodesic with and of length For with take to be the constant loop with value
Note that intersect one another only at finitely many points. Choose then a smooth triangulation of which contains as a sub-complex and is so small that every simplex is contained in some geodesic ball of the Riemannian manifold The latter condition implies that for distinct with the simplices containing are disjoint (for if there were some intersection point then the geodesic between and that between would be two different lifts of the geodesic between but the lift ought to be unique). The finite cover induces then a triangulation of which we use in what follows.
Put and denote by the lifts under of Denote by the length of each Regard as -chains in the free abelian group Denote by the norm on that is, if we denote by the set of edges (or -simplicies) in then for we have We prove
Lemma 3.7.
There exists independent of and so large that
Proof.
Put which is independent of Denote by the greatest length of edges in and by the least length of edges in which are also independent of Fix and denote by the number of edges contained in then On the other hand, the norm of is the multiple count of the relevant edges so
This proves the lemma with ∎
Denote by the subgroup of -cycles. We show that the natural projection is surjective. Let be any element and represent this by some loop based at Then is a loop in based at which is based homotopic to some concatenation of Lifting this to we see that is based homotopic to some concatenation of The corresponding -cycle in has then homology class which proves our claim.
As is a free abelian group, we can choose a subgroup such that the natural projection is surjective. For denote by its homology class modulo torsion.
Lemma 3.8.
For we have
Proof.
Write with Recall that each agrees with some and that the least length of the homology class is equal to the length of the paths So
Lemma 3.7 implies then
because have no common edges and are orthonormal. This completes the proof. ∎
We pass now to the -coefficients. Extend the natural projection to the -linear map Extend also the given homomorphism to Define an -linear map to be the composite of the natural projection and the extended We prove
Corollary 3.9.
For we have
Proof.
The hypothesis on and Lemma 3.8 imply that for This extends readily to because we can cancel out the denominators. Note that is an -linear map between finite-dimensional vector spaces, which is automatically continuous with respect to the norm topologies. Moreover is dense in so the estimate holds with -coefficients as we want. ∎
We introduce a -cocycle of cohomology class Denote by the subspace of -boundaries and by its orthogonal complement. Put and denote by the orthogonal projection. Define an -linear map to be the composite of this and that defined above. Then vanishes on and is accordingly a -cocycle on Since agrees with the orthogonal complement in to it follows that for if we denote by its projection onto then and in particular, so has cohomology class We prove
Lemma 3.10.
For every edge we have
Proof.
We use now the -linear map from the -vector space of -cochains to that of -forms, defined by Singer and Thorpe [57, §6.2, Lemma 1], which maps -cocycles to closed -forms and induces the de Rham isomorphism of the two cohomology groups. Part (4) of their lemma shows that for each edge in the corresponding -form is supported near in The process of assigning the -form is thus local, which we can do essentially in before passing to the finite cover Hence it follows by Lemma 3.10 that the -form corresponding to is bounded uniformly with respect to completing the proof. ∎
We prove also that McLean’s theorem holds uniformly with respect to finite covers.
Proposition 3.11.
Let be a compact Riemannian manifold without boundary, the disc sub-bundle of radius and a standard Calabi–Yau structure on Define a function by Then there exists such that the following holds: let be a finite cover and a closed -form on with and both computed with respect to the induced metric then there exists a function such that is a special Lagrangian section of
Proof.
For an integer denote by the Sobolev space of -forms on and by that of -forms on Denote by the co-dimension one closed linear subspace consisting of with Denote by the linear elliptic operator We regard as a multiplication operator and write for instance. By the definition of the operator is surjective so there exists a right inverse to it, which we denote by The key fact is that we can choose Sobolev norms on these spaces such that is bounded uniformly with respect to the finite cover
Those Sobolev norms we get directly from the pull-back metric will not do for this purpose. For instance, take with using the co-ordinate on Let be an integer and the -fold cover modulo for Then which contributes to the Laplacian and to the Green operator by The latter is clearly unbounded, which we shall not want to happen.
We define now the Sobolev norms we use. Let be less than half the injectivity radius of the Riemannian manifold For denote by the open geodesic ball of radius about As is compact there exists a finite set such that We look now at the finite cover For with denote by the lift of with respect to which exists because is simply connected. Thus is an open geodesic ball in Define then a norm function by
where the latter is the ordinary Sobolev norm; that is, where are computed with respect to We prove now that is bounded below in an -uniform way.
Lemma 3.12.
There exists independent of the finite cover and so large that for every we have
Proof.
Recall that if is a compact Riemannian manifold without boundary and a smooth function then there exists such that for every with we have the constant is the inverse of the first non-zero eigenvalue of the elliptic operator This extends to a compact Riemannian manifold with boundary, because the estimate follows from that on the double (with a modified metric). Suppose therefore that may now have boundary and note that for every with we have
Thus Given an element of not satisfying the vanishing mean condition, add to this a constant and use the result we have just obtained; then we get the same estimate for it. Consequently, for each there exists independent of and so large that for every we have Take any and for put Then
where This completes the proof. ∎
Fix now so that we have the Sobolev embedding We prove that the Sobolev embedding and the elliptic regularity are -uniform too.
Lemma 3.13.
There exists independent of the finite cover and so large that for every we have
Proof.
By the Sobolev embedding on there exists independent of and so large that for every and we have Then for we have
which proves the left inequality of (3.5).
We turn now to the other part. We show first that for with the non-empty intersection is necessarily connected. If there were two components, there would be two distinct geodesics connecting these two. But have been chosen to have the common radius which is less than half the injectivity radius. So lies in a single geodesic ball, which contradicts that the two geodesic exists. Thus is connected. The union is accordingly simply connected.
As is covered by ’s of radius there exists so close to that if we denote by the concentric ball of radius then still covers Then for we have
For and denote by the set of such that and the simply connected set lifts to Exchanging the order of the two sums in (3.6) we see then that
Note that there is an injection defined by so that the number is bounded above by some -independent constant It follows then from (3.7) that
On the other hand, by the ordinary a priori estimates over there exists independent of and so large that for every and we have
Hence it follows by (3.8) that
Thus will do for the right part of (3.5). ∎
Corollary 3.14.
There exists independent of the finite cover and so large that for every we have
Proof.
We complete now the proof of Proposition 3.11. Let be so small that if is a section of with then is a section over of Note that if is a section of with then we can define the pull-back by and in particular the -form Put The left inequality of (3.5) implies then that for every we can define as a section over of By the proof of McLean’s theorem [43, 34] there exists a non-linear operator defined by We solve the equation Standard computation shows that is differentiable at with respect to the Sobolev norms, whose differential agrees with Corollary 3.14 implies then that for with we have so We solve the equation for with Define by We seek then a fixed point of the composite and use therefore the contraction mapping principle. As is the quadratic or higher term made up of there exists independent of and so large that the following holds: let satisfy at every point of and put then
at every point of Integrating this and using the Cauchy–Schwarz inequality we get
Putting and we have
This with (3.9) implies
Hence choosing and using Corollary 3.14 we see that is a contraction map from the closed ball to itself. So there exists with Recall that has been any integer and that fixed points are unique; then we see that for every Thus is and accordingly so is Now so The graph of is the special Lagrangian we want, which proves Proposition 3.11. ∎
Here is a corollary of Proposition 3.6 and the estimates in the proof above.
Corollary 3.15.
There exists independent of the finite cover and so large that every linear map with may be represented by a closed -form on with norm and where the norm and are computed with respect to the induced metric
Proof.
Fix an integer Then by Proposition 3.6 there exist a constant independent of and a closed smooth -form on with Define a -form on by Then and We estimate the norm of By Corollary 3.14 there exist constants independent of and so large that
By Lemma 3.13 there exists a constant independent of and so large that and
Thus which completes the proof. ∎
3.5 Proof of Theorem 1.3 (ii)
We prove Theorem 1.3 in the general case (when is not abelian). We begin by choosing a disc sub-bundle We impose two conditions. Firstly, should be so small that we can use Lemma 3.4. Secondly, let be as in Corollary 3.15 and as in Proposition 3.11; then the radius of should be
Let be a closed -unobstructed special Lagrangian and an object supported on Then is a non-zero object. By Lemma 3.4 the graded vector space is supported in a single degree. Denote by the representation made from the products in By hypothesis the group has no non-abelian free subgroup, and accordingly its image does not have such a subgroup either. Since is finitely generated it follows that so is To this we apply the following two results: one is by Tits [60, Corollary 1] and the other is a Lie–Kolchin theorem stated for instance by Borel [12, Chapter III 10.5]. Both apply to fields of arbitrary characteristic. For the latter statement we need the field to be algebraically closed. We therefore pass to the algebraic closure of as in §2.1. This is possible because the bounding cochain equations will still have solutions with the larger field. We denote by the same the algebraic closure.
Theorem 3.16 (Tits).
Every finitely-generated linear group has either a nonabelian free subgroup or a finite-index solvable subgroup.
Theorem 3.17 (Lie–Kolchin).
Every solvable linear group over an algebraically closed field has a finite-index triangularisable subgroup.
Consequently there exists a finite-index subgroup of on which is triangularisable. Let be the corresponding finite cover, and take a one-dimensional sub-representation of the triangularisable group.
Recall that and denote by the closed -form on that represents the cohomology class and satisfies the equation Denote by the graph of Using the notation of Lemma 2.79 write where is some group homomorphism This defines a rank-one local system over Hence we get a filtered local system over which we denote by the same Lemma 2.79 implies then that is isomorphic to the representation on Denote by the restriction to of As is a sub-representation of there exists a degree-zero non-zero morphism from to Hence it follows by Proposition 2.93 that there exists a degree-zero non-zero morphism from to Thus
By Corollary 2.83 we have for Corollary 3.15 implies then that may be represented by a closed -form on with and Proposition 3.11 implies then that the graph of is Hamiltonian equivalent to some special Lagrangian As this Hamiltonian perturbation does not change the isomorphism class of it follows from Proposition 2.67 (i) that is an irreducible component of Thus is unbranched, which completes the proof of Theorem 1.3 (ii). ∎
Remark 3.18.
Theorem 1.3 holds for some badly immersed special Lagrangians. In the known construction of Fukaya categories, the underlying Lagrangians are required at least to be cleanly immersed [20]. On the other hand, in our nearby problem we want to include as many (special) Lagrangians as possible, so we introduce the following notion [33, before Corollary 4.6].
Let be a closed manifolds and be a Lagrangian immersion. Put We say that an object in is supported near if there exist arbitrarily-small Lagrangian perturbations of which underlie some objects in the same isomorphism class in More concretely, this means that for every there exists a perturbation of the immersion with norm bounded by for all , and which supports an object representing the given isomorphism class
Theorem 1.3 holds for a special Lagrangian which admits such an isomorphism class as above. The reason is that the results of the paper [33] are stated and proved for objects of supported near a special Lagrangian. In particular, Proposition 2.67 and Lemma 3.4 are still valid for these objects. So the proof of Theorem 1.3 will work.
But the condition above on is rather strong. Here is one possible way of verifying it. Suppose given a smooth family such that and the other are all generic and mutually isotopic under domain Hamiltonians on in the sense of Akaho–Joyce [9, Definition 13.14 (ii)] (this is a family version of the classical transversality condition in differential topology). If we could lift such an isotopy to an ambient Hamiltonian isotopy, then it would follow that any isomorphism class in supported on the immersion is also supported on for all generalising the Hamiltonian invariance theoreom by Fukaya, Oh, Ohta and Ono [21]. In general, such a lift does not exist because then those strips which bound arcs between self-intersection points will have constant area under the isotopy. We can perhaps use the methods of [48] to prove invariance statements for this perturbation problem.
4 Tsai–Wang’s Method
We begin by recalling the key formula from geometric measure theory. Let be a Riemannian manifold. Given a vector field on , a point , and an -dimensional plane we denote by the trace of the map where the two arrows are the covariant derivative and the orthogonal projection In particular, if is the gradient vector field of a smooth function then We use also the notion of rectifiable integral varifolds in which we treat, following Simon [56], as Radon measures on Recall that if is a compactly supported stationary integral -varifold in then for every vector field on we have In particular,
for every smooth function
We recall now the following result from the literature [14, 61, 62]. We prove this to show how to use (4.1).
Proposition 4.1.
Let be a Ricci-flat Calabi–Yau manifold of complex dimension and a closed embedded special Lagrangian which has positive Ricci curvature with respect to the induced metric from Then there exists a neighbourhood of in which every compactly supported stationary integral -varifold is supported on
Proof.
Tsai and Wang prove that is strongly stable in their sense [62, Proposition A(iii)] and that the following holds [62, Proposition 4.1]: there exists an open neighbourhood of such that for every point and for every -dimensional plane we have at
where is the distance from and depends only upon Applying (4.1) with and using (4.2) we find on the support of the varifold completing the proof. ∎
Remark 4.2.
The geometric meaning of the proof is as follows: let be a tubular neighbourhood and the flow of the gradient vector field of then does not increase the area of
As is well known, closed special Lagrangian currents are area-minimizing and in particular area-stationary [28]. Combining this with Proposition 4.1 we see immediately that the following holds:
Corollary 4.3.
In the circumstances of Proposition 4.1 every compactly supported closed special Lagrangian current in is necessarily non-singular, of phase and supported on ∎
Example 4.4.
We turn now to the example which shows that abelian is essential to the latter part of Theorem 1.3 (ii). We begin by proving a lemma in slightly more general circumstances.
Lemma 4.5.
Let be integers. Take and give these the flat metrics. Then every compactly supported stationary integral -varifold in is parallel to that is, supported on for some finite set
Proof.
Let be a compactly supported stationary integral -varifold in Applying (4.1) again with we get But is pointwise non-negative and degenerates only in the direction of so -almost everywhere in we have and is parallel to
We recall a lemma of Simon [56, Lemma 19.5]. Let be a point on the support of and we work near this so we can regard as a point of the universal cover of Let be so small that for every point of distance from we can define Fix and let be a point on the support of of distance from and such that is not parallel to Simon [56, Lemma 19.5] proves then that
where denotes the density of and the ball of radius about The original formula has in fact another term on the right-hand side, which contains the integral over of where are the projections of onto respectively. But is parallel to so the additional term vanishes and (4.3) holds.
Remark 4.6.
The hypothesis of Proposition 4.1 is stable under small perturbations of the Ricci-flat Kähler metric (just because positivity is an open condition). On the other hand, Lemma 4.5 will fail under perturbations of the ambient metric. For instance, Kapouleas and Yang [37] construct a sequence of minimal surfaces converging to the flat Clifford torus with multiplicity two; and a suitable rescaling of the ambient metric converges to the flat metric on the tubular neighbourhood of
Corollary 4.7.
Let be a flat Calabi–Yau manifold, a closed embedded totally geodesic special Lagrangian, and a tubular neighbourhood of Then every compactly-supported closed special Lagrangian current in is parallel to that is, a locally constant multi-valued graph over Every such graph is a disjoint union of single-valued graphs if and only if is abelian.
Proof.
Since is a closed flat manifold we get according to Bieberbach a finite Galois cover such that is a maximal abelian subgroup of Denote by the same the finite cover between the cotangent bundles. Identify with a neighbourhood of the zero-section and the pull-back with a neighbourhood of As is a product bundle we can use Lemma 4.5 to show that the pull-back of a compactly supported closed special Lagrangian current is parallel; and accordingly, so is the original Now if is abelian, this is the maximal abelian subgroup of itself and so every special Lagrangian is a disjoint union of single-valued graphs. If is non-abelian, there exist a non-trivial covering transformation and accordingly an element which acts non-trivially upon So there exists a constant -form on that does change under By rescaling we can suppose that is so small that we can define its graph in The push-forward of this defines then a truly multi-valued graph in ∎
Example 4.8.
Let with a Klein bottle. Write as the mapping torus of the automorphism given by modulo for Then there are corresdponding double covers and under which the graph of is pushed forward to a multi-valued graph. Precisely speaking, is not Calabi–Yau and is not even orientable. But is still a special Lagrangian in the extended sense as in Definition 2.32; see also Remark 2.33. The multi-valued graph is also a special Lagrangian in the same sense. It is easy to find multi-valued graphs which are ordinary special Lagrangians. For instance, take to be the mapping torus of the orientation-preserving diffeomorphism given by where then the graph of is pushed forward to a multi-valued graph which is an ordinary special Lagrangian.
5 Branched Examples
Here are examples of branched nearby closed Lagrangians, which will all be Lagrangian surgeries. We begin by recalling the result of Fukaya, Oh, Ohta and Ono [21, Chapter 10] which describes the effect of applying Lagrangian surgeries to pairs of objects in the Fukaya category.
Theorem 5.1 (cf. [21, Chapter 10]).
Let be two closed embedded -graded Lagrangians which intersect each other transversely at least at one point, with the Maslov index function identically equal to Let be a -graded Lagrangian surgery [51, Lemma 2.14]. Then for any supported on respectively, there exist an object supported on and for every other a long exact sequence
where the connecting homomorphism is the right multiplication by some element of ∎
Remark 5.2.
We allow to have two or more intersection points, which is possible as Fukaya, Oh, Ohta and Ono mention [21, after Theorem 55.7] although they deal only with singular intersection points in their formal treatment. So there are many objects of the form above according to different scales of surgeries at different intersection points.
Here are remarks on the proof. Fukaya, Oh, Ohta and Ono use the cobordism method for the moduli spaces of holomorphic discs with boundary on the surgery; the cobordism is associated to stretching the neck (degenerating the almost complex structure) near the intersection points. The method of Biran–Cornea [18] provides an easier proof which does not rely on a degeneration argument: one can associate to the surgery a Lagrangian cobordism in , and the count of holomorphic triangles with boundary on this cobordism yields the desired long exact sequence. Although their results are stated for monotone or exact Lagrangians, these assumptions are imposed only to avoid using the theory of virtual chains in case the Lagrangians bound holomorphic discs. We shall in practice take to be graphical Lagrangians in a cotangent bundle, which do not bound any non-constant holomorphic discs; so their proof can be implemented with no difficulty.
For special Lagrangians we have
Corollary 5.3.
In the circumstances of Theorem 5.1 let be both special Lagrangians. Then is not isomorphic in to any other object supported on a closed special Lagrangian of any phase.
Proof.
By shifting we may suppose that has phase for some and that has phase Then by (2.28) we have so Applying (5.1) to in place of we get two exact sequences
On the other hand and are both nonzero, so and are both nonzero. Suppose contrary to our conclusion that is isomorphic to an object supported on a closed special Lagrangian of phase By perturbing we may suppose that is supported on a nearly-special Lagrangian which intersect and generically. Since we get an intersection point with Maslov index Hence denoting by the phase function and using (2.28) we find and in particular Also since we get an intersection point with and in particular But both and are near so that is, and have all phase Since and are both non-zero it follows by Proposition 2.67 (i) that these three Lagrangians have a common irreducible component. But this is impossible because intersect each other transversely at least at one point. ∎
We recall now a key fact about Morse -forms; a Morse -form on a manifold is a closed -form on such that every point of has a neighbourhood on which for some Morse function
Theorem 5.4 (Honda [32]).
Let be a closed manifold of dimension and a Morse -form on with no critical point of index or Then there exists on a Riemannian metric and in the de Rham class a -harmonic Morse -form which has the same distribution of indices as has; that is, for each the number of index- critical points of is equal to that of ∎
Remark 5.5.
If is equipped with any real analytic structure, we can suppose that and above are real analytic. To see this, perturb the metric from Honda’s theorem to a real analytic Riemannian metric, which we denote still by the same . Choosing the perturbation to be sufficiently small, there exists a unique -harmonic -form near in the same de Rham class, which we denote again by the same This is analytic by elliptic regularity.
Here is another fact we will use. We give the canonical symplectic and Liouville structures.
Lemma 5.6.
Let be a closed oriented real analytic Riemannian manifold. Then there exist a disc sub-bundle and a standard Calabi–Yau structure in the sense of Definition such that the restriction of the Kähler metric to is equal to and the complex structure is compatible with the Liouville structure of in the sense of Definition ∎
Remark 5.7.
The real analytic structure induced from the complex structure above is possibly different from that induced from the real analytic structures of and
Proof of Lemma 5.6.
Guillemin and Stenzel [27, §5] prove this except the part which has to do with that is, there exist a sufficiently small disc sub-bundle and a complex structure on compatible with the Liouville structure of and such that the restriction to of the corresponding Kähler metric is equal to Making smaller if we need, extend to a holomorphic section of which is locally of the form where are local holomorphic co-ordinates with which is defined by This is possible because every is real analytic. Since is oriented we get, making yet smaller if we need, a holomorphic volume form on which is locally of the form and whose restriction to agrees with the volume form of This completes the proof. ∎
Remark 5.8.
Guillemin and Stenzel [27, §5] do not suppose that is oriented. They show also that is invariant under the involution defined by and that is the unique complex structure satisfying this and the other conditions in the lemma.
We give now an example of non-special Lagrangian surgeries. Let be a closed connected orientable manifold of dimension and a Morse -form on whose critical points have the same index. The Poincaré duality of Morse–Novikov homology implies then that must be even and the critical points must have index We prove
Theorem 5.9.
There exist a disc sub-bundle a closed embedded branched -graded Lagrangian and an object of supported on that is not isomorphic to any other object supported on a closed special Lagrangian of any phase relative to any standard Calabi–Yau structure
Proof.
Applying Theorem 5.4 to we get a harmonic -form with the same distributions of critical points as has. Applying McLean’s theorem to we get a nearby special Lagrangian such that the pair has the same distribution of indices as has. Consider now as in Theorem 5.1. Corollary 5.3 implies then that this cannot be represented by a special Lagrangian, completing the proof. ∎
Example 5.10.
Let be an even number and take to be the self-connected sum of that is, is obtained from by eliminating two points of it and attaching the handle between them. Take a Morse function with one critical point of index two of index and one of index Performing the connected sum construction at the critical point of index and that of index we obtain a -form on satisfying the hypothesis of the theorem above. There is accordingly a branched nearby Lagrangian with unobstructed. On the other hand, as there is an isomorphism so the hypotheses of Theorem 1.3 (ii) holds except the special Lagrangian condition.
Let be an object of the Fukaya category supported on and let be any point. Then the Floer cochain group is non-zero in degrees and As these two degrees differ at least by two and no differentials exist between them. So the cohomology group is non-zero in two different degrees, which is a counterexample to Lemma 3.4.
We give now an example of special Lagrangian surgeries.
Theorem 5.11.
Let and let have a Morse -form with at least one critical point of index at least one of index and none of index or Then near there exist a standard Calabi–Yau structure and relative to it a closed branched special Lagrangian.
Proof.
Following the proof of Theorem 5.9 we get a nearby special Lagrangian such that the Maslov indices of the pair has the same distributions of the Morse index of So has at least one pair intersection points, one of index and the other of index which thus satisfies the hypothesis of Joyce [34, Theorem 9.7]. We can define then a special Lagrangian surgery as we want. ∎
The branched special Lagrangian in the preceding theorem is possibly immersed; any pairs of critical points of of indices and may be removed by surgery but the other critical points become transverse double points.
Example 5.12.
Take to be the selfconnected sum of then has a Morse -form with exactly two critical points, one of index and the other of index So with we can apply Theorem 5.11 and if we have so that is -unobstructed.
But there are also -obstructed special Lagrangian surgeries:
Corollary 5.13.
Let be as in Theorem 5.11 and assume that has no nonabelian free subgroup.
If then near there exist a standard Calabi–Yau structure and relative to it a closed embedded -obstructed special Lagrangian.
If then near there exist a standard Calabi–Yau structure and relative to it a closed generically-immersed -obstructed special Lagrangian
Proof.
Example 5.14.
For (i) take and regard it as the selfconnected sum of According to Milnor [44, Lemma 8.2] for every Morse function we can produce, by changing it locally, two additional critical points of any adjacent indices. Applying this to the height function of we get a Morse function with exactly one critical point of index exactly one critical point of index and arbitrarily many pairs of critical points of indices and So has such a Morse -form as in (i) above.
In the same way, on with there is such a Morse -form as in (ii) above with arbitrarily many pairs of critical points of adjacent indices or
Corollary 5.13 implies a statement without Lagrangians in it and perhaps of interest in itself:
Corollary 5.15.
Let be a closed connected orientable manifold of dimension with having no nonabelian free subgroup. Let be a Morse -form on with no critical points of index or Then
if we have and has no critical point at all.
Suppose now that and has at least one critical point of index and one critical point of index Then
if there exists a smooth function such that has no critical point at all; and
if then either or has some critical point of index or
Proof.
We first prove (i). Since has no critical points of index or it follows that is not exact, so If this condition and the hypothesis on imply , but since the Euler characteristic of vanishes, every Morse -form on with no critical point of index or has no critical point at all, as in (i).
We then prove (iii). Since the immersed Lagrangian in Corollary 5.13 (ii) has obstructed, it follows that either or has some self-intersection points of index or So (because ) or has some critical points of index or
We prove (ii) in two different ways. The first proof goes as follows. Let be two critical points of one of index and the other of index such a pair exists by Poincaré duality of Novikov homology. As in Corollary 5.13 (i) there is a double point surgery with obstructed. In dimension the bounding cochains on are critical points of the disc potential function. The relevant homology classes come from those of strips between with the right orientation. For this we use again Fukaya, Oh, Ohta Ono’s result [21, Chapter 10]; the number of holomorphic discs homologous to this strip with multiplicity one is conserved under the surgery. We prove that the number of holomorphic discs homologous to the multiplicity-one strip is exactly one. This is an integer because is an exact symplectic manifold and we can therefore use the result of Fukaya, Oh, Ohta and Ono [23]. We can also work with the Novikov field of any characteristic so the Lagrangian has obstructed in characteristic and in particular, is non-zero modulo But is any, so Denote by the graph of As the Floer complex of the pair and the Novikov complex of are quasi-isomorphic, the number of gradient trajectories from to is also one. We can then use the cancellation theorem in Morse theory [44, Theorem 5.4]; that is, we can cancel out with each other by changing in its de Rham cohomology class. Repeating this process we get a Morse -form with no critical point at all, as claimed.
For the other proof we use the geometrisation theorem on -manifolds, from which we deduce
Proposition 5.16.
Every closed orientable -manifold satisfies one of the following five statements: (i) is finite; (ii) (iii) (iv) there exists a finite cover where is the mapping torus of an orientation-preserving diffeomorphism or (v) has a nonabelian free subgroup.
Proof.
We suppose that is infinite and prove that one of (ii)–(v) holds; we shall refer to the book [11] for those results about -manifolds which we use, without tracing back their origins in the literature. As is infinite, has no Riemannian metric of positive constant sectional curvature, so is not spherical [11, §1.7]. According to Theorem 1.11.1 of [11], either one of (ii)–(iv) above holds; or has no finite-index solvable subgroup. It thus remains to show that, if is an infinite -manifold group with no finite-index solvable subgroup, then it must have a nonabelian free subgroup.
The first step in this direction is to note that [11, p50, (C.6)] asserts that our assumption on the fundamental group implies that one of the following (a)–(d) holds: (a) is reducible [11, p3]; (b) is a hyperbolic manifold and is a linear group over [11, p51, (C.7)]; (c) is a Seifert manifold and is a linear group over [11, p52, (C.10) and (C.11)]; or (d) is a Haken manifold with an incompressible torus and (v) holds [11, p58, (C.24)]. Moreover:
In Case (a) it is well known that either (ii) holds or is a connected sum of two closed -manifolds neither of which is a sphere. In the latter case are both nontrivial (because Poincaré’s conjecture is true). If then and (iii) holds. Otherwise, has so many elements that (v) holds as follows: one of has at least two nontrivial elements and the other of has at least one nontrivial element so there is a nonabelian free group generated by
In Cases (b) and (c) we use Tits’ theorem [60, Theorem 1]: in characteristic every linear group has either a finite-index subgroup or a nonabelian free subgroup (we have used already a positive-characteristic version, Theorem 3.16). Since we have supposed that has no finite-index solvable subgroup, it follows that (v) holds in Cases (b) and (c).
This completes the proof. ∎
Proposition 5.16 implies
Corollary 5.17.
Let be a closed orientable -manifold with having no nonabelian free subgroup. Then every nonzero element of may be represented by a nowhere-vanishing closed -form on
Proof.
By Proposition 5.16 one of (i)–(iv) holds. In Cases (i) and (iii) we have and there is nothing to prove. In Case (ii) we have and letting be the coordinate of we have which has certainly the desired property. In Case (iv) the mapping torus is determined by the monodromy matrix acting upon which we can classify by direct computation into the following three cases; (a) there is an integer such that (b) is conjugate in to for some or (c) has two distinct real eigenvalues, neither equal to We study these respectively as follows:
In Case (a) we can rechoose the finite cover so that from which we get a natural embedding where are the three coordinates of This expression implies the desired property of
In Case (b) we can rechoose the finite cover so that will have two eigenvalues both equal to one. Trivialising over two intervals in and applying to them the Mayer–Vietoris sequence, we get an isomorphism
(5.4) where is obtained from the projection and the right-hand side is the -invariant subspace of As has two eigenvalues both equal to one, there is a coordinate such that is invariant under and lifts to Then by (5.4) we have in which has certainly the desired property.
In Case (c) there is again an isomorphism of the form (5.4) but the right-hand side vanishes because has no eigenvalue equal to one. So in which has certainly the desired property.
This completes the proof. ∎