### v3

# Nearby Special Lagrangians

Abstract
Let be a Calabi–Yau manifold and a closed connected embedded special Lagrangian; closed Lagrangians mean compact Lagrangian submanifolds without boundary. We prove that if the fundamental group is abelian then there exists a Weinstein neighbourhood of in which every closed irreducibly immersed special Lagrangian with unobstructed Floer cohomology is close to We prove also that if is virtually solvable then for every positive integer there exists a Weinstein neighbourhood of in which every closed irreducibly immersed special Lagrangian of degree and with unobstructed Floer cohomology is unbranched; that is, the projection is a covering map. We prove a stronger statement when is finite and a weaker statement when has no non-abelian free subgroups. The conditions, the Floer cohomology condition and the special Lagrangian condition are all essential as we show by counterexamples.
Contents
1 Introduction
We begin by defining the basic notions we will use.
Definition 1.1.
A Calabi–Yau manifold is a complex manifold of complex dimension equipped with a Kähler form and a nowhere-vanishing holomorphic form Joyce [38] and other authors call this an almost Calabi–Yau manifold to indicate that this need not be Ricci-flat. We do more symplectic topology in the present paper and therefore leave out the word almost so as not to be confused with almost complex structures.
A Lagrangian in the Calabi–Yau manifold is a Lagrangian immersion whose intersection points are transverse double points. We refer only to the image when we want to save notation. We say that is closed if is compact without boundary, embedded if is an embedding of manifolds, irreducibly immersed if is connected, and special of phase if is a nowhere-vanishing real -form on We then orient by Plainly by a special Lagrangian we mean a special Lagrangian of phase
Let be a Calabi–Yau manifold and a closed embedded special Lagrangian. McLean [49] determines those closed special Lagrangians which are close to We study the version of this problem; that is, we seek a neighbourhood of in which we can determine as many closed special Lagrangians as possible.
There are some circumstances in which no non-trivial nearby special Lagrangians exist. Tsai and Wang [70, 71] prove indeed that if the Kähler metric on is Ricci-flat and the induced metric on Ricci-positive then there exists a neighbourhood of in which every closed special Lagrangian has image equal to
There are other circumstances in which many nearby special Lagrangians exist. For instance, let be a hyperKähler -manifold and a compact Riemann surface of genus embedded in as a complex Lagrangian with respect to one of the three complex structures; then the hyperKähler rotations of Hitchin spectral curves are nearby special Lagrangians. We make in Example 5.13 a higher-dimensional version of this, using special Lagrangian surgeries. Donaldson and He [16, 32, 33] construct more examples of branched special Lagrangians which are locally modelled upon the product of and a -fold spectral curve. These special Lagrangians are all branched as we define now.
Definition 1.2.
Let be a Calabi–Yau manifold and a closed connected embedded special Lagrangian. A Weinstein neighbourhood of means a neighbourhood which may be embedded, as a compact disc sub-bundle, into the cotangent bundle with its standard symplectic form. For a closed irreducibly immersed special Lagrangian we define its degree to be that of the composite of the immersion and the projection the degree is an integer because special Lagrangians are oriented. We say that is unbranched if the map is a finite cover (of degree equal to what we have just defined). This is equivalent to saying that there exists a finite cover such that the pre-image of under the induced map is a union of perturbations of (if is embedded, this union is disjoint); here a perturbation means the graph over of some closed -form. We say that is branched if is not unbranched.
We prove that no branching occurs under the hypotheses which are weaker than that of Tsai and Wang, except that we can treat only those Lagrangians with unobstructed Floer cohomology. We make a brief definition of the key words about this, referring to §2 for the actual contents.
Definition 1.3.
Let be a Calabi–Yau manifold, a closed connected embedded special Lagrangian, and a Weinstein neighbourhood of Let be a Novikov field in the sense of Definition 2.2. As is an exact symplectic manifold we can use a result of Fukaya, Oh, Ohta and Ono [25] by which we can allow to have characteristic Let be a special Lagrangian with immersion map By a filtered local system of we mean a finite-rank non-zero local system on of valued -vector spaces in the sense of Definition 2.9. We say that has unobstructed Floer cohomology, or for short, unobstructed if there exists a filtered local system of and a bounding cochain of in the sense of Definition 2.49. The zero-section for instance has unobstructed with any choice of local system on this because it has essentially no holomorphic discs. The same holds for its perturbations too.
We state now the main results of the paper.
Theorem 1.4.
Let be a Calabi–Yau manifold, a closed connected embedded special Lagrangian, and a Weinstein neighbourhood of then the following hold.
Let the fundamental group be finite and let be a closed irreducibly immersed special Lagrangian of phase with unobstructed; then and
If is virtually abelian (that is, having a finite-index abelian subgroup) then there exists a neighbourhood of such that every closed irreducibly immersed special Lagrangian with unobstructed and contained in is unbranched; moreover, if is (strictly) abelian then is a perturbation of
If is virtually solvable then for every there exists a neighbourhood of such that the following holds: let be a closed irreducibly immersed special Lagrangian of degree contained in and which has unobstructed with respect to some filtered local system of rank then is unbranched.
If has no non-abelian free subgroups then for every there exists a neighbourhood of such that the following holds: let be a closed irreducibly immersed special Lagrangian of degree contained in and which has unobstructed with respect to some filtered local system of rank over some characteristic-zero Novikov field; then is unbranched.
If is finite as in (i) above then the other hypotheses on in (ii)–(iv) hold too except the strict abelian condition in the latter part of (ii). But the conclusion of (i) is still stronger in two respects: we do not have to shrink the Weinstein neighbourhood and we can include special Lagrangians of any phase. By Myers’ theorem the hypothesis that should be finite follows from that of Tsai and Wang who suppose that has positive Ricci curvature.
We give now a sketch of the proof of Theorem 1.4. The Lagrangian has unobstructed and defines accordingly, as we shall see in §2, an object of the Fukaya category This is a strict category whose hom spaces are valued vector spaces over some Novikov field We recall in §2.1 that the algebraic closure of is itself a Novikov field. As the bounding cochain of may still be defined with coefficients in the larger field, we can suppose that the original Novikov field is already algebraically closed.
We embed into the wrapped Fukaya category which include objects supported on compact Lagrangians with boundary. Fix a point and by an abuse of notation identify the non-compact fibre with the compact fibre in over the same point This defines then an object of which we denote by the same symbol
The hom spaces of and are not only filtered but also -graded -vector spaces. We denote by the hom space from to Since the Lagrangian is special it follows that is non-zero and supported in a single degree; the non-zero property will be proved in Corollary 2.56, and the single degree property in Lemma 3.3.
We prove then that satisfies the split-generation and other properties. These are known for exact Lagrangians. But no non-trivial special Lagrangian perturbations of are exact, so we do include non-exact Lagrangians in and show that the split-generation and other statements are still valid. These imply then that is a certain representation of the fundamental group and that this representation determines the object up to isomorphism in the cohomology category
Suppose now that is strictly abelian as in the last part of Theorem 1.4 (ii). Since is algebraically closed we get then a one-dimensional sub-representation of This may be realized by a special Lagrangian perturbation of with some local system on it. Then by a Thomas–Yau theorem, Proposition 2.67 (i), we can show that the original special Lagrangian agrees with the perturbation of
Theorem 1.4 (i) is proved in §3.1, which is simpler than that of (ii). The chief reason is that the algebraic problem about the representation will be trivial after we pass to the universal cover of
Theorem 1.4 (iii) is proved in the same way as (ii) is proved, but the representation theory argument will be more complex. The new part will be the use of Lie–Kolchin’s theorem and we shall therefore be restricting the representation to a finite-index subgroup of Denote by the corresponding finite cover. Using the upper bound given in (iii) we prove in §3.4 that the Green operator over (a right inverse to the linearized operator in McLean’s deformation theory) is bounded above by a constant independent of the original and This implies that the perturbation argument will work in after pulling back the relevant objects in The result is that is the immersed image of some perturbation of the zero-section So is unbranched as we have to prove.
The proof of Theorem 1.4 (iv) is nearly the same as that of (iii). The new part will be the use of Tits’ alternative theorem. We shall need the characteristic to be zero for an entirely algebraic reason which we explain in §3.6.
The hypotheses on in (ii)–(iv) are essential to Theorem 1.4 (ii)–(iv). Suppose for instance that is a compact Riemann surface of genus Then has non-abelian free subgroups, satisfying none of the hypotheses in (ii)–(iv). Feix [19] and Kaledin [39] construct a hyperKähler metric on a neighbourhood of the zero-section in which we can accordingly define the hyperKähler rotations of spectral curves. These have unobstructed with respect to any local systems over any Novikov field, because they are of real dimension (for which trivial bounding cochains exist).
The strict abelian condition (rather than the virtually abelian condition) is essential to the last part of Theorem 1.4 (ii). We prove in Corollary 4.7 that there are multi-valued graphs over a certain flat manifold This has a non-trivial finite cover by and the fundamental group is accordingly not abelian.
We show that Lagrangians being special is indispensable to Theorem 1.4 (ii). We prove in Example 5.11 that there exists an example of with and branched nearby non-special Lagrangians with unobstructed. We prove in Corollary 5.3 that these are not special even in the weak categorical sense.
We prove in Corollary 5.14 and in Example 5.15 that there exist branched nearby special Lagrangians with obstructed. We verify directly that these properties hold except the obstructed condition. So if they had unobstructed, Theorem 1.4 would be false. Hence it follows that they have obstructed. There is a further consequence, Corollary 5.16, which is entirely about Morse -forms on having nothing directly to do with special Lagrangians.
We begin in §2 with the more formal treatment of Fukaya categories. In §3 we prove Theorem 1.4. In §4 we recall Tsai–Wang’s result and give the examples related to it. In §5 we give the branched examples and prove the relevant results.
Acknowledgements. This paper is based upon the discussions we had in Fall 2016 at the Institute for Advanced Study in Princeton. We are both grateful to David Treumann for pointing out that spectral curves are the typical counter example to our results, to Dominic Joyce for helping us to find the examples with obstructed, and to the referees for pointing out the mistakes in the earlier versions of the paper and helping us to improve the presentation. The second author would also like to thank Mark Haskins and Martin Li for interesting conversations which inspired him to begin this work.
The first author was supported by the Erik Ellentuck Fellowship, the IAS Fund of Math, NSF grants DMS-1308179, DMS-1609148 and DMS-1564172, DMS-2103805 and by the Simons Foundation through its “Homological Mirror Symmetry” Collaboration grant. The second author was supported by the WPI project at the Kavli Institute for Physics and Mathematics of Universe, by the JSPS grants 16K17587, 18J00075 and 21K13788, and by the NSFC grant 11950410501.
2 Fukaya Categories
In this section we explain the definition and properties of the Fukaya categories concerned with the proof of Theorem 1.4. We use pseudo-holomorphic curve moduli spaces to define them. As we include non-exact Lagrangians we cannot follow the direct perturbation method of Seidel [61]. We follow instead the virtual perturbation method of Akaho–Joyce [9] and Fukaya, Oh, Ohta and Ono [23]. Our treatment is however slightly different from theirs. On the one hand, ours is simpler in that we shall need to define Floer cohomology groups only as -vector spaces rather than as modules as in [9, 23]. The latter has more information but we shall not need it. On the other hand, for our theorem (Theorem 1.4) to hold, we need to include filtered local systems (in the sense of Definition 2.12) in the Fukaya categories. Although this is not explicitly done in [9, 23] it is a straightforward modification.
Also, as we are concerned with wrapped Fukaya categories we need to include countably many Lagrangians (in [9, 23] the authors deal with single or finitely many Lagrangians). We do this by the same induction process as in [9, 23]. Recall that we can deal only with finitely many kinds of Kuranishi spaces in doing the virtual counts [23, §7.2.3] and the results are accordingly a finite approximation of what we want. These are called gapped categories in our simplified notation, Definition 2.14 (or gapped algebras in their notation [9, 23]). By the obstruction theory arguments of [9, 23] we can pass to the limit called gapped categories. In this method the issue [23, §7.2.3] above arises every time we have infinitely many things. We do have a new parameter which is the number of Lagrangians. To deal with this we modify the obstruction theory arguments of [9, 23], which we do in Theorem 2.26.
We thus obtain the gapped categories we shall need. (These are curved categories, including the terms. We make them into strict categories by adding bounding cochains, following the references [23, 9].) We define then the wrapped Fukaya categories by the localization method, which has been used already in the literature [29] (the idea is due to Paul Seidel and the first author). The advantage of this is that we can make the definition only from the algebraic properties of Hamiltonian continuations, without going back to the study of pseudo-holomorphic curve moduli spaces.
We prove the split-generation theorem, Theorem 2.86, by a method of Paul Seidel and the first author (taken from their unpublished work). This is close to the proof of Fukaya, Seidel and Smith [26] for exact Lagrangians. They use Lefschetz fibrations, Seidel long exact sequences, and equivariant Fukaya categories; and we replace the last ingredient by the use of Viterbo restriction functors. When working with non-exact Lagrangians, the first two ingredients have already appeared in the literature, so we focus on the discussion of Viterbo restriction functors in the non-exact setting. After providing the argument for this generalisation, the rest of the proof is straightforward.
There is another method [2] of proving the split-generation theorem, using symplectic cohomology (for time-one periodic Hamiltonian orbits). One should then prove that it has the same relation as in the exact case to the algebraic invariants obtained from the Fukaya category. One could do this in closed symplectic manifolds (as will appear in [6]) and for monotone Lagrangians (as is done already in [57]) but not yet in the circumstances above.
We begin in §2.1 by defining Novikov fields and explaining their properties we shall need. In §2.2 we define the gapped categories and provide the obstruction theory arguments. Apart from the issue mentioned above, the major difference from the references [23, 9] is that we extend the definition of gapped algebras; see Remark 2.15.
The gapped structure will be used only for the process of defining the curved categories, which we shall in practice be able to treat as filtered categories (in the sense of Definition 2.14) forgetting the gap conditions.
In §§2.3 and 2.4 we define the curved category of a symplectic manifold In §2.5 we define the strict categories including the bounding cochains. In §2.6 we recall the Thomas–Yau theorems we shall need. In §2.7 we explain the effect of symplectic and Hamiltonian diffeomorphisms upon the Fukaya categories. In §2.8 we introduce the wrapped Fukaya category by the localisation method. In §2.9 we define the exact Fukaya category and show that its cohomology category is equivalent to that in the literature.
We turn in §2.10 to the proof of results in cotangent bundles. We prove two lemmas we shall need to deal with perturbations of in the proof of Theorem 1.4 (ii)–(iv). In §2.11 we prove the split-generation theorem including non-exact Lagrangians. In §2.12 we deal with the Yoneda functors obtained from the split-generation theorem.
2.1 Novikov Fields
We begin by recalling the definition of non-Archimedean valued fields. We call them valued fields for short.
Definition 2.1.
By a valued field we mean the pair of a field and a function called a valuation, such that and for we have and By a valued field extension of we mean a valued field such that is a sub-field of and the two valuations agree on We use the same symbol for different valuations.
Every valued field has a sub-ring defined by which is a local ring with maximal ideal in other words, is a unit if and only if
Let be a valued field and define a function by for Then for the strong triangle inequality holds and there is a distance function defined by making into a metric space We call a complete valued field if is complete.
Definition 2.2.
Define the minimum Novikov field to be the Novikov field in the literature [9, 23] without the Maslov index factor in their notation. We begin by recalling this. Let be a field and its group ring. Denote by the formal variable corresponding to so that every element may be written as with a finite subset of Define a function by Then the function defined by for satisfies the strong triangle inequalities, making into a metric space; and is its completion. The convolution products extend to the completion, making into a ring. This is in fact a field because of the inversion formula for with The function extends to the completion, making it into a complete valued field We denote an element of by an infinite sum with a discrete subset of
By a Novikov field we mean a complete valued field extension of We call the ground field of the Novikov field.
We prove now that algebraically closed Novikov fields exist for arbitrary characteristic. We give two proofs: one is Krasner’s lemma, Lemma 2.5, a general fact about valued fields, and the other is an explicit construction in Proposition 2.6 (ii).
Let be a valued field. Define a subring defined by This has a unique maximal ideal and the residue field The value group is a subgroup of defined by We recall the following basic facts:
Proposition 2.3.
Let be a complete valued field, and an algebraic extension; then the valuation of extends uniquely to so that:
if has finite degree, the valuation on is complete; and
if has infinite degree, the valuation of is strictly not complete.
Also, if is finite then
More precisely, there exist two intermediate fields of the maximal unramified field and the maximal tamely-ramified field such that:
if has characteristic then and if then is a power of
Proof.
We consult Neukirch [51]; the referred statements in what follows are all from Chapter II of this book.
The unique extension is by Theorem 4.8. By Proposition 4.9 the extended valuation is equivalent to the norm on a direct sum of whose elements have only finitely many non-zero components even if they are infinite dimensional. So they are complete if and only if finite-dimensional.
The inequality (2.1) is from Proposition 6.8. The field is defined in Definition 7.4, and in Definition 7.10 but under the hypothesis that has positive characteristic; if has characteristic we define The parts (i), (ii) and (iv) are then obvious. The part (iii) follows from Proposition 7.7; this is stated again in the positive-characteristic case, but its proof applies also to the characteristic-zero case. ∎
Here the extension is called the wild part of which is in general difficult to study. But it has degree a power of and for Galois extensions of such a degree, there is a classification result:
Proposition 2.4 (Artin–Schreier).
Let be any field of characteristic Then for every either:
the polynomial has a root in which case its roots are all in or
is irreducible, in which case is a degree- Galois extension of
Conversely:
every degree- Galois extension of is of the form for some or more generally
every finite Galois extension of of degree equal to a power of is a repeated extension of this kind.
Proof.
Finally we recall:
Lemma 2.5 (Krasner).
Every valued field has a complete algebraically-closed extension.
Proof.
We turn now to the second method. Let be a field. In Definition 2.2 we have defined the minimum Novikov field with ground field We define the maximum Novikov field by
where Recall that being well ordered means that every non-empty subset has a minimum. In particular, is half discrete in the sense that for every there exists such that
because we can take to be the minimum of The sums and products in are defined in the natural way: for we can define
because is well ordered; and
because is finite, which follows since are half discrete.
The valuation is defined by setting and by taking in (2.3) the least Then is a complete valued field. We have
We prove now the following proposition. Part (ii) with algebraically closed is the result we have promised to prove. Recall that a field is a perfect field if either this has characteristic or the ring homomorphism defined by is surjective.
Proposition 2.6.
are both perfect fields.
is maximal in Krull’s sense [44]; that is, has no non-trivial valued field extension with and In particular, if is algebraically closed, so is
If is algebraically closed of characteristic so is
If has positive characteristic then is not algebraically closed, and its algebraic closure is strictly not complete.
Remark 2.7.
These results do not assert that is the algebraic closure of For rational-power series, the algebraic closure is described in Kedlaya [42].
Proof of Proposition 2.6.
For (i) let be either or We suppose and prove that for every there exists with For we can take For we can write For each choose a root of the polynomial which exists because is algebraically closed. Put As in (2.6) the binomial expansion is valid so that
where the second equality follows since has characteristic Thus is an element we want.
The former part of (ii) is an old result [40, 44]. For the latter, if is algebraically closed, then is algebraically maximal and is already maximal in so has no non-trivial finite extension, or equivalently is algebraically closed.
The former part of (iii) is proved by Fukaya, Oh, Ohta and Ono [24, Lemma A.1] but we give another proof, using Proposition 2.3 with Suppose therefore that is a finite extension, and note that:
since is algebraically closed it follows that the inclusions in (ii) are all equalities, so
since it follows that the inclusions in (iii) are all equalities, so and
because has characteristic zero.
Consequently, that is, has no non-trivial finite extension, or equivalently, is algebraically closed.
For the latter part of (iii) let and consider the Artin–Schreier polynomial
If this had a root of the form then by induction on it must be of the form
so is not algebraically closed.
We show next that the algebraic closure of has infinite degree over By an Artin–Schreier element over we shall mean a root of an Artin–Schreier polynomial over Let be the field extension which contains all the Artin–Schreier elements over We prove the following three facts in this order:
if with and if are respectively roots of the polynomials then
if is an Artin–Schreier element over then where is a root of the polynomial for some with and every a root of the polynomial for some and and
if the algebraic closure of had finite degree over we should have where are such as in (II) and are under the same condition as in (II) but that may of course be different.
For (I) note that is a root of the polynomial so for some which implies as we want. For (II) write as a root of the polynomial and recall that has by definition only finitely many terms of negative powers with respect to that is, and Applying to this the part (I) repeatedly, we see that as we want. For (III) it is clear from definition that On the other hand, by our hypothesis is a finite extension and so is obtained from by adding finite Artin–Schreier elements. Hence using (II) repeatedly we get the other inclusion we want.
Now in (III) write as a root of the polynomial for some and Then for some the algebraic closure of so
But as is an infinite extension there is some that is not in Then by (2.11) the Artin–Schreier element a root of the polynomial is not in which contradicts the definition of
So the algebraic closure of has infinite degree over and by Proposition 2.3 (b) it is strictly not complete. ∎
Remark 2.8.
The rôle of Artin–Schreier polynomials is essential to the proof above. In fact, if is algebraically closed then every finite normal extension is a repeated Artin–Schreier extension, as we show now. In the notation of Proposition 2.3 with we have again On the other hand, by (i) above, is automatically Galois. To this we can apply Proposition 2.4 so is a repeated Artin–Schreier extension as we want.
2.2 Gapped Categories
We begin by defining the notion of valued vector spaces. As in Definition 1.3 we will use these to define filtered local systems of Lagrangians.
Definition 2.9.
Let be a Novikov field. By a valued -vector space we mean the pair of a -vector space and a function called also the valuation, such that for we have and for we have where we use both and
Define a function by for Then the strong triangle inequality holds for and in particular, there is a distance function defined by making into a metric space We call a complete valued vector space if is a complete metric space. The valued vector space structure of extends to the metric space completion of making it into a complete valued vector space. We call this the completion of
Let be valued -vector spaces and a -linear map which is continuous with respect to the metric space structures above. Then there exists by that is, is the operator norm of the bounded operator We say that is filtered if We call it a filtered isomorphism if is a linear isomorphism with filtered. The latter is equivalent to or to because for linear maps between valued vector spaces.
Let be valued -vector spaces and a multi-linear map. Define then by
We say that is filtered if We denote by the set of filtered multi-linear maps from to The pair has the obvious structure of a valued -vector space. If is complete then so is
Let be a family of complete valued -vector spaces. Then there exists a valuation defined by The complete direct sum is the completion of the direct sum
We give a basic example of finite-dimensional valued -vector space.
Example 2.10.
Let be an integer and an -dimensional valued -vector space. We give another valuation. Let be a basis with which exists because for every Then has a valuation defined by for Note that by the strong triangle inequality we have for every We call the valuation with respect to if the equality holds for every
More generally, when the basis is not specified, we call plainly an valuation if it is so with respect some basis of with Recall from Definition 2.2 that we have the subring Regard as an module and define a submodule by Then is an valuation if and only if is a free module. More precisely, given a finite set with we have if and only if is the free module with basis
We give two other equivalent conditions. Note that is a submodule of and the quotient has the vector space structure over We prove that the following three conditions are equivalent: (i) is a rank- free module; (ii) is a finitely generated module; and (iii) the -vector space has dimension Clearly (i) implies (ii). We prove that (ii) implies (iii). Take such that their images are a -basis. Then as is finitely generated we can use Nakayama’s lemma [45, Chapter X, Lemma 4.3] (for non-Noetherian local rings) which implies So and which is the condition (iii). We prove now that (iii) implies (i). Take such that their images are linearly independent over Then and we show that the equality holds. Otherwise there is some By re-ordering we can write where with Then
and have positive valuation. So in But none of is zero, which contradicts that are linearly independent. Thus (i)–(iii) are equivalent.
Remark 2.11.
Consider the category whose objects are -dimensional valued -vector spaces, whose morphisms are filtered homomorphisms. Then there is an equivalence from the category of rank- free modules to that of -dimensional valued vector spaces with valuations. The function between the object sets is defined by Each morphism is mapped to
When every valuation is the valuation. So the category of rank-one free modules is equivalent to that of one-dimensional valued vector spaces.
We define then the filtered local systems.
Definition 2.12.
By a filtered local system over a topological space we mean one of non-zero finite-dimensional valued vector spaces whose parallel transports are filtered homomorphisms of valued vector spaces.
We make now the definitions about discrete sub-monoids of These will be concerned with the areas of pseudo-holomorphic curves with Lagrangian boundary conditions.
Definition 2.13.
Denote by the set of non-negative integers. Let be a discrete sub-monoid, that is, a discrete subset containing and closed under addition. By a decomposition of we mean with and with such that By a reduced decomposition of we mean and with and see Equation (2.16) below for a typical formula where this notion is used. Every decomposition with and for some defines indeed the reduced decomposition
Choose a total order on such that there is an ordered set isomorphism which we denote by and if are as above a decomposition of then
with equality if and only if at least of are equal to Setting and we get that is, is the least element of In other words, if and only if
We choose the order for every discrete sub-monoid so that if are two discrete sub-monoids with then the induced map will preserve the orders. Here is an explicit way of doing this. Set and for set
The maximum makes sense because is discrete and has the least positive element. It is clear that for For define by either or and Then we can check all the conditions above.
Let be a Novikov field, which we shall fix in this section. We define the notion of gapped categories, over The morphisms spaces are -graded. The structures are curved, including the terms.
Definition 2.14.
We begin with a filtered curved category which consists of an object set a family of -graded complete valued -vector spaces, and a family
and such that
where and are any integers with For we have for every and (2.16) means
Suppose now given By a gapped category we mean an object set a family of -graded complete valued -vector spaces, a discrete sub-monoid and a family with and such that
where are as in (2.16) and the sum is taken over reduced decompositions s of Here the hom space
consists of filtered vector space homomorphisms (as in Definition 2.9); the superscript of refers to the degree, which has nothing to do with the filtrations. In (2.17) the operators are well-defined because by (2.13).
When the gapped category defines a filtered category of the same object set, the same morphism spaces, and the operators defined by where for and for The sum is well-defined because is complete.
Remark 2.15.
The definition of implies that every is a filtered homomorphism between valued vector spaces. This will be crucial to defining the Fukaya categories with filtered local systems. In the references [9, 23] they define to be linear over the residue field determined uniquely from the sum this is not true in our circumstances.
They also work with gapped algebras which are gapped categories with a single object; or they are expressible as the pair of a -graded complete valued -vector space and a family with satisfying the same equations. If is a gapped category then the complete direct sum, has a gapped algebra structure defined in the obvious way. Conversely, let be a gapped algebra with such that the induced map is zero unless and if this is the case then has image in Then and define a gapped category.
We define the truncations of gapped functors, and the full subcategories of gapped functors.
Definition 2.16.
Let with and a gapped category. The truncation of means the same object set the same morphism spaces and the subfamily these define a gapped category. By a full subcategory of we mean a subset the subfamily and the subfamily these define a gapped category.
We define the gapped functors,
Definition 2.17.
Let be two filtered categories. By a filtered functor from to we mean a function between the object sets and a family
and such that for every the element has positive valuation, that is, and for with we have
where are as in (2.16) and are any integers with The right-hand side of (2.20) converges because the terms have positive valuation.
Suppose now that are gapped categories, Note that any two discrete sub-monoids of are contained in some single discrete sub-monoid of By a gapped functor from to we mean a function between the object sets; a discrete sub-monoid with respect to which are gapped; and a family
with and such that
where are as in (2.17) and any decomposition of The operators are well-defined because by (2.13).
When the gapped functor defines a filtered functor between the corresponding filtered categories, with where for and for The sum is well-defined because is complete.
We define the identity functors of gapped categories,
Definition 2.18.
Let be a filtered category; we leave out for short. The identity functor is defined by taking the function to be the identity; for to be the identity for and for with
Suppose now that is a gapped category, Then the identity functor is defined by taking the function to be the identity; to be the identity for and for with
We define the composites of gapped functors,
Definition 2.19.
Let be filtered categories and filtered functors. Then the composite functor means the ordinary composite and the family with defined by
where are as in (2.20).
Suppose now that are filtered categories, and filtered functors. Then the composite functor means the ordinary composite a discrete sub-monoid with respect to which are gapped; and the family with and defined by
where are as in (2.22). Whether or not (2.23) holds is independent of the choice of with respect to which are gapped; for we can first make the smallest choice of and then show that for not in this smallest
We define the homotopies of gapped functors, We make a rather strong hypothesis on the functions between object sets.
Definition 2.20.
Let be filtered categories and filtered functors. Suppose that induce the same function between the object sets, which we denote by for short. By a homotopy we mean a family
and such that (i) for every the element has positive valuation, that is, and (ii) for with we have
where are as in (2.20); and any integers with The right-hand side converges by (i) above.
Suppose now that are gapped categories, and gapped functors with the same function between the object sets. By a homotopy we mean a discrete sub-monoid with respect to which are both gapped and a family
with and such that
where are as in (2.22) and any decomposition of The operators are well-defined because are all by (2.13).
When the gapped homotopy defines a filtered homotopy between the corresponding filtered functors, with where for and for The sum is well-defined because is complete.
Remark 2.21.
We define the notion of being homotopic for gapped functors,
Definition 2.22.
Let be gapped categories and gapped functors which induce the same function between the object sets. We say that are homotopic if there is a homotopy from to
The following is a straightforward extension of Fukaya, Oh, Ohta and Ono’s result for gapped algebras.
Lemma 2.23 (cf. [23, Proposition 4.2.37]).
Let be gapped categories and fix a function Then being homotopic is an equivalence relation for those gapped functors from to inducing the given function ∎
We define the homotopy equivalences of gapped categories,
Definition 2.24.
Let be gapped categories of the same object set and a gapped functor which induces the identity on the object set. We call a homotopy equivalence if there exists a gapped functor which induces the identity on the object set and such that are homotopic to the identity functors of respectively. We say that are homotopy equivalent if there exists a homotopy equivalence
The following is also a straightforward extension of Fukaya, Oh, Ohta and Ono’s result for gapped algebras.
Lemma 2.25 (cf. [23, Corollary 4.2.44 and Remark 7.2.71]).
Being homotopy equivalent is an equivalence relation for gapped categories of the same object set.
Let be gapped categories of the same object set and a gapped functor which induces the identity on the object set. Then is a homotopy equivalence if and only if for any the co-chain map induces an isomorphism of the cohomology groups. ∎
Here is the obstruction theory result for gapped categories which extends that of Fukaya, Oh, Ohta and Ono [23, Theorem 7.2.72].
Theorem 2.26.
Let be a gapped category, and a full subcategory of the truncation of Let be a gapped category with and a homotopy equivalence which induces the identity on the object set. Then there exist a gapped category with which contains as a full subcategory in its truncation; and a gapped functor which induces the identity on the object set, whose truncation agrees with on
Proof.
Recall from definition that the truncation of is a family of co-chain complexes. We assign to a co-chain complex and a co-chain map For set and For with or set and define to be the identity map. In both cases the co-chain map induces an isomorphism of the cohomology groups.
We extend these co-chain complexes to a gapped category and these co-chain maps to a gapped functor where is the truncation of We do this by an induction on Suppose therefore that we have defined for with and for We extend these to a gapped category and a gapped functor
Denote by the element with We assign to some filtered homomorphisms which satisfy the equations. For we use the given and otherwise we introduce some new Define a degree-one -linear map from to itself by
This is well-defined because implies And Define an obstruction co-cycle by
where or Doing the same computation as Fukaya and others [23, (4.5.2.1)] we find that is, is certainly a co-cycle. Their computation [23, (4.5.2.4)] shows also that the homotopy equivalence induces an isomorphism from the -cohomology group to the -cohomology group mapping the cohomology class of to that of But the latter vanishes for the obvious reason: is already an category. Thus is a co-boundary, written as for some
Define then an obstruction co-cycle in by sending to
where or and are a decomposition of with so that As is a co-chain equivalence we can modify if we need, so that the corresponding obstruction co-cycle will vanish. So there is a gapped functor as we want, completing the induction.
We have thus obtained a gapped category and a gapped functor that is, are given for every with Extend these to with by repeating the obstruction theory arguments above; then the proof is complete. ∎
2.3 Fukaya Categories
We begin by recalling the definition of symplectic manifolds with contact boundary. As in Definition 1.2 the Weinstein neighbourhoods are the typical example.
Definition 2.28.
A compact symplectic manifold with contact boundary is a symplectic manifold which is given a -form near such that and the vector field given by is outwards pointing on . We call a Liouville -form of
In these circumstances there exists near a unique smooth function with values in outside with values on and such that We call the radius function near We say that an almost complex structures is compatible with the contact boundary if it is compatible with in the usual sense, if near and if is invariant near under the flow of
If there exists a global -form with and satisfying the conditions above then we call a Liouville domain. By a Liouville subdomain of a Liouville domain we mean a compact co-dimension submanifold with smooth boundary at every point of which the vector field points outwards. Then is itself a Liouville domain. Its radius function is well defined in near and in fact extends to a neighbourhood in of
We consider the following class of Lagrangians:
Definition 2.29.
Let be a compact symplectic manifold with contact boundary, and a Liouville -form of it. Let be a compact manifold possibly with boundary and a Lagrangian immersion. We say that is generic if its every self-intersection point is a transverse double point outside We say that has Legendrian collar if near and has regular value (this is automatically true if is closed).
Suppose now that and are two Lagrangians in given by immersions and We say that are a generic pair if the immersion is generic.
We recall now the definition of -graded symplectic manifolds which uses almost complex structures on the ambient symplectic manifold; for other equivalent definitions see for instance Seidel [59].
Definition 2.30.
Let be a symplectic manifold of dimension Then a -grading of is the homotopy class of a pair where is a compatible almost complex structure and a nowhere-vanishing section of the complex line bundle where is the complex line bundle over of forms relative to We say that is -graded if it is given a -grading.
Remark 2.31.
A -grading of exists if and only if where denotes the first Chern class of
We recall next the definition of -graded Lagrangians, which is a generalization of special Lagrangians in the sense of Definition 1.1.
Definition 2.32.
Let be a -graded symplectic manifold of dimension and let represent the -grading. It is known then that for each Lagrangian immersion there exists a unique function such that where is some nowhere-vanishing section of the real line bundle where is the real line bundle over of -forms. We call the phase function of
By a -grading of we mean a lift of to the universal cover By a -graded Lagrangian we mean the pair of and a -grading of it. We say that is special of phase with respect to if is locally constant. In particular, a connected -graded special Lagrangian is equivalent to the pair of and a constant phase Define an operator by sending the -graded Lagrangian to where is the lifted phase function so that we can do define This will induce the shift operator of the triangulated category once has been defined.
Remark 2.33.
The Lagrangian immersion admits a -grading if and only if the group homomorphism vanishes.
We define next the Maslov indices.
Definition 2.34.
Let be a -graded symplectic manifold of dimension and let represent the -grading. Let be a generic Lagrangian immersion. We consider the fibre product
which consists of the diagonal identified with and the self-intersection pairs in Suppose that is -graded by a phase function and we define then a Maslov index function Set on the diagonal We associate to each pair an integer called the Maslov index. Put and take a -linear isomorphism which maps to to and to for some These are unique up to order so we can define
The definition of phase functions implies readily that is in fact an integer.
Suppose now that are a generic pair of Lagrangians, -graded by and respectively. The fibre product is then a finite set and we define the Maslov index function by sending to the right-hand side of (2.29) above with in place of respectively.
Remark 2.35.
Since it follows by (2.29) that
In particular, if then the integer must lie in the closed interval and if then
Example 2.36.
Take with co-ordinates and Define a -grading by Take and For define
so Define a -grading by starting with on and lifting the path Then and in particular The Maslov index of at the intersection point is now equal to This agrees with the Morse index of the quadratic function on Our convention is thus the same as Seidel’s [59, §2d, (v)].
We extend the definition of relative spin structures to countably many Lagrangians. These will equip the pseudo-holomorphic curve moduli spaces with natural orientations.
Definition 2.37.
Let be a symplectic manifold, a compact manifold and a generic Lagrangian immersion; we allow to have boundary. By a relative spin structure of we mean the data where is a triangulation of an oriented vector bundle over the -skeleton of an orientation of a triangulation of such that induces a simplicial map and a spin structure of the bundle restricted to the -skeleton of
Suppose now that is a sequence of compact manifolds which may have boundary, and a generic Lagrangian immersion. By a relative spin structure of we mean a sequence such that each is a relative spin structure of each is a subcomplex of and the restriction of each to agrees with
Remark 2.38.
Suppose that that is defined over the whole and that each is a spin structure of over the whole Then each Stiefel–Whitney class is the pull-back of In this case we refer to as the background class of the relative spin structure.
We can in practice take to be embedded in the cotangent bundle over a closed oriented manifold and take where is the projection. So the background class is and the zero-section will be given a spin structure of
We introduce now the notion of branes:
Definition 2.39.
Let be a compact symplectic manifold with contact boundary, and a Liouville -form of it. By a brane on we mean a pair where
is a compact -graded generically-immersed Lagrangian with Legendrian collar; and
is a filtered local system over the domain of the immersion
We write only in place of when we want to save notation.
We choose countably many branes and a Novikov field.
Hypothesis 2.40.
Let be a compact symplectic manifold with contact boundary, and a Liouville -form of it. Let be a countable set of branes such that any two distinct Lagrangian immersions in are a generic pair in the sense of Definition 2.29; we allow some of the Lagrangian immersions to be the same. Fix a Novikov field When has characteristic we suppose that has no non-constant pseudo-holomorphic sphere (cf. [25]). When has characteristic we suppose that and the underlying Lagrangians have a relative spin structure with respect to some order of the Lagrangians.
We introduce also the homologically perturbed Floer complexes:
Definition 2.41.
Let be as in Hypothesis 2.40 and two elements. Denote by and the projections. The homologically perturbed Floer complex associated to the pair is the -graded finite-dimensional valued -vector space
where denotes the degree shift by Maslov indices. Since the fibres of are valued vector spaces it follows that so are and
Remark 2.42.
The definition (2.31) is actually the result of applying the filtered homological perturbation lemma [23, Theorem A] to the singular or de Rham model in the literature. More precisely, in any construction of Lagrangian Floer cohomology groups we have to choose a chain model of the Floer complex with Akaho–Joyce [9] and Fukaya, Oh, Ohta and Ono [23] both take the singular chain model. Fukaya [21, 22] takes the de Rham model assuming that the ground field contains
We come now to the construction of the curved Fukaya category This is not yet the Fukaya category used in the Introduction.
Theorem 2.43.
Let be as in Hypothesis 2.40 and put Then has the structure of a gapped category whose morphism spaces are given by (2.31) and which has the following properties:
the structure is curved in such a way that if if has no self-intersection point of index or and if does not bound any non-constant pseudo-holomorphic discs, then has curvature
for a pair , the differential has leading term and
for a triple which are supported on the same Lagrangian, the leading term of the product map
(2.32) vanishes with the following exceptions: on the diagonal component of this agrees with the usual cup product map
and on each self-intersection pair it is given by the map
2.4 Proof of Theorem 2.43
Definition 2.44.
Let be a compact manifold with boundary, which we suppose oriented if has characteristic Let be a filtered local system on The smooth singular chain complex is the module generated by pairs where is a smooth singular -simplex in and a (flat) section of of non-negative valuation, modulo the relation
The differential on is defined by
where is the face inclusion map.
Let be a compatible almost complex structure on in the sense of Definition 2.28. We introduce now the moduli spaces of -holomorphic discs. Let be underlying Lagrangians in
Definition 2.45.
We denote by the compact moduli space of finite-area stable -holomorphic discs in with boundary points, whose corresponding segments are ordered counterclockwise and lifted to More precisely, for each genus-zero prestable bordered Riemann surface we choose a continuous map as Akaho–Joyce do [9, Definition 4.1]; this is essentially unique. We distinguish and because the evaluations maps, which we will define shortly in (2.36), are different.
Denote by the points corresponding to the marked points on and by the domains of the immersed Lagrangians. For each holomorphic curve we choose a continuous map as Akaho–Joyce [9, Definition 4.2]. We require that the connected components of should map to respectively.
For we denote by the subset made from -holomorphic curves of area
We recall the maximum principle on Riemann surfaces with Neumann boundary conditions, under the assumption that the almost complex structure is compatible with the contact boundary:
Lemma 2.46 (Abouzaid–Seidel [8, Lemma 7.2]).
There exists a compact set of which contains the image of every non-constant -holomorphic curve included in ∎
Hence by the Gromov compactness theorem we get
Corollary 2.47.
Every is compact. ∎
Suppose now that are given local systems respectively. Note that for every there is an evaluation map
defined by where is the lifted map and the marked point on Define by So has local systems of the form for Using the parallel transport map we get a natural map of local systems
Let be a discrete sub-monoid which contains the areas of every pseudo-holomorphic disc with boundary on the underlying Lagrangians of The gapped categories in what follows will be gapped with respect to We fix an order on the countable set and use the same order on the underlying Lagrangians of it. Here are the next three steps:
For each construct a gapped category whose objects are the first Lagrangians and whose morphism space between is (2.35). We denote this morphism space by As Fukaya, Oh, Ohta and Ono [25, Remark 13.5] point out we can work with the whole complex (2.35), not with countably generated sub-complexes of it as in the earlier works [9, 23].
Construct a homotopy equivalence from to the truncation of
Lift the gapped structure of to the gapped structure.
We do (I) as follows. By Corollary 2.47 every moduli space of the form is a finite union of Kuranishi spaces, which are oriented naturally by the relative spin structure. Suppose that for each we are given We choose a multivalued perturbation of the fibre product
where whenever Here we can in fact make a single-valued perturbation when the symplectic manifold is spherically positive in the sense of Fukaya, Oh, Ohta and Ono [25]; this is possible mainly because of Remarks 12.6 and 12.7 of their paper. Combining this perturbation with (2.37) we get an element of which we call the virtual chain for It has coefficients in the ground field by the result of Fukaya, Oh, Ohta and Ono [25]. Varying we get a map
One can prove that there is a consistent choice of the virtual chains above such that defines the gapped structure.
Step (II) is done by a standard cobordism (or bifurcation) argument. Step (III) is a consequence of Theorem 2.26. Once these have been done, we can define a gapped category whose truncation is homotopy equivalent to Applying the homological perturbation lemma to we get a gapped category that we want. ∎
2.5 Bounding Cochains
We define the units and bounding cochains of filtered categories.
Definition 2.48.
Let be a filtered category. By units of this we mean a family of degree-zero morphisms such that for for with for with and for with and for containing at least one of
We call a strict category if for every If it is strict then define the cohomology category to be the associative category consisting of the same objects; the morphism space for and the product map induced from the operators. Define also the category to be the subcategory of consisting of the same objects and whose every morphism space is the degree-zero part
Let be filtered strict categories and a filtered functor. We call a strict functor if This induces then a functor between the cohomology categories, in the obvious way.
Let be any object and a degree-one element with We call a bounding cochain if The sum converges because We define then a filtered strict category Define to be the set of pairs with and a bounding cochain. For set For define by
The sum converges because Since it follows that that is, is a filtered homomorphism. The equations of imply those of Since it follows that Thus is a filtered strict category.
If is a filtered category with units then the strict category corresponding to it has units too. This is a straightforward extension of the algebra case [23, Lemma 5.2.17].
We introduce now the strict category where ‘nc’ stands for not necessarily closed Lagrangians.
Definition 2.49.
We define the strict category to be that made from by using bounding cochains as above. We denote an object of by a triple where is an object of and a bounding cochain in We say that is supported on or that underlies We define the -graded -linear associative category to be the cohomology category of We denote by its morphism space between two objects
The following is a straightforward extension of the work of Fukaya, Oh, Ohta and Ono [23, §7.3] for embedded Lagrangians. For or the de Rham version of this is made already by Fukaya [22].
Theorem 2.50.
The filtered categories have units. ∎
Remark 2.51.
For or we can make the units directly as Fukaya [22] does, using the de Rham model for (2.31). Strict unitality fails for the singular chain model because the perturbations of Kuranishi spaces of the form (2.38) need not be compatible with forgetting marked points. One can however prove that for there exists a homotopy unit of whose leading term is equal to the fundamental class of This may be done by taking fundamental chains of the underlying Lagrangians in and constructing in a consistent way the homotopies between those perturbed Kuranishi spaces for which some of in (2.38) are the fundamental chains, and those obtained from the forgetful maps. The homotopy units become the ordinary units after the homological perturbations.
We recall the notion of zero objects in and give a corollary to Theorem 2.50.
Definition 2.52.
We call a zero object if the unit of is zero.
Corollary 2.53.
If is a zero object, the Lagrangian has at least one self-intersection point of index
Proof.
Since is -graded it follows that the unit in may be cancelled out only by elements of which come from self-intersection points of of index ∎
Corollary 2.53 gives another proof of the following fact.
Corollary 2.54 (Fukaya, Oh, Ohta and Ono [23, Theorem E]).
Every supported on an embedded Lagrangian is non-zero. ∎
Take as in Definition 2.30 so that we can speak of special Lagrangians in Consider nearly special Lagrangians which have phase sufficiently close (in the sense) to a constant function. The estimate (2.30) implies then
Proposition 2.55.
If is supported on a generically immersed nearly special Lagrangian then the Floer cochain group is supported in degrees and accordingly, so is the cohomology group ∎
This and Corollary 2.53 imply
Corollary 2.56.
Every supported on a nearly special Lagrangian is non-zero. ∎
We define the Fukaya category which we use most often.
Definition 2.57.
We denote by the full subcategory of objects supported on closed Lagrangians; that is, the underlying Lagrangian of an object of is a Lagrangian immersion from a compact manifold without boundary.
We prove now a Poincaré duality theorem. We are in the right context for doing so; but we shall not need the result for the later treatment, and the reader in a hurry can proceed safely to Proposition 2.66 below.
Theorem 2.58.
Let be an object of whose local system has one-dimensional fibres. Then there exists a non-degenerate pairing
Proof.
We begin by recalling the notion of filtered bimodules. Let be a -graded -linear filtered curved algebra. By a bimodule over we mean a -graded module equipped with operations
which satisfy the bimodule version of the equation. An bimodule bimodule is filtered if it is equipped with a valuation, so that bimodule operations do not decrease valuations.
Example 2.59.
The algebra itself has the structure of a diagonal bimodule with operations
Example 2.60.
If is an bimodule over then its dual has the structure of a bimodule given by
and
Next let and apply to the single brane the proof of Theorem 2.43. Then we get a gapped algebra We use the same notation for the diagonal bimodule, and denote by its dual bimodule. Recall that the of has leading term equal to the usual differential, whose cohomology group is
We prove
Proposition 2.61.
There exists a -graded -linear natural isomorphism of graded vector spaces
Proof.
Recall that Poincaré duality makes sense because of the relative spin structure data, so that there is a pairing
where we do not have to change the local system term because For an off-diagonal self-intersection pair there is also a pairing
which completes the proof. ∎
The following is the heart of the proof of Theorem 2.58.
Lemma 2.62.
Proof.
Recall that is by construction an inverse limit of some gapped algebras, which we denote by Regard this as a diagonal bimodule with operations of the form
and For each we construct a bimodule map which consists of a collection of
indexed by non-negative integers and . We regard the output as an input by duality, so that we have to construct a compatible collection of maps
Here we use the hypothesis that the local system of has one-dimensional fibres.
In order to construct this map, we use the stable moduli spaces of pseudo-holomorphic discs with boundary points, ordered as in Figure 2.1: the marked point on the left arrow corresponds to the input, and that on the right arrow corresponds to the output.
We choose inductively a family of perturbation data for the fibre products of the form
where denotes the evaluation maps at the boundary marked points ordered along the boundary and are singular simplices in with We make these compatible with the previous perturbations of (2.38) and those for the lower-order maps relative to Then we get a bimodule homomorphism, which we denote by
From the bimodule maps of this form we construct an bimodule map, by an induction and obstruction theory argument similar to that in the proof of Theorem 2.43. We give a brief account of this; for more details see Fukaya, Oh, Ohta and Ono [23, §5.2] (who use their results to prove the deformation invariance of Floer cohomology groups [23, Theorem D] which is in fact a homotopy equivalence of bimodules).
Recall that we have constructed an homotopy equivalence between and the truncation of From this we get an algebra which is homotopy equivalent to and whose truncation is The bimodule map which induces an bimodule map (over the algebra )
which we denote still by In the previous stage we have constructed an bimodule map
This is homotopic to the reduction of (2.53) by a choice of homotopy between the defining perturbation data. Using homological perturbation, and proceeding by induction, we get the desired system of bimodule maps. ∎
Corollary 2.63.
is a homotopy equivalence of bimodules.∎
The bimodule homotopy equivalence induces for every bounding cochain another bimodule homotopy equivalence
relative to the -deformed structure; and for the term
is given by
Hence, passing to the cohomology groups, we get an isomorphism
which completes the proof of Theorem 2.58. ∎
Theorem 2.58 implies
Corollary 2.64.
Let be a non-zero object of whose local system has one-dimensional fibres. Then ∎
Remark 2.65.
This is proved already for embedded Lagrangians without local systems by Fukaya, Oh, Ohta and Ono [23, Theorem E]. The second author [36, Lemma 4.4] gives two other proofs, including immersed Lagrangians. One uses the cyclic symmetry [21] assuming that the ground field contains The other uses open-closed maps [23, Theorem 6.4.2].
Corollary 2.64 gives a sufficient condition for the hypothesis of the following proposition.
Proposition 2.66.
Let be isomorphic in and supported on closed special Lagrangians of phase respectively which need not be modulo Suppose Then ∎
2.6 Thomas–Yau Theorems
We recall a few generalizations of Thomas–Yau’s theorem proved by the second author [36]. In this subsection we work with Calabi–Yau manifolds and integrable complex structures.
Proposition 2.67.
Let be a Calabi–Yau manifold of complex dimension Let be supported on closed irreducibly immersed special Lagrangians of phase respectively which need not be modulo Suppose that either or embedded. Then the following hold.
If then either or is supported in degrees
If are isomorphic in with then
Remark 2.68.
The point of (i) is that the degrees are strictly and It follows immediately from (2.30) that they belong to
Proof of Proposition 2.67.
Let be embedded. Then we can work in a Weinstein neighbourhood of which is a Stein manifold. This is the condition (ii) of the second author’s result [36, Theorem 6.2], which implies (i) above. We prove (ii) now. By Proposition 2.66 the two special Lagrangians have the same phase. Hence it follows by (i) that either or is supported in But the latter is impossible because the graded vector space has two non-trivial degrees with difference So ∎
We recall also the following version. We say that a Calabi–Yau manifold is real analytic if the Kähler form of is real analytic with respect to the underlying real analytic structure of the complex manifold
Proposition 2.69.
Let be a real analytic Calabi–Yau manifold and two closed irreducibly-immersed Lagrangians which are, near both special and graded by Then either agree with each other near or any pair of objects supported on them have Floer cohomology supported in degrees
Proof.
Suppose and denote by the set of intersection points at which have at least one common tangent space. The second author [36, Theorem 5.1(i)] proves then that there exists a neighbourhood of and arbitrarily-small Hamiltonian perturbations of respectively which intersect generically each other with no intersection point in of index or On the other hand, these have no intersection point of index or outside [36, Lemma 2.7]. So they have no intersection point at all of index or Since are both graded by it follows by (2.30) that every intersection point of the two perturbed Lagrangians has degree in So their Floer complex is supported in degrees which completes the proof. ∎
2.7 Hamiltonian Continuations
We return now to the symplectic topology context. We recall the definition of homotopies between strict functors.
Definition 2.70.
Let be strict categories. Then we can define the functor category which is a strict category whose objects are strict functors from to We say that two objects are homotopic if they are isomorphic in Here we do not impose any conditions on the object sets as we do in Definition 2.20.
We say that a strict functor is a homotopy equivalence if there exists such that is isomorphic in to the identity on and isomorphic in to the identity on Seidel [61, Theorem 2.9] proves that is a homotopy equivalence if and only if the induced functor is an equivalence of categories.
We explain now the effect of symplectomorphisms of upon the Fukaya categories. Recall from Definition 2.28 that the radius function is defined near Note also that symplectomorphisms act naturally upon the objects of
Lemma 2.71.
Let be a symplectomorphism isotopic to the identity and which agrees near with the Hamiltonian diffeomorphism of for some Suppose that induces a bijection of Then there exists an equivalence
such that for every the image is given by a bounding cochain for The functor is compatible with composition in the sense that, for a pair and of boundary-linear symplectomorphisms, the composite is naturally isomorphic to
Proof.
The strict category is made from a gapped category whose definition involves the choice of an almost complex structure and other data. We can then define another gapped category by using We want to show that the two gapped categories are homotopy equivalent. This is done by Fukaya, Oh, Ohta and Ono [23, Theorem A] and by Akaho–Joyce [9, Theorem 11.2(a)] for gapped algebras by interpolating between the almost complex structures; we do so while maintaining compatibility with the contact boundary, thus ensuring compactness. They use the obstruction theory arguments for gapped functors [23, Lemma 7.2.129] for gapped algebras. We extend them to gapped categories in the same way as in the proof of Theorem 2.26. Then we get a homotopy equivalence of the two gapped categories. Fukaya [22, Proposition 13.18] proves that this induces a homotopy equivalence of the corresponding strict categories. There is thus an equivalence (2.59) of the cohomology categories.
To find the natural isomorphism between and we begin with homotopies of gapped categories. We lift them to gapped categories by an obstruction theory for homotopies of homotopies. This is done by Fukaya, Oh, Ohta and Ono [23, Corollary 7.2.219] for gapped algebras. We extend it in the same way as in the proof of Theorem 2.26. Then we get a homotopy between and as gapped functors. This induces again a homotopy between the strict categories, which induces in turn a natural isomorphism that we want. ∎
There is more to say if the symplectomorphism in Lemma 2.71 is a Hamiltonian diffeomorphism:
Proposition 2.72.
Let be the Hamiltonian diffeomorphism of a smooth function which agrees near with for some Suppose that induces a bijection of Then there exists a natural transformation from to the identity functor; that is, for any there is a commutative diagram
If either or are supported on a compact Lagrangian, then these arrows are isomorphisms. If there is another such Hamiltonian diffeomorphism then
Proof.
We explain how to deduce the statement from the result of Fukaya, Oh, Ohta and Ono [23, Theorem G (G.4)] who prove that pairs of Hamiltonian diffeomorphisms induce bimodule homomorphisms of Floer groups (Akaho–Joyce explain in [9, Theorem 13.6] how to extend this to immersed Lagrangians, and it is also straightforward to include local systems). As in Lemma 2.71, the construction of the bimodule homomorphisms involves pseudo-holomorphic curve equations with respect to varying almost complex structures on the target, and we require their compatibility with the contact boundary to ensure compactness.
Define to be the image of the unit of under the bimodule homomorphism induced by the pair of Hamiltonian diffeomorphisms. We prove then that the diagram (2.60) commutes. The pair of Hamiltonian diffeomorphisms induces the bimodule homomorphism In particular, this is a left module homomorphism and there is a commutative diagram
Consider now the direct sum If these are supported on distinct Lagrangians respectively then regard as an object of supported on If they are supported on the same Lagrangian then regard it as an object of with local system where are the local systems of respectively. In either case the pair induces a bimodule homomorphism In particular, this is a right module homomorphism and there is a commutative diagram
We look at the following component, which commutes too:
This combined with (2.62) implies the commutative diagram (2.60).
We finally prove (2.61). Using (2.63) with in place of in place of and in place of we get a commutative diagram
Consider the unit of and and look at the image of this pair. Then we see that
But Fukaya, Oh, Ohta and Ono [23, Theorem G (G.4)] prove so the right-hand side of (2.65) is equal to This completes the proof. ∎
2.8 Wrapped Fukaya Categories
In general, if is an category and a set of morphisms, there exists a universal pair consisting of an category and an functor such that every element of becomes invertible in The key property that we shall use is the fact that this construction is compatible with passage to triangulated categories in the sense that there is a natural equivalence of triangulated categories
where the right-hand side is the localisation of a triangulated category.
The next result is a key computational tool for computing Fukaya categories from localisations, and also implies the above equivalence of triangulated categories in this specific context:
Proposition 2.73.
Let be an category, and an equivalence. For objects write the morphism space of the pair Let be a natural transformation from to the identity functor; and the localisation of by the set . Then for any two objects there exist natural isomorphisms
where is an inverse to Also for there exists a commutative diagram
where the horizontal arrows are products maps and the vertical isomorphisms are obtained from those of (2.67).
Proof.
Assuming that the category is triangulated, the localisation may be constructed as the categorical quotient of by the subcategory consisting of the cones of all morphisms . Morphisms in the quotient category are then computed by the direct limit of the groups as proved in [63, Lemma 7.18]. The fact that this is isomorphic to the direct limit of the groups is then a consequence of iteratively applying the functor, and the expression in terms of the direct limit of the groups is implied by the constancy of the colimit with respect to either variable. The expression for the product is an immediate consequence of functoriality. ∎
We shall also need the following observation:
Lemma 2.74.
Assume that are two sets of morphisms such that for each element , there exists a morphism such that is a composition of morphisms in . Then the natural functor is an equivalence.
Proof.
It suffices to prove that satisfies the universal property of the localisation away from , i.e. that every morphism in is already invertible in . This is immediate from the assumption because a composition of morphisms in becomes invertible in , and the only way for a product of morphisms to be invertible is if both factors are so. ∎
We now arrive at the definition of the wrapped category in the contact-type setting:
Definition 2.75.
Let be a set of Hamiltonians which are linear at the boundary in the sense that they agree with for some constant near . Assume that this set contains Hamiltonians whose slopes are arbitrarily large. Identify these with the time-one maps they generate, and suppose that if and then Then is defined as the localisation of by the set of continuation morphisms.
As a consequence of Lemma 2.74 this definition does not depend on the set of Hamiltonians which are chosen: if we add or subtract a Hamiltonian diffeomorphism to the set , then the resulting localisation does not change because of the existence of continuation maps to Hamiltonians of larger slope.
The closed Fukaya category is embedded naturally as a full subcategory of because the continuation morphisms are invertible in
2.9 Exact Lagrangians
Suppose now that the Liouville -form is given globally on so that we can speak of exact Lagrangians in and we can make the following definition.
Definition 2.76.
Denote by the full subcategory of those objects for which
the Lagrangian is embedded and exact; that is, has no self-intersection point and is an exact -form;
the local system is of the form where is the ground field of and some local system on and
is given the trivial bounding cochain, which exists because is embedded and does not bound any non-constant pseudo-holomorphic discs.
For each choose and fix a primitive function with For define the action function
by where are the primitives of respectively.
We recall the following fact.
Proposition 2.77.
Let and their primitives respectively. As in the proof of Theorem 2.43 take on a compatible almost-complex structure define the moduli space and define the evaluation map
Then the fibre of this map over a point consists of -holomorphic discs of area
Proof.
This implies that the structure of is reducible to :
Corollary 2.78.
has the structure of a -linear category whose hom space between two objects is given by
and whose structure maps are defined by the same disc counts but without the factor. The -linear extension of this category is isomorphic to that of the subcategory with the same objects, under the maps
defined as the multiplication by on the component of each ∎
Proof.
Choose now a countable set of Hamiltonians as in Definition 2.75. Denote by the localisation of with respect to this. Using (2.74) we regard as a full subcategory of
Seidel [61] makes another definition of Fukaya categories of exact Lagrangians. This is different in two respects from that definition made in §2.4. One is that Seidel defines the hom space of a single exact Lagrangian by introducing a Hamiltonian perturbation which intersects transversely, and does Lagrangian Floer theory for the pair The other is that Seidel perturbs the moduli spaces of pseudo-holomorphic curves by direct perturbations without using the Kuranishi space machinery.
We compare now the two definitions. Denote by the Fukaya category Seidel defines with the same object set as This is already a strict category because the underlying Lagrangians are exact. Denote by its localization with respect to the same Hamiltonians as we use for We prove then the following proposition, in which we consider only the cohomology categories rather than the original categories; this will do for our later treatment.
Proposition 2.79.
is equivalent to
Proof.
We define a functor We leave out the local systems to save notation. For two exact embedded Lagrangians we define a -linear map Suppose either or has no boundary. Then is defined as for some Hamiltonian diffeomorphisms with compact support in More precisely, the Floer group is defined by the direct perturbations. But these may be regarded as a different choice of virtual chains, so the resulting Floer groups will be canonically isomorphic. This is done by Fukaya, Oh, Ohta and Ono [23, Theorem G] (in the context of the present paper the similar arguments have appeared in the proof of Lemma 2.71). There is thus a canonical isomorphism There is another isomorphism from which we get an isomorphism that we want.
Suppose have both Legendrian collar. Then the morphisms of form the wrapped Floer groups and Since the isomorphism exists for every we get an isomorphism
We prove that the -linear maps thus defined are compatible with the products. Let and suppose first that at least two of have no boundary. Write and Put Recall that the isomorphism is a bimodule isomorphism with respect to and It is in particular a left module isomorphism with respect to so the diagram
commutes. The upper arrow contains the product map and the lower arrow contains the product map These are thus compatible.
Suppose now that at least two of have Legendrian collar. Then and The isomorphisms and are compatible with the product maps, completing the proof. ∎
2.10 Results in Cotangent Bundles
Let be a closed connected manifold, and a compact disc sub-bundle. Identify again the non-compact fibre over a point with the compact fibre in over the same point. Recall that the wrapped Floer group may be computed by using a result of Abbondandolo, Portaluri and Schwarz [1, Theorem 7.1]: there is a -graded -algebra isomorphism
where denotes the based loop space and is non-positively graded. We shall not need an lift of (2.75) as treated by the first author [4].
To define the Fukaya categories we do the following:
We take the background class on to be the pull-back by the projection of the Stiefel–Whitney class So the zero-section has a relative spin structure, and we give it a natural one which we shall define shortly below.
Take the -grading of to be that by the complexification of a squared real volume form on So there exist -gradings of We include in the object supported on with the trivial rank-one local system and the trivial bounding cochain.
We give the standard Liouville -form so that is an exact Lagrangian. We fix a point and include in the object supported on the fibre with the trivial relative spin structure, an arbitrary -grading, the trivial rank-one local system and the trivial bounding cochain.
Here is the natural relative spin structure on Note that the operator and the isomorphism defines a representation We often put for short.
Lemma 2.80.
The Lagrangian has a relative spin structure such that the representation on is trivial. If has characteristic this representation is automatically trivial.
Proof.
If has characteristic then defines an object of the exact Fukaya category over and the representation on makes sense over But this is automatically trivial because has only one unit.
Suppose now that has characteristic and choose at first any relative spin structure on . Then defines an object of the exact Fukaya category over (it is only here in the present paper that we use Fukaya categories over a ring). The object defines a representation and such representations are classified by elements of Since the set of relative spin structures is a torsor over the same group we can then choose a possibly different relative spin structure which makes this representation trivial. ∎
We compute the representations corresponding to perturbations of the zero-section
Lemma 2.81.
Let be a closed -form on whose graph lies in the sub-bundle Let be a filtered local system over and regard this as an object of which makes sense because and no non-constant pseudo-holomorphic curve exists for Denote by the one-dimensional representation Regard as the representation of Then is isomorphic to the representation on corresponding to the operator under the isomorphism
Proof.
Let be any element. Then there exists a corresponding element of under the isomorphism Choose a Hamiltonian diffeomorphism such that intersects transversely at in the zero-section and there exists an element of corresponding to
Denote by the fibrewise translation by Put where is the projection This induces a category isomorphism and accordingly a commutative diagram
where we have written for the morphism spaces of Write for the morphism spaces of Since has no differential it follows then that the operator
agrees with the parallel transport map of The areas of the -holomorphic discs are equal to which we explain next.
The first path from to is the diagonal from the bottom left to the top right in Figure 2.2, which lies on and the integral over this is equal to The second path from to is the vertical line in Figure 2.2, which lies on and the integral over this vanishes. The third path from to is the horizontal line in Figure 2.2, which lies on and the integral over this is equal to Thus the map
maps to So acts by upon As the functor is an isomorphism, has the same action upon which completes the proof. ∎
We need some more preparation for the proof of Theorem 1.4 (ii)–(iv). We will give a general enough treatment which applies to all the three cases. There would in fact be a simpler way if we wanted to prove only (ii); we could then focus on a single loop space whereas we shall in what follows be concerned with many path spaces The latter will be needed, however, at the end of §2.12; and the results in it will be used for the proof of (iii),(iv).
Fix an arbitrary triangulation of which contains the base-point and such that the fibre over each vertex of is an underlying exact Lagrangian in the category Denote by the space of paths from one vertex to another vertex and recall that there is a -vector space isomorphism Given three vertices there is a map defined by concatenating paths. Using the isomorphisms for we get then a -bilinear map
which in fact agrees with the product map in
Let be an object of and put a finite-dimensional -vector space. Denote by the -linear representation defined by the operator and the isomorphism Let have a sub-representation on some one-dimensional subspace
We define then an approximate local system of one-dimensional valued -vector spaces. More precisely, this is a contravariant functor from the full subcategory of the fundamental groupoid whose objects are the vertices of We define the target category by saying that its objects are one-dimensional valued -vector spaces and that its morphisms are plain vector space homomorphisms (which need not be filtered).
For each vertex the one-dimensional valued -vector space is defined as follows. Let be any connected component and define to be the image of under the operator We also write for short, which is in fact independent of the choice of For if is another component then as is associative we can write using a homotopy inverse Noting that and recalling that is a representation of we find So We have also which implies that the multiplication by is bijective and that is one-dimensional over We give the filtration induced from that of Thus is a one-dimensional valued -vector space.
Note that for any second vertex and for any morphism from to the product map induces a -vector space homomorphism This defines the desired functor because the operator (2.76) is given by concatenation of paths.
Note that the valuation of linear maps between one-dimensional -vector spaces may be defined in a natural way such that there is a functor where the latter is regarded as a groupoid of one object whose hom space is the additive group composed by the ordinary addition Composing the two functors and we get a functor
We introduce another piece of notation: if is a directed edge in from a vertex to another vertex then regard as a path and denote by its homotopy class relative to the end points. We prove now
Lemma 2.82.
In the circumstances above there exists a constant independent of such that for every directed edge of we have
Proof.
Recall that the wrapped category is defined by an localization after choosing a countable set of Hamiltonians which are linear near the boundary of the disc sub-bundle Suppose first that for any two vertices of no Reeb chords from to have integer length. Then for any its Hamiltonian chords stay in the region where the Hamiltonian is not linear, and accordingly, away from the boundary of Using this fact we can find a constant such that for any Hamiltonian for any two vertices and for any Hamiltonian chord with and we have
For each directed edge in between vertices fix a finite sum where and ’s are Hamiltonian chords from to that maps to the class under the isomorphism Let be so large that for every such and we have always which is possible because is a finite complex.
Recall now that and that is a filtered homomorphism. Then and
Replacing by its reverse edge we get also
The integer length condition on Reeb vector fields is a generic condition which may be satisfied by re-scaling the symplectic form and the Liouville form as Abouzaid–Seidel [8, last pargraph of §3.1] do. That is, consider the fibrewise re-scale of by The pull-back by this of the canonical -form on is Re-scale also by the Hamiltonians used above. Replacing above with we get then the same conclusion ∎
Given a functor we can define a -cochain by for every directed edge of We show that is a cocycle. Let be directed edges of such that the starting point of agrees with the endpoint of where Suppose that the concatenation is a -boundary of Then may be computed under the group homomorphism where is the starting point of But this may be pushed down to the homology group on which the -boundary vanishes; and accordingly, so does Define a norm on by Define also by Lemma 2.82 may then be re-stated as follows:
Corollary 2.83.
Let be as in Lemma Then ∎
We re-state the effect of fibrewise re-scales of
Corollary 2.84.
In the circumstances of Lemma 2.82 suppose that the underlying Lagrangian of lies within distance from the zero-section Then there exists independent of and such that ∎
Lemma 2.82 and Corollary 2.84 are only a few pieces of action-complete Floer theory [62, 72, 73] and local Fukaya categories [7]. In particular, the estimates above (which will do for our purpose) would be far from optimal; for instance, given a homotopy class we could try to bound by means of the least length or energy of after giving a Riemannian metric.
2.11 The Split-Generation Theorem
We recall the definitions we shall need about split-generations of categories.
Definition 2.85.
Let be a strict category with units and full subcategories. We say that is split-generated by if the image of under lies in the full subcategory split-generated by Here is defined for instance by Seidel [61, Lemma 4.7]. The full subcategory split-generated by is defined by Seidel [61, after Lemma 4.8].
We state and prove now the split-generation theorem including non-exact Lagrangians.
Theorem 2.86.
Let be a closed connected manifold, and a compact disc sub-bundle. The Fukaya category is then split-generated by the full subcategory of consisting of the single object
We begin by explaining the relevant properties of Viterbo restriction functors. We work in a slightly more general context than above. Let be a Liouville domain and a Liouville subdomain. We assume that we have fixed a background class in for the construction of Fukaya categories on , and use the restriction of this class to when discussing Fukaya categories on We assume that every underlying Lagrangian in that is not contained in intersects transversely and is equipped with a primitive function for the Liouville -form, which is constant near . Note that the embedding induces an functor
because of the maximum principle for pseudo-holomorphic discs in bounded by closed Lagrangians in In this setting, we prove the following result in Appendix A.6
Proposition 2.87.
There is a Viterbo restriction functor , extending the functor on exact Fukaya categories defined in [8].
Given the functoriality of restriction maps, we conclude:
Corollary 2.88.
If is an object of whose image in lies in the category split-generated by a collection of exact Lagrangians, then lies in the subcategory of split-generated by their images under the restriction functor. ∎
We return now to the case where is a compact disc sub-bundle in In [26] Fukaya, Seidel and Smith construct a Lefschetz fibration with total space a Liouville subdomain of (which they call ), whose critical points are of either disjoint from the zero-section or correspond to the distinct critical points of some Morse function on (see the discussion preceding [26, Lemma 7]). Taking the inverse image of a small region in the base of the Lefschetz fibration produces a Liouville subdomain containing the zero-section and which has the property that all its Lefschetz thimbles intersect transversely at respectively. We change by some exact -form on the Liouville -form on it so that there exist on the Lefschetz thimbles some primitives of the new Liouville form which are constant near respectively. This is possible because the old Liouville -form is exact on some neighbourhoods of
We use with the new Liouville -form and its primitives on the Lefschetz thimbles. Applying to them the doubling construction by Seidel [61, §18] we get
a Liouville manifold (which Seidel denotes by ) containing as a compact Liouville subdomain, and as a closed exact embedded Lagrangian;
a collection of embedded exact Lagrangian spheres (which Seidel denotes by respectively) which are given primitives that are constant near respectively, and whose intersections with are the respective Lefschetz thimbles associated to and
a symplectomorphism such that is disjoint from and which is Hamiltonian isotopic to a composition of Dehn twists around see the first paragraph in the proof of [61, Lemma 18.15].
Since we are interested in studying Fukaya categories that are possibly twisted by background classes, we need to know that any background class on extends to a background class on :
Lemma 2.89.
The restriction map admits a splitting which vanishes on every sphere.
Proof.
Seidel’s construction exhibits as a branched double cover of a Liouville domain containing as a deformation retract. A splitting is then provided by the double covering map. ∎
The data (i)–(iii) above and Lemma 2.89 imply that the following holds.
Lemma 2.90.
If is a sufficiently small Weinstein neighbourhood of , then
the intersection of each Lagrangian with is Hamiltonian isotopic to a cotangent fibre, and
the image of every closed Lagrangian in under is disjoint from .∎
Fix a background class on and the class induced in by the splitting above. Introduce the Fukaya category which contains the Lagrangians with the trivial relative spin structures (which exists because every is the double of a disc in ), the trivial rank-one local systems and the trivial bounding cochains.
Remark 2.91.
The key result about this Fukaya category is:
Lemma 2.92.
Every object of supported in lies in the category split-generated by the spheres .
Proof.
The proof is the same as that for exact Lagrangians by Seidel [61, Proposition 18.17], using Oh’s result [52, Section 9] for Seidel sequences in the non-exact case. Oh proves indeed that given two objects and a Lagrangian sphere , with corresponding Dehn twist , we have a null homotopy for the composition
where the first map is given by composition and the second by multiplication with a distinguished closed morphism in . So the cone of this morphism lies in the category generated by .
Using this sequence repeatedly we see that if is an object of supported in then there is some closed morphism from to whose cone lies in the category generated by But the Lagrangians underlying and may be made disjoint after a Hamiltonian isotopy, so is exact and the cone of is isomorphic to a direct sum of and (with shifts). This implies that is a summand of the object lying in the category generated by . ∎
We are ready now to prove the split-generation result:
Proof of Theorem 2.86.
Lemma 2.89 implies that any background class on is the restriction of a background class on . If , the Lagrangian spheres are automatically relatively spin. In the remaining case, we note that the only possible background class is represented by the Poincaré dual of the zero section, which corresponds to a cotangent fibre; and its pull-back to is represented by the inverse image of a cotangent fibre, whose intersection with vanishes with coefficients because the submanifolds intersect an even number of times. The result then follows by applying Lemma 2.92 and Corollary 2.88, using as well the fact that all cotangent fibres are equivalent in the wrapped Fukaya category, so that one suffices. ∎
2.12 Yoneda Functors
We recall now some standard facts about Yoneda functors. We follow Seidel [61] with minor modifications. For a strict category with units denote by the category of left modules (Seidel [61, (2f)] works with right modules in place of left modules, and with cohomology units in place of strict units). Seidel [61, (1l) and (2g)] defines the Yoneda functor which is a strict functor, and proves that this induces a fully faithful functor between the cohomology categories [61, Corollary 2.13].
Seidel [61, (2.13)] proves also that if are strict categories and a cohomologically fully faithful functor then the pull-back functor and the Yoneda functors fit into the diagram
which commute after passing to the cohomology categories.
Lemma 2.93.
Let be a strict category with units. Let be full subcategories such that is split-generated by Consider the composite of the inclusion functor the Yoneda functor and the pull-back functor Then the induced functor is fully faithful.
Proof.
We use a result of Seidel [61, Lemma 4.7] and its proof. Let be a split-closure of the triangulated envelop of as Seidel defines [61, after Lemma 4.8]. Denote by the full subcategory split-generated by Then the composite functor has image in Using (2.80) twice we get now a commutative diagram
Here is induced from the Yoneda functor and accordingly fully faithful. As (2.81) commutes the composite factoring through and is fully faithful. But is split-generated by so is fully faithful too. The arrow is induced from the Yoneda functor and therefore fully faithful too. The composite is thus fully faithful. ∎
We return now to the circumstances of Theorem 2.86. Denote by the endomorphism algebra of the single object Lemma 2.93 implies then
Corollary 2.94.
The induced functor is fully faithful. ∎
We show in Proposition 2.95 below that this functor is compatible in a certain sense with a finite cover Denote by the same the induced map between the cotangent bundles, which is also a finite cover. Fix a point Identify again the non-compact fibre with the compact fibre in over the same point
We define Fukaya categories of the pre-image We pull back to the branes on or more precisely, the pull-back of an underlying Lagrangian in is possibly disconnected, in which case we include all the connected components of it; for instance, we include the fibre We pull back to the relative spin structure on with which we introduce the Fukaya category
We define then an algebra homomorphism which acts as the set-theoretic push-forward upon the underlying Lagrangians. Choose a compatible almost complex structure on and pull this back to Recall that are made from gapped categories, We define gapped algebra homomorphisms between these which respect all the operators, simply by pushing forward the pseudo-holomorphic discs. We define in the same way an inverse system of homotopy equivalences with which we can pass to the limits. Then we get a gapped algebra homomorphism between the curved complexes Including the bounding cochains we get a strict algebra homomorphism that we want.
We pass now to the wrapped categories. By definition there exist a localization functor and accordingly a composite functor Restricting this to the single object we get an algebra homomorphism We pull back to the set of Hamiltonians used to define with which we define the wrapped category We show that the algebra homomorphism descends to the hom space in the localized category Let be the time-one diffeomorphism of a Hamiltonian we have chosen, which is the pull-back of some Hamiltonian Denote by the same the time-one diffeomorphism Then the push-forward of is equal to and the push-forward of the continuation morphism is the continuation morphism The latter is already invertible in and hence we get by the universal property an algebra homomorphism that we want.
We define now a pull-back functor between the Fukaya categories whose underlying Lagrangians are compact without boundary. Take with a Lagrangian, a local system and a bounding cochain. We define a gapped homomorphism
that assigns to each singular chain in the sum of all possible inverse images in Using the fact that discs are simply connected, we see that the moduli space of discs in is a cover of the moduli space of discs in and hence that a choice of virtual chains for determines the corresponding choice for by passing to the cover. This implies that (2.82) is compatible with the gapped algebra structures on the singular chain complexes The same argument shows that the homotopies required to construct the inverse system are also compatible with the pull-back by So we can take the limit with respect to and (2.82) is a gapped homomorphism. This maps the bounding cochain to a bounding cochain in which defines an functor that we want.
Note that the push-forward algebra homomorphism induces a pull-back functor Hence we get a diagram
where the vertical arrows are the Yoneda functors. We prove
Proposition 2.95.
The diagram (2.83) commutes in the sense that for every object the two Yoneda modules and are isomorphic modules over
Proof.
This follows from the fact that the relevant holomorphic curve moduli spaces are bijective under the covering of the moduli spaces used above; the underlying Lagrangians in include which specifies the lift of each disc in ∎
Finally we extend Lemma 2.82 to the finite cover of Suppose therefore that we are given again a triangulation of and pull back this to a triangulation of which we denote by Suppose given an object and using the same notation as in Lemma 2.82, put and denote by the representation given by the products. Suppose that the restriction of to the subgroup has a sub-representation on some one-dimensional subspace Then we get a functor as in Lemma 2.82 with in place of and with in place of The estimate in Lemma 2.82 holds uniformly with respect to that is, the following holds.
Lemma 2.96.
There exists a constant independent of independent of and such that
Proof.
The estimate is uniform with respect to because the Hamiltonians may be chosen first in We pull them back to so as to follow the proof of Lemma 2.82. Then for every directed edge in between vertices the isomorphism maps the class to the finite sum where we can use the same ’s as in the proof of Lemma 2.82. We can use accordingly the same constant as in the proof of Lemma 2.82, which implies the estimate above we want. ∎
The re-scaling argument in Corollary 2.84 implies
Corollary 2.97.
In the circumstances of Lemma if the underlying Lagrangian of lies within distance from the zero-section then we have ∎
3 Proof of Theorem 1.4
We begin in §3.1 with the proof of Theorem 1.4 (i). In §3.2 we prove the key lemma for Theorem 1.4 (ii). For this we do not need any hypothesis on In §3.3 we prove Theorem 1.4 (ii) when is virtually abelian. In §3.4 we prove a few facts we shall need to deal with finite covers of In §3.5 we prove Theorem 1.4 (iii),(iv).
3.1 Proof of Theorem 1.4 (i)
Let be a Calabi–Yau manifold, a closed connected embedded special Lagrangian, and a Weinstein neighbourhood of in the sense of Definition 1.2. We begin by proving
Proposition 3.1.
If the fundamental group is trivial then the Fukaya category is generated by the single object the zero-section.
Proof.
Let be an object of and consider its image under the Yoneda functor Applying the homological transfer lemma, we may replace by an algebra whose underlying graded vector space is given by its cohomology, which is isomorphic to and hence is supported in non-positive degrees. It follows (by the degree-filtration argument [5, Appendix A]) that the Yoneda image of is a repeated extension of ungraded modules over the degree algebra But is isomorphic to so the image of is a repeated extension of ungraded -vector spaces. As is supported on a compact Lagrangian without boundary, the number of these extensions is finite and each ungraded piece is finite-dimensional over Recall now that intersects transversely and exactly at one point and that is one-dimensional over Then each ungraded piece of the Yoneda image of may be written as for some finite-dimensional -vector space As the functor is fully faithful, the object is isomorphic in the cohomology Fukaya category to a repeated extension of objects of the form This completes the proof. ∎
We use an algebraic lemma by the first author [5, Lemma A.4] who proves the first general results on Arnold’s nearby Lagrangian problem [27]. The original statement is over but the proof shows that it applies to any field including :
Lemma 3.2.
Let be a -graded -linear triangulated category generated by one object whose endomorphism algebra is supported in non-negative degrees. Then every object with supported in non-negative degrees is isomorphic in to for some integer and for some finite-dimensional -vector space ∎
We complete now the proof of Theorem 1.4 (i).
Proof of Theorem 1.4 (i).
Suppose first that is trivial. Let be supported on a closed special Lagrangian Then is a non-zero object and by Proposition 2.55 the graded group is supported in non-negative degrees. Proposition 3.1 and Lemma 3.2 imply then that is isomorphic in to some with the trivial bounding cochain. Since is non-zero it follows that is a non-zero vector space. Note also that there are vector space isomorphisms
which is non-zero (by Corollary 2.64 or because is isomorphic to the ordinary cohomology group). Hence it follows by Proposition 2.67 (ii) that
If is not trivial but finite, taking the universal cover we see that the pre-image agrees with So as we have to prove. ∎
3.2 The Key Lemma
We prove the single degree property of nearby special Lagrangians.
Lemma 3.3.
Let be a Calabi–Yau manifold of complex dimension a closed connected embedded special Lagrangian, and a Weinstein neighbourhood of Fix a point and a -grading of with phase at Then there exists a neighbourhood of such that for every object whose underlying Lagrangian is special of phase and contained in the graded vector space is supported in degree
Remark 3.4.
Here the speciality is essential because otherwise may be supported in two or more degrees; see Example 5.11.
There is in fact an analytical theorem for those nearby special Lagrangians in a given homology class as in Theorem 1.4 (iii),(iv). These are close to the zero-section as varifolds in the sense of geometric measure theory, so we can use Almgren’s theorem (for area-minimizing currents). As explained by De Lellis and Spadaro [15, Theorem 2.4] there exists a small set such that outside the fibre over the nearby special Lagrangians ’s are close to the zero-section This implies that the Maslov index of the pair is always zero, so that the chain complex is already supported in degree
Proof of Lemma 3.3.
Suppose first that the Calabi–Yau structure near is real analytic. For each we define a compactly-supported Hamiltonian perturbation of which is special Lagrangian near To start with, take a -linear isomorphism which maps to to the symplectic form to and the holomorphic volume form to where are the coordinates of Take and with so that the plane is special Lagrangian with respect to the flat Calabi–Yau structure. Take a small ball in about and embed it into near mapping to Denote the image of the ball by the same for short. Take a neighbourhood of in over which is a product bundle. We can suppose that the ball lies in the product bundle Denote by the vertical projection of the product bundle. Put which is a ball in about the origin. Choose a neighbourhood of which is so small that here means that the closure of is contained in
We perturb into a special Lagrangian with respect to the Calabi–Yau structure of There are two ways of doing this. The first is to use Cartan–Kähler’s theorem as Harvey and Lawson do [31, Chapter III, Proof of Theorem 5.5], which we can do because we have supposed that our Calabi–Yau structure is real analytic. The second is to use the implicit function theorem again as Harvey and Lawson do [31, Chapter III, Proof of Corollary 2.14]. In either case we get a special Lagrangian near We can suppose that this is a graph over Using a cut-off function we get then a compactly-supported Hamiltonian perturbation of such that is a special Lagrangian in
Grade the Lagrangians and by phase and the Lagrangians and by phase at For define an affine function by For define an affine function by Using these we get a Hamiltonian isotopy from to the plane which does not cross As there is a graded Hamiltonian isotopy from to This induces a -graded Hamiltonian isotopy from to Suppose now that is supported on a closed special Lagrangian Since it follows then, by the unique continuation principle, that the special Lagrangian part of does not agree with We can thus apply Proposition 2.69 to and consequently is supported in degrees So is supported in degrees Noting that this holds for every and that it follows that is supported in degree
Finally, if the Kähler form near is not real analytic we can approximate it by a real analytic one (after taking a Kähler potential) and find a compactly-supported Hamiltonian diffeomorphism such that is real analytic. Applying to this the result above, we get a corresponding neighbourhood of Put Let be supported on a closed special Lagrangian (relative to ) and denote by the push-forward. Then is supported in the single degree But as is compactly supported, we have which completes the proof. ∎
The effect of the lemma above is that we do not have to deal with the extension problem which we explain now. Let be an object of As in the proof of Theorem 1.4 (i) the Yoneda module is a repeated extension of ungraded modules over that is, representations of These extensions may be non-trivial. In Example 5.11, for instance, the direct sum of and the graph of are different from their surgery. More precisely, these define non-isomorphic objects of They are made up of the same representations but with different extensions.
3.3 Proof of Theorem 1.4 (ii)
Let be a Calabi–Yau manifold of complex dimension complex structure Kähler form and nowhere-vanishing holomorphic form Let be a closed connected embedded special Lagrangian with virtually abelian, and a Weinstein neighbourhood of
We make several definitions we will use for a version [38, Proposition 2.13] of McLean’s theorem. Denote by the metric on induced from Define a function by
Denote by the linear operator where is computed with respect to Then is an elliptic operator and we have where the direct sums are orthogonal with respect to This implies that the de Rham cohomology group is isomorphic to
We fix a triangulation of and use the norm on introduced before Corollary 2.83. The de Rham theorem implies then that there exists such that every -cocycle may be represented on by some -form with where the pointwise norm is computed with respect to The McLean theorem [38, Proposition 2.13] implies in turn that there exists with the following property: for every with the graph of is Hamiltonian equivalent to some special Lagrangian in Let be as in Corollary 2.84. Denote by the disc sub-bundle of radius which is so small that we can use Lemma 3.3.
Let be a closed special Lagrangian with unobstructed and contained in We use an algebraically closed Novikov field to define the Fukaya category Let be an object supported on Corollary 2.56 implies then that is a non-zero object. By Lemma 3.3 the graded vector space is supported in a single degree. Denote by the -linear representation made from the products in
Suppose first that is abelian so that has a one-dimensional sub-representation Recall that and denote by the closed -form on which represents the de Rham class and satisfies the equation Using the notation of Lemma 2.81 write where is some group homomorphism This defines a rank-one local system over As in Remark 2.11 the category of rank-one free modules is equivalent to that of one-dimensional valued vector spaces. Hence we get a filtered local system over which we denote by the same Lemma 2.81 implies then that is isomorphic to the representation on As is a sub-representation of there exists a degree-zero non-zero morphism from to Thus
By Corollary 2.84 we have We have chosen so that Then the graph is Hamiltonian equivalent to some special Lagrangian As this Hamiltonian perturbation does not change the isomorphism class of it follows from Proposition 2.67 (i) that
For virtually abelian we have only to pass to a finite cover of which has abelian fundamental group. This completes the proof. ∎
3.4 Finite Covers and McLean’s Theorem
We begin by proving a version of the de Rham theorem which is uniform with respect to finite covers.
Proposition 3.5.
Let be a closed connected Riemannian manifold, a triangulation of and a finite cover. Let be a -cocycle of the simplicial complex and put Then for every integer there exists independent of the finite cover and so large that the de Rham class may be represented on by some closed -form with norm with respect to the induced metric
Proof.
We use the -linear map from the -vector space of -cochains to that of -forms, defined by Singer and Thorpe [65, §6.2, Lemma 1], which maps -cocycles to closed -forms and induces the de Rham isomorphism of the two cohomology groups. Part (4) of their lemma shows that for each edge in the corresponding -form is supported near in The process of assigning the -form is thus local, which we can do essentially in before passing to the finite cover More precisely, denote by the dual basis to Write The -form corresponding to is then bounded uniformly with respect to On the other hand, we have for every These imply the estimate we want. ∎
We prove then that McLean’s theorem holds uniformly with respect to finite covers.
Proposition 3.6.
Let be a Calabi–Yau manifold of complex dimension Let be a closed connected embedded special Lagrangian, and a Weinstein neighbourhood of Define a function by (3.1) using the Kähler form and the nowhere-vanishing holomorphic form of Then for every there exists such that the following holds: let be a finite cover of degree and a closed -form on with and both computed with respect to the induced metric then there exists a function such that the graph of is a closed embedded special Lagrangian in the pre-image whose Calabi–Yau structure is induced by
Proof.
For an integer denote by the Sobolev space of -forms on and by that of -forms on Denote by the co-dimension one closed vector subspace consisting of with where is the volume form of Denote by the modified Laplacian We often regard as a multiplication operator and write We prove
Lemma 3.7.
There exists depending on but independent of the finite cover and so large that for every we have
Proof.
Gallot [28, 3.1] proves the following: let be a compact Riemannian manifold of dimension and of Ricci curvature then there exists a constant which depends only on and is such that for every with we have
where the Sobolev norms and are computed with respect to We apply this to the pull-back by As the volume and diameter of may be estimated in terms of those of combined with the there exists a constant independent of such that for every with we have
Suppose now that is any and put Then
where the products and the Sobolev norms are computed with respect to Note that is independent of The estimate (3.3) implies then
Since the volume of is bounded above by a constant depending only on and it follows that so is Hölder’s inequality, on the other hand, implies This combined with (3.4) completes the proof. ∎
Remark 3.8.
If is a finite cover of then the factor tends to infinity as the covering degree increases.
Fix now so that we have the Sobolev embedding We prove that the Sobolev embedding and the elliptic regularity are -uniform too.
Lemma 3.9.
There exists independent of the finite cover and so large that for every we have
Proof.
Let be less than half the injectivity radius of the Riemannian manifold For each point denote by the open geodesic ball of radius about As is compact there exists a finite set such that By the Sobolev embedding on each there exists independent of and so large that for every we have Note that is also independent of We look now at the finite cover Take any with and denote by the lift of with respect to which exists because is simply connected. Then for any we have
which is bounded above by Since is covered by ’s with it follows by (3.6) that proving the left inequality of (3.5).
We turn now to the other part. We show first that for with the non-empty intersection is necessarily connected. If there were two components, there would be two distinct geodesics connecting these two. But have been chosen to have the common radius which is less than half the injectivity radius. So lies in a single geodesic ball, which contradicts that the two geodesic exist. Thus is connected. The union is accordingly simply connected.
As is covered by ’s of radius there exists so close to that if we denote by the concentric ball of radius then still covers Then for we have
For and denote by the set of such that and the simply connected set lifts to Exchanging the order of the two sums in (3.7) we see then that
Note that there is an injection defined by so that the number is bounded above by some -independent constant It follows then from (3.8) that
On the other hand, by the ordinary a priori estimates over there exists independent of and so large that for every and we have
Hence it follows by (3.9) that
Thus will do for the right part of (3.5). ∎
Remark 3.10.
For the proof above we have not used the degree bound as we did after (3.2). This may be justified by recalling that the latter (3.2) is a global property of Riemannian manifolds. The current Sobolev inequality, which is of the form would hardly be such a global property. Interior estimates about elliptic regularity are also of local nature.
The definition of implies that is surjective, so there exists a right inverse to this, which we denote by We prove then that the following holds.
Corollary 3.11.
There exists independent of the finite cover and so large that for every we have
Proof.
We complete now the proof of Proposition 3.6. We follow the standard proof of McLean’s theorem [38, Proposition 2.13] using the uniform estimates above. Denote by the nowhere-vanishing holomorphic form of Let be so small that if is a -form on with then the graph of lies in the pre-image Put The Sobolev inequality (3.5) implies then that for every the graph of lies in the pre-image and we can therefore define on the -form Define an operator by for We have because depends only on the homology class whose pairing with vanishes because is a special Lagrangian. So is well defined.
Standard computation shows that is differentiable at with respect to the Sobolev norms. By hypothesis we have and hence we see that the differential agrees with Corollary 3.11 implies then that for with we have so We solve the equation for with Define by We seek then a fixed point of the composite and use therefore the contraction mapping principle. As may be computed pointwise on there exists independent of and so large that the following holds: let satisfy at every point of and put then
at every point of Integrating this and using the Cauchy–Schwarz inequality we get
Putting and we have
This with (3.10) implies
Hence choosing and using Corollary 3.11 we see that is a contraction map from the closed ball to itself. So there exists a fixed point of the operator and Thus Now satisfies the elliptic system and is therefore The graph of is thus the special Lagrangian we want. ∎
Here is a corollary of Proposition 3.5 and the estimates in the proof above.
Corollary 3.12.
Fix a constant and suppose that in the circumstances of Proposition 3.5 the finite cover has degree Then there exists independent of and so large that the de Rham class may be represented on by some closed -form with norm and where the norm and are computed with respect to the induced metric
Proof.
Fix an integer Then by Proposition 3.5 there exist a constant independent of and a closed smooth -form on with Define a -form on by Then and We estimate the norm of As is again uniformly bounded, there exist constants independent of and so large that
By (3.5) there exists a constant independent of and so large that and
Thus which completes the proof. ∎
3.5 Proof of Theorem 1.4 (iii),(iv)
We prove Theorem 1.4 (iii) now. For of dimension the hypothesis that should be virtually solvable implies that virtually abelian. So we can use the result of (ii).
Suppose therefore that has dimension so that we can use the results of 3.4. Suppose also that is solvable at first. We begin by choosing a disc sub-bundle We impose two conditions. Firstly, should be so small that we can use Lemma 3.3. Secondly, let be as in Lemma 2.82 or Corollary 2.97; as in Corollary 3.12; and as in Proposition 3.6. Then the radius of should be
Let be a closed special Lagrangian with unobstructed and contained in We use again an algebraically closed Novikov field to define the Fukaya category Let be an object supported on Then is a non-zero object. By Lemma 3.3 the graded vector space is supported in a single degree. Denote by the representation made from the products in By hypothesis the group is solvable; and accordingly, so is its image We apply to this a version with estimate of the Lie–Kolchin theorem.
Theorem 3.13 (Lie–Kolchin).
There exists a function such that for and for an algebraically closed field (which need not be the Novikov field) every solvable subgroup of has a finite-index triangularizable subgroup of index
By hypothesis the dimension of is bounded above by a given constant, say Consequently there exists a finite-index subgroup of of index and on which is triangularisable. Let be the corresponding finite cover, and take a one-dimensional sub-representation of the triangularisable group.
Recall that and denote by the closed -form on that represents the cohomology class and satisfies the equation Denote by the graph of Using the notation of Lemma 2.81 write where is some group homomorphism This defines a rank-one local system over Hence we get a filtered local system over which we denote by the same Lemma 2.81 implies then that is isomorphic to the representation on Denote by the restriction to of As is a sub-representation of there exists a degree-zero non-zero morphism from to Hence it follows that there exists a degree-zero non-zero morphism from to Thus
Choose now a triangulation of so that we can define the functor By Corollary 2.97 we have Lift to a -cocycle of with coefficients. Using the notation of Proposition 3.5 we have then We can therefore apply Corollary 3.12 to the cohomology class which may now be represented on by some closed -form with and Proposition 3.6 implies then that the graph of is Hamiltonian equivalent to some special Lagrangian As this Hamiltonian perturbation does not change the isomorphism class of it follows from Proposition 2.67 (i) that is an irreducible component of Thus is unbranched as we want to prove.
In the general case in which is virtually solvable, all we have to do is to pass to a finite cover of whose fundamental group is solvable. This completes the proof of Theorem 1.4 (iii).
We turn now to the proof of Theorem 1.4 (iv). We follow the proof above till we use Theorem 3.13. We give then a new argument. By hypothesis the group has no non-abelian free subgroup; and accordingly, its image has no such subgroups either. We apply to this the following version of Tits’ theorem [69, Corollary 1].
Theorem 3.14.
For and for any field (of any characteristic) every finitely generated subgroup of that has non-abelian free subgroups has a finite-index solvable subgroup.
Remark 3.15.
This will be the most convenient to us because we are concerned only with finitely generated groups. The paper [69] deals with those linear groups which need not be finitely generated.
Theorem 3.14 and a few other known facts imply
Corollary 3.16.
There exists a function such that for and for a field of characteristic every finitely generated subgroup of that has non-abelian free subgroups has a finite-index solvable subgroup of index
Proof.
We consult Wehrfritz [74, 75] without tracing back the original papers of older results. It is known [75, Corollary 3.8] that every linear group has a unique normal solvable subgroup Let be finitely generated as above. It is then known [74, Lemma 2] that the quotient group is a finite subgroup of for some This comes from other results [75, Theorems 6.3 and 6.4] which show indeed that There is an old theorem of Jordan [75, Theorem 9.2] which says that every finite subgroup of has a finite-index abelian subgroup of index where depends only on So there exists a finite-index abelian subgroup of index where depends only on Denote by the pull-back of by the projection There is then an exact sequence which implies that is solvable by abelian and accordingly solvable. Also is a finite-index subgroup of index less than that of which is as we have to prove. ∎
Recall from hypothesis that the dimension of is bounded above by a given constant Corollary 3.16 and Theorem 3.13 imply then that there exists a finite-index subgroup of on which the representation is triangularisable and such that the index of the subgroup is bounded above by a constant depending only upon The rest of the proof is the same as that for Theorem 1.4 (iii). We have thus proved Theorem 1.4 (iv). ∎
3.6 Versions of Theorem 1.4 (iii), (iv)
We discuss first the difference between (iii) and (iv) in Theorem 1.4. In the proof of (iv), if we did not suppose that has characteristic then there would be a problem in the proof of Corollary 3.16. Jordan’s theorem in fact fails for of characteristic Take for instance to be the algebraic closure of the finite field and take There is then a sequence of subgroups of and as is well known, their normal subgroups will lie in the centre of except for or So the indices of them are nearly the order of which tend to infinity as increases.
We make a positive characteristic version of Theorem 1.4 (iv), supposing that there exist infinite finitely presented torsion groups, although no such examples are known. We need these to be finitely presentable because they are with a compact manifold.
Proposition 3.17.
Let be as in Theorem 1.4, and a prime number. Let be an infinite torsion group with no elements of order Then for every there exists a neighbourhood of such that the following holds: let be a closed irreducibly immersed special Lagrangian of degree contained in and which has unobstructed with respect to some filtered local system of rank over some Novikov field of characteristic then is unbranched.
Proof.
Recall that Theorem 3.14 holds in characteristic We make a characteristic- version of Corollary 3.16: there exists a function such that for for a prime and for a field of characteristic every subgroup of that has non-abelain free subgroups and has no elements of order has a finite-index solvable subgroup of index As we have mentioned above, we have only to be careful about the last step to the proof of Corollary 3.16. We replace Jordan’s theorem [75, Theorem 9.2] by Schur’s theorem [75, Corollary 9.4]: every finite subgroup of with no elements of order has a finite-index abelian subgroup of index where depends only on Once this has been done we can follow the proof of Theorem 1.4 (iv) to show that Proposition 3.17 holds. ∎
We make another version of Theorem 1.4 (iii),(iv) which is closer to Theorem 1.4 (ii), not referring to the constant but making again a strong hypothesis on
Proposition 3.18.
Let be as in Theorem 1.4. Let have no non-abelian free subgroups and satisfy the property Then there exists a neighbourhood of such that there exists a neighbourhood of such that every closed irreducibly immersed special Lagrangian with unobstructed and contained in is unbranched.
Proof.
The proof is in fact easier. The property implies indeed a uniform eigenvalue estimate [48, Proposition 2.5] for the Laplacians over finite covers of More precisely, as the linear operator is a modified Laplacian we shall need the following computation: in the circumstances of Proposition 3.6 introduce a conformal change of the Riemannian metric on then where is the ordinary Laplacian of Proposition 3.6 will then hold automatically; and moreover, we shall not need the index estimates as in the proof of Theorem 1.4 (iii),(iv). ∎
There is however a problem: no known examples of strictly infinite, finitely presentable groups satisfy the hypotheses above. Here we are interested in strictly infinite groups because of Theorem 1.4 (i), in which we have proved already a stronger result for finite.
At least -manifold groups do not satisfy all the conditions above. Every -manifold group is always residually finite; and as we shall see in Lemma 5.19 those -manifold groups which have no non-abelian free subgroups are amenable. These two contradict the property [48, Proposition 1.32] (see also Brooks [13] and Sunada [67]).
There are examples [37, 46, 55] of infinite, finitely presentable non-amenable groups without non-abelian free subgroups (counterexamples to von Neumann’s problem) which however do not satisfy the property (because they have surjections onto ). There are, on the other hand, infinite torsion groups [18] with the stronger property T, called Golod–Shaferevich groups, and simple groups [54] with the property T which has no non-abelian free groups; but these would hardly be finitely presentable.
3.7 Bad Immersions
Theorem 1.4 holds for some badly immersed special Lagrangians. In the known construction of Fukaya categories, the underlying Lagrangians are required at least to be cleanly immersed [22]. On the other hand, in our nearby problem we want to include as many (special) Lagrangians as possible, so we introduce the following notion [36, before Corollary 4.6].
Let be a closed manifolds and be a Lagrangian immersion. Put We say that an object in is supported near if there exist arbitrarily-small Lagrangian perturbations of which underlie some objects in the same isomorphism class in More concretely, this means that for every there exists a perturbation of the immersion with norm bounded by for all , and which supports an object representing the given isomorphism class
Theorem 1.4 holds for a special Lagrangian which admits such an isomorphism class as above. The reason is that the results of the paper [36] are stated and proved for objects of supported near a special Lagrangian. In particular, Proposition 2.67 and Lemma 3.3 are still valid for these objects. So the proof of Theorem 1.4 will work.
But the condition above on is rather strong. Here is one possible way of verifying it. Suppose given a smooth family such that and the other are all generic and mutually isotopic under domain Hamiltonians on in the sense of Akaho–Joyce [9, Definition 13.14 (ii)] (this is a family version of the classical transversality condition in differential topology). If we could lift such an isotopy to an ambient Hamiltonian isotopy, then it would follow that any isomorphism class in supported on the immersion is also supported on for all generalising the Hamiltonian invariance theoreom by Fukaya, Oh, Ohta and Ono [23]. In general, such a lift does not exist because then those strips which bound arcs between self-intersection points will have constant area under the isotopy. We can perhaps use the methods of [56] to prove invariance statements for this perturbation problem.
4 Tsai–Wang’s Method
We recall now Tsai–Wang’s method and explain then the relevant examples of nearby special Lagrangians. In this section we do not use Floer cohomology groups or Fukaya categories.
We begin by recalling the key formula from geometric measure theory. Let be a Riemannian manifold. Given a vector field on a point , and an -vector subspace we denote by the trace of the map where the two arrows are the covariant derivative and the orthogonal projection In particular, if is the gradient vector field of a smooth function then We use also the notion of rectifiable integral varifolds in which we treat, following Simon [64], as Radon measures on Recall that if is a compactly supported stationary integral varifold in then for every vector field on we have In particular,
for every smooth function
We recall now the following result from the literature [47, 70, 71]. We prove this to show how to use (4.1).
Proposition 4.1.
Let be a Ricci-flat Calabi–Yau manifold of complex dimension and a closed embedded special Lagrangian which has positive Ricci curvature with respect to the induced metric from Then there exists a neighbourhood of in which every compactly supported stationary integral -varifold is supported on
Proof.
Tsai and Wang prove that is strongly stable in their sense [71, Proposition A(iii)] and that the following holds [71, Proposition 4.1]: there exists an open neighbourhood of such that for every point and for every -dimensional plane we have at
where is the distance from and depends only upon Applying (4.1) with and using (4.2) we find on the support of the varifold completing the proof. ∎
Remark 4.2.
The idea of the proof is that the negative gradient flow of decreases the area of The computation above is more convenient for the rigorous treatment.
As is well known, closed special Lagrangian currents are area-minimizing and in particular area-stationary [31]. Combining this with Proposition 4.1 we see immediately that the following holds:
Corollary 4.3.
In the circumstances of Proposition 4.1 every compactly supported closed special Lagrangian current in is necessarily non-singular, of phase and supported on ∎
Example 4.4.
We turn now to the example which shows that strictly abelian is essential to the last part of Theorem 1.4 (ii). We begin by proving a lemma in slightly more general circumstances.
Lemma 4.5.
Let be integers. Take and give these the flat metrics. Then every compactly supported stationary integral -varifold in is parallel to that is, supported on for some finite set
Proof.
Let be a compactly supported stationary integral -varifold in Applying (4.1) again with we get But is pointwise non-negative and degenerates only in the direction of so -almost everywhere in we have and is parallel to
We recall a lemma of Simon [64, Lemma 19.5]. Let be a point on the support of and we work near this so we can regard as a point of the universal cover of Let be so small that for every point of distance from we can define Fix and let be a point on the support of of distance from and such that is not parallel to Simon [64, Lemma 19.5] proves then that
where is the Euclidean volume of the unit ball in the density of and the ball of radius about The original formula has in fact another term on the right-hand side, which contains the integral over of where are the projections of onto respectively. But is parallel to so the additional term vanishes and (4.3) holds.
Remark 4.6.
The hypothesis of Proposition 4.1 is stable under small perturbations of the Ricci-flat Kähler metric (just because positivity is an open condition). On the other hand, Lemma 4.5 will fail under perturbations of the ambient metric. For instance, Kapouleas and Yang [41] construct a sequence of minimal surfaces converging to the flat Clifford torus with multiplicity two; and a suitable rescaling of the ambient metric converges to the flat metric on the tubular neighbourhood of
Corollary 4.7.
Let be a flat Calabi–Yau manifold, a closed embedded totally geodesic special Lagrangian, and a tubular neighbourhood of Then every compactly-supported closed special Lagrangian current in is parallel to that is, a locally constant multi-valued graph over Every such graph is a disjoint union of single-valued graphs if and only if is abelian.
Proof.
Since is a closed flat manifold we get according to Bieberbach a finite Galois cover such that is a maximal abelian subgroup of Denote again by the same the finite cover between the cotangent bundles. Identify with a neighbourhood of the zero-section and the pull-back with a neighbourhood of As is a product bundle we can use Lemma 4.5 to show that the pull-back of a compactly supported closed special Lagrangian current is parallel; and accordingly, so is the original Now if is abelian, this is the maximal abelian subgroup of itself and so every special Lagrangian is a disjoint union of single-valued graphs. If is non-abelian, there exist a non-trivial covering transformation and accordingly an element which acts non-trivially upon So there exists a constant -form on that does change under By rescaling we can suppose that is so small that we can define its graph in The push-forward of this defines then a truly multi-valued graph in ∎
Example 4.8.
Let with a Klein bottle. Write as the mapping torus of the automorphism given by modulo for Then there are corresponding double covers and under which the graph of is pushed forward to a multi-valued graph. Strictly speaking, is not Calabi–Yau and is not even orientable. But is still a special Lagrangian in the extended sense as in Definition 2.32; see also Remark 2.33. The multi-valued graph is also a special Lagrangian in the same sense. It is easy to find multi-valued graphs which are ordinary special Lagrangians. For instance, take to be the mapping torus of the orientation-preserving diffeomorphism given by where then the graph of is pushed forward to a multi-valued graph which is an ordinary special Lagrangian.
5 Branched Examples
Here are examples of branched nearby closed Lagrangians, which will all be Lagrangian surgeries. We begin by recalling the result of Fukaya, Oh, Ohta and Ono [23, Chapter 10] which describes the effect of applying Lagrangian surgeries to pairs of objects in the Fukaya category.
Theorem 5.1 (cf. [23, Chapter 10]).
Let be a compact -graded symplectic manifold with contact boundary and two closed embedded -graded Lagrangians which intersect each other transversely at least at one point. Let the Maslov index function be identically equal to so that we can define a -graded Lagrangian surgery [59, Lemma 2.14]. Then for any supported on respectively, there exist an object supported on and for every other a long exact sequence
where the connecting homomorphism is the right multiplication by some element of ∎
Remark 5.2.
We allow to have two or more intersection points, which is possible as Fukaya, Oh, Ohta and Ono mention [23, after Theorem 55.7] although they deal only with single intersection points in their formal treatment. So there are many objects of the form above according to different scales of surgeries at different intersection points.
Here are remarks on the proof. Fukaya, Oh, Ohta and Ono use the cobordism method for the moduli spaces of holomorphic discs with boundary on the surgery; the cobordism is associated to stretching the neck (degenerating the almost complex structure) near the intersection points. The method of Biran–Cornea [20] provides an easier proof which does not rely on a degeneration argument: one can associate to the surgery a Lagrangian cobordism in , and the count of holomorphic triangles with boundary on this cobordism yields the desired long exact sequence. Although their results are stated for monotone or exact Lagrangians, these assumptions are imposed only to avoid using the theory of virtual chains in case the Lagrangians bound holomorphic discs. We shall in practice take to be graphical Lagrangians in a cotangent bundle, which do not bound any non-constant holomorphic discs; so their proof can be implemented with no difficulty.
For special Lagrangians we have
Corollary 5.3.
In the circumstances of Theorem 5.1 let be a Calabi–Yau manifold and both special Lagrangians. Then is not isomorphic in to any other object supported on a closed special Lagrangian of any phase.
Proof.
By shifting we may suppose that has phase for some and that has phase Then by (2.30) we have so Applying (5.1) to in place of we get two exact sequences
On the other hand and are both non-zero, so and are both non-zero. Suppose contrary to our conclusion that is isomorphic to an object supported on a closed special Lagrangian of phase By perturbing we may suppose that is supported on a nearly special Lagrangian which intersect and generically. Since we get an intersection point with Maslov index Hence denoting by the phase function and using (2.30) we find and in particular Also since we get an intersection point with and in particular But both and are near so that is, and have all phase Since and are both non-zero it follows by Proposition 2.67 (i) that these three Lagrangians have a common irreducible component. But this is impossible because intersect each other transversely at least at one point. ∎
We recall now a key fact about Morse -forms; a Morse -form on a manifold is a closed -form on such that every point of has a neighbourhood on which for some Morse function
Theorem 5.4 (Honda [35]).
Let be a closed manifold of dimension and a Morse -form on with no critical point of index or Then there exists on a Riemannian metric and in the de Rham class a -harmonic Morse -form which has the same distribution of indices as has; that is, for each the number of index- critical points of is equal to that of ∎
Remark 5.5.
If is equipped with any real analytic structure, we can suppose that and above are real analytic. To see this, perturb the metric from Honda’s theorem to a real analytic Riemannian metric, which we denote still by the same . Choosing the perturbation to be sufficiently small, there exists a unique -harmonic -form near in the same de Rham class, which we denote again by the same This is analytic by elliptic regularity.
We make now the definition of standard Calabi–Yau structures in cotangent bundles.
Definition 5.6.
Let be a closed connected orientable Riemannian manifold. Then a standard Calabi–Yau structure near the zero-section is a pair consisting of a compatible (integrable) complex structure and a nowhere-vanishing -holomorphic form relative to which is a special Lagrangian. When is oriented, we require this to agree with that induced by
We show that standard Calabi–Yau structures exist as follows.
Lemma 5.7.
Let be a closed oriented real analytic Riemannian manifold. Then there exist a disc sub-bundle and a standard Calabi–Yau structure on such that the restriction of the Kähler metric to is equal to and the complex structure is compatible with the Liouville structure of in the sense of Definition ∎
Remark 5.8.
The real analytic structure induced from the complex structure above is possibly different from that induced from the real analytic structures of and
Proof of Lemma 5.7.
Guillemin and Stenzel [30, §5] prove this except the part which has to do with that is, there exist a sufficiently small disc sub-bundle and a complex structure on compatible with the Liouville structure of and such that the restriction to of the corresponding Kähler metric is equal to Making smaller if we need, extend to a holomorphic section of which is locally of the form where are local holomorphic co-ordinates with which is defined by This is possible because every is real analytic. Since is oriented we get, making yet smaller if we need, a holomorphic volume form on which is locally of the form and whose restriction to agrees with the volume form of This completes the proof. ∎
Remark 5.9.
Guillemin and Stenzel [30, §5] do not suppose that is oriented. They show also that is invariant under the involution defined by and that is the unique complex structure satisfying this and the other conditions in the lemma.
We give now an example of non-special Lagrangian surgeries.
Theorem 5.10.
Let be a closed connected oriented manifold of and a Morse -form on whose critical points have the sof dimension which must be by the Poincaré duality of Morse–Novikov homology Then there exist a compact disc sub-bundle a closed embedded branched -graded Lagrangian and an object of supported on that is not isomorphic to any other object supported on a closed special Lagrangian (of any phase relative to any standard Calabi–Yau structure).
Proof.
Applying Theorem 5.4 to we get a harmonic -form with the same distributions of critical points as has. Applying McLean’s theorem to we get a nearby special Lagrangian such that the pair has the same distribution of indices as has. Consider now as in Theorem 5.1. Corollary 5.3 implies then that this cannot be represented by a special Lagrangian, completing the proof. ∎
Example 5.11.
Let be an even number and take to be the self-connected sum of that is, is obtained from by eliminating two points of it and attaching the handle between them. Take a Morse function with one critical point of index two of index and one of index Performing the connected sum construction at the critical point of index and that of index we obtain a -form on satisfying the hypothesis of the theorem above. There is accordingly a branched nearby Lagrangian with unobstructed. On the other hand, as there is an isomorphism so the hypotheses of Theorem 1.4 (ii) holds except the special Lagrangian condition.
Let be an object of the Fukaya category supported on and let be any point. Then the Floer cochain group is non-zero in degrees and As these two degrees differ at least by two and no differentials exist between them. So the cohomology group is non-zero in two different degrees. This shows that Lagrangians being special is essential to Lemma 3.3.
We give next an example of special Lagrangian surgeries.
Theorem 5.12.
Let and let have a Morse -form with at least one critical point of index at least one of index and none of index or Then near there exist a standard Calabi–Yau structure and relative to it a closed branched special Lagrangian.
Proof.
Following the proof of Theorem 5.10 we get a nearby special Lagrangian such that the Maslov indices of the pair has the same distributions of the Morse index of So has at least one pair intersection points, one of index and the other of index which thus satisfies the hypothesis of Joyce [38, Theorem 9.7] (one side of the equation (62) in that theorem corresponds to index critical points, and the other side of the same equation corresponds to index critical points). We can define then a special Lagrangian surgery as we want. ∎
The branched special Lagrangian in the preceding theorem is possibly immersed; any pairs of critical points of of indices and may be removed by surgery but the other critical points become transverse double points.
Example 5.13.
Take to be the selfconnected sum of then has a Morse -form with exactly two critical points, one of index and the other of index So with we can apply Theorem 5.12 and if we have so that is -unobstructed.
But there are also -obstructed special Lagrangian surgeries:
Corollary 5.14.
Let be a compact connected real analytic Riemannian manifold of dimension whose fundamental group has no non-abelian free subgroup. Let be a Morse -form on with at least one critical point of index at least one of index and none of index or Let be a disc sub-bundle equipped with a standard Calabi–Yau structure as in Lemma
If then there exists in a closed embedded -obstructed special Lagrangian which has obstructed for every Novikov field of characteristic Moreover, if is virtually solvable then the same Lagrangian has obstructed for every Novikov field.
If then there exists in a closed generically-immersed special Lagrangian which has obstructed for every Novikov field of characteristic Moreover, if is virtually solvable then the same Lagrangian has obstructed for every Novikov field.
Proof.
For (i) recall from Poincaré–Hopf’s theorem that the numbers of critical points of indices and are the same. So by Theorem 5.12 there are branched embedded special Lagrangians. Note that we can re-scale to make these special Lagrangians as close to as we like. We can then use Theorem 1.4 (iii),(iv) to show that they have obstructed in such a way as stated above. Part (ii) may be proved in the same way. ∎
Example 5.15.
For (i) take and regard it as the selfconnected sum of According to Milnor [50, Lemma 8.2] for every Morse function we can produce, by changing it locally, two additional critical points of any adjacent indices. Applying this to the height function of we get a Morse function with exactly one critical point of index exactly one critical point of index and arbitrarily many pairs of critical points of indices and So has such a Morse -form as in (i) above.
In the same way, on with there is such a Morse -form as in (ii) above with arbitrarily many pairs of critical points of adjacent indices or
Corollary 5.14 implies a statement without Lagrangians in it and perhaps of interest in itself:
Corollary 5.16.
Let be a closed connected orientable manifold of dimension with having no non-abelian free subgroup. Let be a Morse -form on with no critical points of index or Then
if we have and has no critical point at all.
Suppose now that and has at least one critical point of index and one critical point of index Then
if then either or has some critical point of index or
Proof.
We first prove (i). Since has no critical points of index or it follows that is not exact, so If this condition and the hypothesis on imply , but since the Euler characteristic of vanishes, every Morse -form on with no critical point of index or has no critical point at all, as in (i).
We then prove (ii). Since the immersed Lagrangian in Corollary 5.14 (ii) has obstructed, it follows that either or has some self-intersection points of index or So (because ) or has some critical points of index or ∎
We apply the same method to and prove
Proposition 5.17.
Let be a closed connected orientable manifold of dimension with virtually solvable. Let be a Morse -form on with no critical points of index or Then there exists a smooth function such that has no critical points at all.
Proof.
Let be two critical points of one of index and the other of index such a pair exists by Poincaré duality of Novikov homology. As in Corollary 5.14 (i) there is a double point surgery with obstructed. In dimension the bounding cochains on are critical points of the disc potential function. The relevant homology classes come from those of strips between with the right orientation. For this we use again Fukaya, Oh, Ohta Ono’s result [23, Chapter 10]; the number of holomorphic discs homologous to this strip with multiplicity one is conserved under the surgery. We prove that the number of holomorphic discs homologous to the multiplicity-one strip is exactly one. This is an integer because is an exact symplectic manifold and we can therefore use the result of Fukaya, Oh, Ohta and Ono [25]. We can also work with the Novikov field of any characteristic so the Lagrangian has obstructed in characteristic and in particular, is non-zero modulo But is any, so Denote by the graph of As the Floer complex of the pair and the Novikov complex of are quasi-isomorphic, the number of gradient trajectories from to is also one. We can then use the cancellation theorem in Morse theory [50, Theorem 5.4]; that is, we can cancel out with each other by changing in its de Rham cohomology class. Repeating this process we get a Morse -form with no critical points at all, as claimed. ∎
The hypothesis in Proposition 5.17 that should be virtually solvable is stronger than that in Corollary 5.16. It is therefore natural to ask whether the same conclusion holds under the same hypothesis on as in Corollary 5.16. This is in fact true as we show now.
Proposition 5.18.
Let the hypothesis of Corollary 5.16 hold with Then there exists a smooth function such that has no critical points at all.
For the proof we use the geometrisation theorem for -manifolds. We begin with
Lemma 5.19.
Every closed orientable -manifold satisfies one of the following five statements: (i) is finite; (ii) (iii) (iv) there exists a finite cover where is the mapping torus of an orientation-preserving diffeomorphism or (v) has a non-abelian free subgroup.
Proof.
We suppose that is infinite and prove that one of (ii)–(v) holds; we shall refer to the book [11] for those results about -manifolds which we use, without tracing back their origins in the literature. As is infinite, has no Riemannian metric of positive constant sectional curvature, so is not spherical [11, §1.7]. According to Theorem 1.11.1 of [11], either one of (ii)–(iv) above holds; or has no finite-index solvable subgroup. It thus remains to show that, if is an infinite -manifold group with no finite-index solvable subgroup, then it must have a non-abelian free subgroup.
The first step in this direction is to note that [11, p50, (C.6)] asserts that our assumption on the fundamental group implies that one of the following (a)–(d) holds: (a) is reducible [11, p3]; (b) is a hyperbolic manifold and is a linear group over [11, p51, (C.7)]; (c) is a Seifert manifold and is a linear group over [11, p52, (C.10) and (C.11)]; or (d) is a Haken manifold with an incompressible torus and (v) holds [11, p58, (C.24)]. Moreover:
In Case (a) it is well known that either (ii) holds or is a connected sum of two closed -manifolds neither of which is a sphere. In the latter case are both non-trivial (because Poincaré’s conjecture is true). If then and (iii) holds. Otherwise, has so many elements that (v) holds as follows: one of has at least two non-trivial elements and the other of has at least one non-trivial element so there is a non-abelian free group generated by
In Cases (b) and (c) we use Theorem 3.14. Since we have supposed that has no finite-index solvable subgroup, it follows then that (v) holds in Cases (b) and (c).
This completes the proof. ∎
Lemma 5.19 implies
Corollary 5.20.
Let be a closed orientable -manifold with having no non-abelian free subgroup. Then every non-zero element of may be represented by a nowhere-vanishing closed -form on
Proof.
By Lemma 5.19 one of (i)–(iv) holds. In Cases (i) and (iii) we have and there is nothing to prove. In Case (ii) we have and letting be the coordinate of we have which has certainly the desired property. In Case (iv) the mapping torus is determined by the monodromy matrix acting upon which we can classify by direct computation into the following three cases; (a) there is an integer such that (b) is conjugate in to for some or (c) has two distinct real eigenvalues, neither equal to We study these respectively as follows:
In Case (a) we can rechoose the finite cover so that from which we get a natural embedding where are the three coordinates of This expression implies the desired property of
In Case (b) we can rechoose the finite cover so that will have two eigenvalues both equal to one. Trivialising over two intervals in and applying to them the Mayer–Vietoris sequence, we get an isomorphism
(5.4) where is obtained from the projection and the right-hand side is the -invariant subspace of As has two eigenvalues both equal to one, there is a coordinate such that is invariant under and lifts to Then by (5.4) we have in which has certainly the desired property.
In Case (c) there is again an isomorphism of the form (5.4) but the right-hand side vanishes because has no eigenvalue equal to one. So in which has certainly the desired property.
This completes the proof. ∎
Appendix A Viterbo Restriction Functors
The purpose of this appendix is to prove Proposition 2.87. The strategy for the proof is essentially the same as the construction of the Viterbo functor from [8], and we shall see that there are essentially no new ideas. However, since [8] only involves exact Lagrangians, we go through the process of explaining how the energy estimates still apply to prove well-definedness of operations when allowing one of the Lagrangians to be non-exact, as long as it lies entirely inside the target of the restriction map. We discuss the energy estimates in a very general context, then explain the algebraic arguments required to reach the desired conclusions, which includes in particular recalling the formulation of the wrapped Fukaya category using linear Hamiltonians, as this will allows us to more directly use the existing construction of the Viterbo functor.
A.1 Topological and Geometric energies
In the definition of wrapped Floer homology using linear Hamiltonians, Floer complexes are defined to have generators given by Hamiltonian chords:
Definition A.1.
Let be a Liouville domain and a function. Denote by its Hamiltonian vector field on and define an -chord in to be a smooth map with and where are some Lagrangians in
The flow of gives a bijection between -chords with endpoitns and intersection points between the image of under the time- flow of and . It is possible to then recast all definitions in terms of such intersection points, but we shall not do so to keep consistency with [8].
The differential in Floer complexes in this setting is defined to be the count of solutions to the Floer equation
on a strip , with interval coordinates , boundary conditions and , and which have finite energy in the sense that
An application of Stokes’s theorem shows that this expression depends only on the homotopy class of the strip, and a second application shows that, in the case of exact Lagrangian it in fact depends only on the input and output.
To define operations on Floer complexes generated by Hamiltonian chords, one removes boundary marked points of Riemann surfaces (which would be asked to map to intersection points between Lagrangians in the framework of Section 2.28), and one counts solutions to a specific class of inhomogeneous Cauchy-Riemann equations, which in a neighbourhood of the punctures agree with the Floer equation. To be more precise, we need some terminology: for a punctured disc with ends is a non-compact bordered Riemann surface isomorphic to the closed unit disc in minus distinct boundary points indexed by in the counter-clockwise order. In particular, is given a complex structure and hence orientation; and we index the boundary arcs also by in a counter-clockwise order, which we fix by saying that the boundary arc is between the and boundary points.
Note that there exist a holomorphic embedding which parametrizes the end of and for a holomorphic embedding which parametrizes the end of We make a universal choice of these that is compatible with isomorphisms, deformations and gluings of punctured discs [8, §2.2]. Denote by the first co-ordinate and the second co-ordinate, which we call the co-ordinates. Define in the same way the co-ordinates of
We now arrive at the key additional datum which we must choose:
Definition A.2.
Let be a punctured disc with ends. We say that a -form on is sub-closed if with respect to the orientation of that is, where is some -form and an orientation -form of We assume as well that agrees, along the strip-like end labelled by an integer , with a non-negative multiple of the -form .
Most of the discussion in this Appendix can be performed using closed forms. The generality of subclosed forms is required for continuation elements (quasi-units) that enter in the choice of localisation.
The definition of the operation on the wrapped Floer complexes in this setting is given by maps
mapping the boundary conditions on a sequence of Lagrangians, which solve the equation
on punctured discs with ends, for a specific family of subclosed -forms and a family of compatible almost complex structures on parametrised by this domain, and which have finite energy in the sense that the geometric energy
is finite. We now arrive at the main notion which we shall later use as the Novikov exponent of operations on wrapped Floer complexes:
Definition A.3.
The topological energy of a solution to Equation (A.4) is
By Stokes’s theorem, the topological energy depends only on the homotopy class of the map . A manipulation using the definition of the Hamiltonian flow further shows that
Our assumptions that is non-negative, and the fact that , together imply that the topological energy is positive.
A.2 The homotopy method
Consider a pair of paths of exact Lagrangians parametrised by an interval , as well as a -parameter family of Hamiltonians on the same domain, which satisfy the following stringent assumption:
| the space of Hamiltonian -chords with endpoints on is regular for all values of , and is proper over the interval . | (A.8) |
This means in particular that we have a canonical identification between the -chords connecting and the -chords connecting for different points in the interval.
In this setting, one would like to define an associated map of Floer complexes
for different parameter values. Even in the most basic case, this map is not given by the naive identification of generators, but requires a correcting scaling factor:
Definition A.4.
Given a chord of , the relative action difference for parameters and in is the sum of the area of the rectangle swept by the -parameter family of chords interpolating between and and the difference
If are exact with primitive and respectively, there is a standard notion of given by
The relative action difference agrees in this case with the difference .
Given the above notion, we have:
Lemma A.5.
If the moduli spaces of solutions to Floer’s equation submerse over the interval , there is a chain-isomorphism of Floer complexes given by
The homotopy method (going back to [23, Section 19]) allows one to drop the condition that all solutions over the are regular: in this case, one considers a sequence of intermediate times, and a sequence of solutions to Floer’s equations at these times, with asymptotic conditions which successively match in the sense that if the output of is , then the input of is . In this case, the expression for Equation (A.9) is a sum of terms, one for each such configuration, with the property that the curves have virtual dimension . The -adic valuation of the contribution of such a configuration is the sum of the topological energies of the strips , and the contribution of Equation (A.12) for each subinterval. Using the positivity of topological energy, we conclude:
Lemma A.6.
If a sequence as above contributes to the homotopy map, then the valuation of its contribution is bounded below by the sum of the relative action differences , over all intermediate orbits. ∎
This result, and its generalisations, will be used to show that certain operations must be -adically convergent, as a consequence of geometric choices that ensure that the relative action differences are bounded below by a positive multiple of . The purpose of the next sections is to indicate which configurations of curves are used to define the operations of interest, at which point a version of the above Lemma will yield the desired convergence.
A.3 Viterbo restriction: the linear part
In Section 2.9, we considered a wrapped Fukaya category , obtained by localisation. Here, we give a slightly different account, replacing Lagrangian isotopies with Hamiltonian continuation maps. For the later discussion, it is also convenient to allow to be either a Liouville manifold (with complete Liouville flow) or a compact Liouville domain (with boundary along which the flow is inward pointing). We use the term end of to refer either to an infinite end or to a neighbourhood of the boundary.
We begin by fixing a finite set of Lagrangians in , which along the end are invariant under the Liouville flow. We then choose a Hamiltonian on so that, for all non-negative integers , all Hamiltonian -chords with endpoints on a pair of Lagrangians in this set are non-degenerate and are disjoint from the end. We can then define a variant of the wrapped category: objects are pairs with a natural number, and, before localisation, there are no morphisms from to unless either (i) or (ii) and , in which case it is given by the Floer group (we are omitting the choices of bounding cochains and local systems for the notation). We then obtain a category by localising at continuation elements in , thought of as the morphisms form to .
Let us now consider the inclusion of a subdomain in a complete Liouville domain. We fix now a finite set of Lagrangians in whose intersection with are objects of , so that if an element of this set has non-empty boundary, it is exact outside of , and admits a primitive which is constant near the boundary. For the choices required to define the Viterbo restriction functor, we start by restating a basic result from [8], which ensures positivity of actions:
Lemma A.7 (Abouzaid–Seidel [8, Lemma 7.3]).
There is a function be a function which on the collar agrees with the radius function of and near agrees with the radius function of , and such that at every point of where is the vector field on defined in Definition 2.28.
Corollary A.8.
Given any pair of Lagrangians in which are exact in the complement of , there exists a constant such that for every and for every -chord with we have ∎
Up to rescaling, we may assume that in the above result.
The key idea in the construction of the Viterbo functor in [8, §4] is a family of Hamiltonian parametrised by , for which the actions of chords in are scaled by , while those outside of remain constant. This is achieved by rescaling along the Liouville flow inside . We shall instead consider a family parameterised by , of Hamiltonians which are constant in and have the effect of rescaling outside . For the next definition, denote by the flow of at time
Definition A.9.
Given a function on which is linear near and on the collar of , and a real number define a smooth function by on and on
Given a Lagrangian which is stable under the Liouville flow on a neighbourhood of and on the end of , we similarly define to agree with in the region , and to be given by the image of under in its complement.
Note that, if lies in , this means that does not depend on . We have the following basic computations:
Lemma A.10.
Given a pair of Lagrangians in satisfying the properties in Definition A.9, and a function so that there is no -chord with endpoints and in a neighbourhood of and in the cylindrical end of , the flow maps -chords with endpoints on bijectively to -chords with endpoints on . ∎
In this setting, every chord from to lies either in or in We call this an inner chord in the former case and an outer chord in the latter case.
We may thus use the notation of Section A.2, and write for the image of an -chord , with endpoints and . Given how the definitions of and involve the Liouville flow, we have:
Lemma A.11.
For any inner -chord with endpoints on , the relative action difference associated to an interval vanishe. There is a constant , so that the relative action difference is bounded below by if is an outer chord. ∎
Assuming that all such chords are non-degenerate, we may now setup the homotopy map as a map of chain complexes
While there is some flexibility in the definition of such a map, we assume that our the homotopy restricts to constant data on ; this is a statement about the choice of almost complex structure not varying with respect to in this region..
Lemma A.12.
Assuming that satisfies the conditions of Definition A.7, there is a value so that the outer chords generate a subcomplex of whenever . The quotient complex is isomorphic to , and the map
obtained by composing the homotopy map with the projection does not depend on in this range.
Proof.
Lemma A.11 implies that there is a choice of past which all outer chords have actions larger than inner chords. In that case, there can be no holomorphic strip with outer input and inner output, so we conclude both that outer chords form a subcomplex and that the components of the continuation map taking value in inner chords do not depend on past this value. The last statement is an application of the integrated maximum principle. ∎
A.4 Viterbo restriction: compatibility with products
Let us now consider a compact, not necessarily exact, Lagrangian , equipped with brane data and a bounding cochain which we supress from the notation, as well as a exact Lagrangian in for which the Viterbo restriction map to the wrapped Fukaya category of is defined. Consider a Hamiltonian as in Definition A.9, and pick an almost complex structure on , of contact type along the end, which extends an almost complex structure on of contact type near . The integrated maximumm principle of [8] implies:
Lemma A.13.
The homotopy method isomorphism is also given by the canonical identification of generators. This identification also induces an isomorphism for all . ∎
We next assume that we have a pair of exact Lagrangians , for which Viterbo restriction is defined, and consider the diagram
The method of Section A.2 extends to give a homotopy in this diagram. The key moduli spaces to consider are moduli spaces of pseudo-holomorphic triangles, with boundary conditions for , and equipped with the subclosed -form which agrees, in strip-like ends, with and near the punctures with adjacent edges labelled and . We choose this -form to be independent of the parameter in the interval , i.e. to be the one used to define the Floer product. The definition requires as well a choice of a family of almost complex structure, whose restriction to we assume not to depend on . With this, we obtain a moduli space for each triple of Hamiltonian chords with starting on and ending on , and starting on and ending on . The following result then follows from the integrated maximum principle:
Lemma A.14.
If lies in the subdomain , then is the product of the interval with the space of pseudo-holomorphic triangles. ∎
The homotopy in diagram (A.15) is given by considering a sequence of real numbers , a sequence of chords with endpoints and , and a collection of maps from Riemann surfaces to so that the following properties hold:
1.
For , the map is a solution of Floer’s equation, with boundary conditions and asymptotic conditions and .
2.
The map is an element of the moduli space a pseudo-holomorphic triangle with boundary conditions , and asymptotic conditions .
The -adic valuation of the contribution of each such configuration is given by the sum of the topological energy of each disc with the sum of the relative action differences for each intermediate chord . Using the fact that the topological energy is positive, as well as Lemma A.7, we now arrive at the main estimate:
Lemma A.15.
The -adic valuation of a configuration contributing to the homotopy is bounded below by a constant multiple of (the element of corresponding to the last disc. ∎
The above result implies that the homotopies in Diagram (A.15) satisfy the following properties: if , then the difference between the homotopies associated to and has valuation bounded below by a constant multiple of . This means that we obtain a well-defined map by letting go to infinity:
Corollary A.16.
The following diagram, in which the vertical map on the right is the Viterbo restriction map, commutes up to homotopy:
The above result says that, at the cohomological level, the Viterbo restriction map is compatible with the Floer product for a triple with compact and exact for , extending the result of [8] for triples of exact Lagrangians. There is a completely analogous argument showing compatibility of restriction with the Floer product for triples . The remaining cases, in which two of the three Lagrangians are compact, or in which the triple is , are even easier to address, as the restriction map is just given by the identity on all Floer groups. We conclude:
Corollary A.17.
The cohomology-level restriction functor from the exact Fukaya category of to that of constructed in [8] extends to the cohomological wrapped Fukaya category with objects which are either (i) exact Lagrangians equipped with primitive which is constant on , or (ii) compact Lagrangians in . ∎
A.5 Wrapped categories using linear Hamiltonians
In order to extend Corollary A.17 to a statement about categories, need an auxiliary definition of wrapped Fukaya categories. To this end, we introduce variants of the category from Hypothesis 2.40:
Definition A.18.
The curved categories and have objects pairs consisting of an object of and a non-negative integer. Let be underlying Lagrangians of and regard these as branes, leaving local systems out of the notation. For define
Define to be the same for and to be for Denote by the localization of with respect to continuation element from Definition 2.75, as well as the elements associated to the homotopy from to . We denote by the localization of with respect to the same continuation morphisms.
We define on and the structures in the same way as in §2.4, choosing domain-dependent almost complex structures and using (A.4) in place of the plain Cauchy–Riemann equation.
There is then an inclusion mapping each underlying Lagrangians to , and which is the identity on morphism spaces. Theh construction of homotopy units [23, Chapter 8] defines an functor acting as the identity between the same object sets and whose operations factor through the identity map for each pair of distinct objects , and whose linear term maps to a representative of the unit.
Using Proposition 2.73 we have:
Lemma A.19.
The functors induce, after adding bounding cochains and passing to localizations, homotopy equivalences . ∎
Composing the functor and a homotopy inverse we get a homotopy equivalence from to the subcategory of with object set The latter will in practice be useful in the next subsection. Denote by and the subcategories with object set which we identify with
Remark A.20.
Here we use localizations repeatedly, and for the definition of or we only use sub-closed -forms for the definition of the continuation elements, so that we do not require popsicles as in the original paper [8]. This is one of the main advantages of the localization approach.
A.6 Proof of Proposition 2.87
We now outline how the results of Appendix A.4 can be extended to obtain an restriction functor:
Proof of Proposition 2.87.
The restriction functor is given by a parametrized count of pseudo-holomorphic curves, defining maps
for each sequence of Lagrangians defining objects of our category.
The geometric construction is essentially the same as that of Abouzaid–Seidel [8, §4] except that we use virtual counts and that we change the Hamiltonians in [8, §4.2] as discussed in Appendix A.3, where we defined the linear term of the functor. The quadratic term is the homotopy introduced in Appendix A.4.
The homotopy is produced from counts of cascades, which are configuration of pseudo-holomorphic curves associated to the vertices of a directed tree, so that the curve labelled by a vertex has boundary on a subsequence , for some parameter . The key point to have in mind from the construction in [8] is that, if with respect to the ordering associated to the tree, then there is chord so that the label of the output at is , and the corresponding input at is .
Exactly as in [8], the integrated maximum principle together with a choice of data that is -independent in ensures that the only non-trivial contributions are associated to vertices where at least one input is an outer orbit.
The proof that is -adically convergent then amounts to showing that the valuation of the contribution of any cascade is bounded below by a constant multiple of . This is exactly the same argument as in Appendix A.4: each vertex contributes a non-negative amount to the valuation, because of the positivity of topological energy, while each edge contributes the action differences of the corresponding homotopy map. Since every vertex must have at least one incoming outer edge, the result then follows from the action estimate in Appendix A.3.
Using the functors and choosing homotopy inverses we get an functor . ∎